### v1

# ROSE: Register Assisted General Time Series Forecasting with Decomposed Frequency Learning

Abstract
With the increasing collection of time series data from various domains, there arises a strong demand for general time series forecasting models pre-trained on a large number of time-series datasets to support a variety of downstream prediction tasks. Enabling general time series forecasting faces two challenges: how to obtain unified representations from multi-domian time series data, and how to capture domain-specific features from time series data across various domains for adaptive transfer in downstream tasks. To address these challenges, we propose a Register Assisted General Time Series Forecasting Model with Decomposed Frequency Learning (ROSE), a novel pre-trained model for time series forecasting. ROSE employs Decomposed Frequency Learning for the pre-training task, which decomposes coupled semantic and periodic information in time series with frequency-based masking and reconstruction to obtain unified representations across domains. We also equip ROSE with a Time Series Register, which learns to generate a register codebook to capture domain-specific representations during pre-training and enhances domain-adaptive transfer by selecting related register tokens on downstream tasks. After pre-training on large-scale time series data, ROSE achieves state-of-the-art forecasting performance on 8 real-world benchmarks. Remarkably, even in few-shot scenarios, it demonstrates competitive or superior performance compared to existing methods trained with full data.
1 Introduction
Time Series Forecasting plays a pivotal role across numerous domains, such as energy, smart transportation, weather, and economics [1]. However, training specific models in deep learning for each dataset is costly and requires tailored parameter tuning, whose prediction accuracy may be limited due to data scarcity [2]. One solution is to pre-train a general model on diverse time series datasets and fine-tune it with a few data for different downstream scenarios. However, applying this paradigm to heterogeneous time series data presents challenges.
Obtaining a unified representation from time series data across various domains is challenging. Time series from each domain involve complex temporal patterns, composed of multiple frequency components combined with each other [3], which is frequency superposition. As shown in Figure 1(a), different frequency components contain distinct semantic information. For example, low and high-frequency components represent long-term trends and rapid variations, respectively [4]. Additionally, different frequency components also contain cross-period information, revealing the interactions between short and long-term patterns in the time series. Thus, the frequency superposition leads to coupled semantic and periodic information in the time domain. Furthermore, large-scale time series data from different domains introduce even more complex temporal patterns and frequency diversity. Existing pre-training frameworks [5, 6, 7], such as mask modeling and contrastive learning, were proposed to learn a unified representation from time series data. However, these methods overlook the frequency diversity and complexity exhibited in heterogeneous time series that come from various domains, making it difficult to capture intricate patterns, thus limiting their generalization capabilities.
Adaptive transferring information from multi-domain time series to specific downstream scenarios presents a challenge. Multi-source time series data originate from various domains [8], whose data exhibit domain-specific information [2]. Information from the same or similar domain as the target domain is useful for improving the model’s effectiveness in the target task [9]. However, as shown in Figure 1(a), existing time series pre-training frameworks [8, 2, 10] focus mainly on capturing generalized features during pre-training and overlook domain-specific features. Thus, they only transfer the same generalized representation to different target domains, called direct transfer, which limits the model’s effectiveness in specific downstream tasks, as shown in Figure 1(b). Therefore, it is necessary to learn domain-specific information during pre-training and adaptively transfer the specific representations to each target domain, called adaptive transfer. Realizing adaptive transfer poses two difficulties: 1) capturing domain-specific information in pre-training. 2) adaptive use of domain-specific information in various downstream tasks.
Figure 1: (a) Pre-training on multi-domain datasets that exhibit combined frequency. Existing general time series forecasting models only extract generalized representations for direct transfer to various downstream target domains. We propose to learn generalized and specific representations during pre-training, and adaptively transfer them to each target domain. (b) The t-SNE visualization of the hidden representations after direct transfer and adaptive transfer: In direct transfer, representations of different domains are mixed, but in adaptive transfer, they show a clear clustering pattern.
To address these challenges, we propose a register assisted general time series forecasting model with decomposed frequency learning (ROSE). First, we propose Decomposed Frequency Learning that learns generalized representations to solve the issue with coupled semantic and periodic information. We decompose individual time series using the Fourier transform with a novel frequency-based mask method, and then convert it back to the time domain to obtain decoupled time series for reconstruction. It makes complex temporal patterns disentangled, thus benefiting the model to learn generalized representations. Second, we introduce Time Series Register (TS-Register) to learn domain-specific information in multi-domain data. By setting up a register codebook, we generate register tokens to learn each domain-specific information during pre-training. In a downstream scenario, the model adaptively selects vectors from the register codebook that are close to the target domain of interest. During fine-tuning, we incorporate learnable vectors into the selected register tokens to complement target specific information to perform more flexible adaptive transfer. As shown in Figure 1(b), adaptive transfer successfully utilizes domain-specific information in multi-domain time series, which contributes to the model’s performance in target tasks. The contributions are summarized as follows:
  • We propose ROSE, a novel general time series forecasting model using multi-domain datasets for pre-training and improving downstream fine-tuning performance and efficiency.
  • We propose a novel Decomposed Frequency Learning that employs multi-frequency masking to decouple complex temporal patterns, empowering the model’s generalization capability.
  • We propose TS-Register to capture domain-specific information in pre-training and enable adaptive transfer of target-oriented specific information for downstream tasks in fine-tuning.
  • Our experiments with 8 real-world benchmarks demonstrate that ROSE achieves state-of-the-art performance in full-shot setting and achieves competitive or superior results in few-shot setting, along with impressive transferability in zero-shot setting.
2 Related work
2.1 Time series forecasting
Traditional time series prediction models like ARIMA [11], despite their theoretical support, are limited in modeling nonlinearity. With the rise of deep learning, many RNN-based models [12, 13, 14] have been proposed, modeling the sequential data with an autoregressive process. CNN-based models [15, 16] have also received widespread attention due to their ability to capture local features. MICN [17] utilizes TCN to capture both local and global features, while TimesNet [18] focuses on modeling 2D temporal variations. However, both RNNs and CNNs struggle to capture long-term dependencies. Transformer-based models [3, 6, 19, 20, 21], with their attention mechanism, can capture long dependencies and extract global information, leading to widespread applications in long-time series prediction. However, this case-by-case paradigm requires meticulous hyperparameter design for different datasets, and its predictive performance can also be affected by data scarcity.
2.2 Time series pre-training
Masked modeling [5, 22, 23] and contrastive learning [24, 25, 26] have attracted significant attention in the time series domain. PatchTST [6] proposes to capture local semantic information by predicting masked subsequence patches. LaST [27] explores using variational inference to decompose time series into representations of seasonality and periodicity within a latent space. PITS [7] combines masked modeling and contrastive learning to obtain better generalization representations. However, the above methods are limited to one-to-one pre-training and fine-tuning, overlooking the frequency diversity and complexity within multi-domain data, thus making them challenging to directly apply to pre-training for multi-domain time series.
Multi-source time series pre-training has recently garnered widespread attention [28, 29, 30]. MOMENT [31] and MOIRAI [8] adopt a BERT-style pre-training approach, while Timer [2] and PreDcT [32] use a GPT-style pre-training approach, giving rise to improved performance in time series prediction. However, the above methods overlook domain-specific information from multi-source data, thus limiting the performance of the models. Different from previous approaches, ROSE pre-trains on large-scale data from various domains and it considers both generalized representations and domain-specific information, which facilitates flexible adaptive transfer in downstream tasks.
3 Methodology
Problem Definition. Given a multivariate time series $\mathbf{X}_{t}=\{\mathbf{x}_{t-L:t}^{i}\}_{i=1}^{C}$, where each $\mathbf{x}_{t-L:t}^{i}\in\mathbb{R}^{L}$ is a sequence of observations. $L$ denotes the look-back window and $C$ denotes the number of channels. The forecasting task is to predict the future values $\mathbf{\hat{Y}}_{t}=\{\mathbf{\hat{x}}_{t:t+F}^{i}\}_{i=1}^{C}$, where $F$ denotes the forecast horizon. $\mathbf{Y}_{t}=\{\mathbf{x}_{t:t+F}^{i}\}_{i=1}^{C}$ is the ground truth of future.
The general time series forecasting model is pre-trained with multi-source datasets $\mathbf{D}_{\text{pre-train}}=\{(\mathbf{X}_{t}^{j},\mathbf{Y}_{t}^{j})\}^{N}_{j=1}$, where $N$ is the number of datasets. For the downstream task, the model is fine-tuned with a training dataset $\mathbf{D}_{\text{train}}=\{(\mathbf{X}_{t}^{\text{train}},\mathbf{Y}_{t}^{\text{train}})\}$, and is tested with $\mathbf{D}_{\text{test}}=\{(\mathbf{X}_{t}^{\text{test}},\mathbf{Y}_{t}^{\text{test}})\}$ to predict $\mathbf{\hat{Y}}_{t}^{\text{test}}$, where $\mathbf{D}_{\text{pre-train}}$, $\mathbf{D}_{\text{train}}$ and $\mathbf{D}_{\text{test}}$ are pairwise disjoint. Alternatively, the model could be directly tested using $\mathbf{D}_{\text{test}}$ without fine-tuning with $\mathbf{D}_{\text{train}}$ to predict $\mathbf{\hat{Y}}_{t}^{\text{test}}$.
3.1 Architecture
Figure 2: The model architecture of ROSE.
The time series forecasting paradigm of ROSE contains two steps: pre-training and fine-tuning. First, the model is pre-trained on large-scale datasets from various domains, with two pre-training tasks: reconstruction and prediction. We set up the self-supervised reconstruction task to help the model understand time series comprehensively and set the supervised prediction task to enhance the model’s few-shot learning ability. Then, ROSE fine-tunes with a target dataset in the downstream scenario.
As shown in Figure 2, ROSE employs the Encoder-Decoder architecture to model time series. The backbone consists of multiple Transformer layers to process sequential information and effectively capture temporal dependencies [33]. ROSE is pre-trained in a channel-independent way, which is widely used in time series forecasting [6].
Input Representations. To enhance the generalization of ROSE for adaptive transferring from multi-domains to different target domains, we model the inputs $\mathbf{x}$ with patch tokens and register tokens. Patch tokens are obtained by partitioning the time series using patching layers [6], to preserve local temporal information. Register tokens are obtained by linear mapping and clustering each of the entire time series input into discrete embedding, to capture domain-specific information, which will be introduced in Section 3.3.
3.2 Decomposed frequency learning
As shown in Figure 1, time series data are composed of multiple combined frequency components, resulting in overlapping of different temporal variations. Furthermore, low-frequency components typically contain information about overall trends and longer-scale variations, and high-frequency components usually contain information about short-term fluctuations and shorter-scale variations, therefore, understanding time series from low and high frequencies separately benefits general time series representation learning. Based on the observations above, we propose decomposed frequency learning based on multi-freq masking to understand the time series from multiple-frequency perspectives, which enhances the model’s ability to learn generalized representations.
Figure 3: An illustration of decomposed frequency learning. Based on the pre-defined thresholds, we randomly apply low/high-frequency masking to the time series in the frequency domain and then transform it back to the time domain for reconstruction.
Multi-freq masking. Given a time series $\mathbf{x}\in\mathbb{R}^{L}$, we utilize the Real Fast Fourier Transform (rFFT) [34] to transform it into the frequency domain, giving rise to $\mathbf{x}_{\mathrm{freq}}\in\mathbb{R}^{L/2}$.
$\mathbf{x}_{\mathrm{freq}}=\mathrm{rFFT}(\mathbf{x}).$ (1)
To separately model high-frequency and low-frequency information in time series, we sample $K_{\mathrm{f}}$ thresholds ${\tau_{1},\tau_{2},\tau_{3},...,\tau_{K_{\mathrm{f}}}}$ and $K_{\mathrm{f}}$ random numbers ${\mu_{1},\mu_{2},\mu_{3},...,\mu_{K_{\mathrm{f}}}}$ for multi-frequency masks, where $\tau\in\mathrm{Uniform}(0,a)$, $a<L/2$ , and $\mu\in\mathrm{Bernoulli}(1,p)$. Each pair of $\tau_{i}$ and $\mu_{i}$ corresponds to the $i_{th}$ frequency mask. This generates a mask matrix $\mathbf{M}\in\{0,1\}^{K_{\mathrm{f}}\times L/2}$, where each row corresponds to the $i_{th}$ frequency mask, each column corresponds to the $j_{th}$ frequency, and each element $m_{ij}$ is 0 or 1, meaning that the $j_{th}$ frequency is masked with the $i_{th}$ frequency mask or not.
$m_{ij}=\left\{\begin{aligned} &\quad\mu_{i}&,if\,\ \mathbf{x}_{\mathrm{freq}_{j}}<\tau_{i}\\ &(1-\mu_{i})&,if\,\ \mathbf{x}_{\mathrm{freq}_{j}}>\tau_{i}\end{aligned}\right.,$ (2)
where $\tau_{i}$ and $\mu_{i}$ denote the threshold and random number for the $i_{th}$ frequency domain mask, $\mathbf{x}_{\mathrm{freq}_{j}}$ represents the $j_{th}$ frequency of $\mathbf{x}_{\mathrm{freq}}$. If $\mu_{i}=1$, it indicates masking of the high frequencies, whereas if $\mu_{i}=0$, it signifies the masking of the low frequencies.
After obtaining the mask matrix $\mathbf{M}$, we replicate $\mathbf{x}_{\mathrm{freq}}$ $K_{\mathrm{f}}$ times to get the $\mathbf{X}_{\mathrm{freq}}\in\mathbb{R}^{K_{\mathrm{f}}\times L/2}$ and perform element-wise Hadamard product with the mask matrix $\mathbf{M}$ to get masked frequency of time series. Then, we use the inverse Real Fast Fourier Transform (irFFT) to convert the results from the frequency domain back to the time domain and get $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}=\{\mathbf{x}_{\mathrm{mask}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$, where each $\mathbf{x}_{\mathrm{mask}}^{i}\in\mathbb{R}^{L}$ corresponding to masking with a different threshold $\tau_{i}$.
$\mathbf{X}_{\mathrm{mask}}=\mathrm{irFFT}(\mathbf{X}_{\mathrm{freq}}\odot\mathbf{M}).$ (3)
Representation learning. After obtaining the $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}$, we divide each sequence $\mathbf{x}_{\mathrm{mask}}^{i}$ into $P$ non-overlapping patches, and use a linear layer to transforming them into $P$ patch tokens, and thus we get $\mathcal{X}_{\mathrm{mp}}=\{\mathbf{X}_{\mathrm{mp}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$ to capture general information, where each $\mathbf{X}_{\mathrm{mp}}^{i}\in\mathbb{R}^{P\times D}$, and $D$ is the dimension for each patch token. We replicate the register tokens $\mathbf{X}_{\mathrm{u}}$ $K_{\mathrm{f}}$ times to get $\mathcal{X}_{\mathrm{u}}\in\mathbb{R}^{K_{\mathrm{f}}\times N_{\mathrm{r}}\times D}$, where $\mathbf{X}_{\mathrm{u}}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ is obtained by inputting the original sequence into the TS-Register, as detailed in Section 3.3. Then, we concatenate the patch tokens $\mathcal{X}_{\mathrm{mp}}$ with the register tokens $\mathcal{X}_{\mathrm{u}}$, and feed them into the transformer encoder to obtain the representation of each masked series. These representations are then aggregated to yield a unified representation $\mathbf{S}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$:
$\mathbf{S}=\mathrm{Aggregator(Encoder(Concatenate(\mathcal{X}_{mp},\mathcal{X}_{u})))}.$ (4)
Reconstruction task. After obtaining the representation $\mathbf{S}$, we feed it into the reconstruction decoder and ultimately reconstruct the original sequence $\mathbf{\hat{x}}\in\mathbb{R}^{L}$ through the reconstruction head. As frequency domain masking affects the overall time series, we compute the Mean Squared Error (MSE) reconstruction loss for the entire time series.
$\mathcal{L}_{\mathrm{reconstruction}}=\mathrm{||\mathbf{x}-\hat{\mathbf{x}}||_{2}^{2}}.$ (5)
3.3 Time series register
By decomposed frequency learning, we can obtain the domain-general representations, and additionally, we propose the TS-Register that learns register tokens as the domain-specific information from the multi-domain datasets for adaptive transfer. It clusters domain-specific information from the multi-domain datasets into register tokens and stores such domain-specific information in the register codebook during pre-training. Then, it adaptively selects similar information from the register codebook to enhance the target domain during fine-tuning.
We set up a randomly initialized register codebook $\mathbf{E}\in\mathbb{R}^{H\times D_{\mathrm{r}}}$ with $H$ cluster center vectors $\mathbf{e}_{i}\in\mathbb{R}^{D_{\mathrm{r}}},i\in\{1,2,…,H\}$. Each of input time series $\mathbf{x}\in\mathbb{R}^{L}$ is projected into a data-dependent embedding $\mathbf{x}_{\mathrm{e}}\in\mathbb{R}^{D_{\mathrm{r}}}$ through a linear layer.
Pre-training stage. As shown in Figure 2(b), we use the register codebook to cluster these data-dependent embeddings, which generate domain-specific information, and store them in pre-training. Specifically, We find a cluster center vector $\mathbf{e}_{\delta}$ from the register codebook $\mathbf{E}$ where we use $\delta$ to denote the cluster that the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ belongs to, as shown in Equation 6.
$\delta=\underset{j=1:H}{\arg\operatorname{min}}\left\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{j}\right\|_{2}.$ (6)
$\mathcal{L}_{\text{register}}=\|\text{sg}(\mathbf{x}_{\mathrm{e}})-\mathbf{e}_{\delta}\|^{2}_{2}+\|\mathbf{x}_{\mathrm{e}}-\text{sg}(\mathbf{e}_{\delta})\|^{2}_{2}.$ (7)
To update the cluster center vectors in the register codebook $\mathbf{E}$ that represent the domain information of the pre-trained datasets, we set the loss function shown in Equation 7 that minimizes the distance between the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ and the cluster center $\mathbf{e}_{\delta}$. $\text{sg}()$ refers to the stop gradient operation, and is used to solve the problem that the gradient of the $\arg\operatorname{min}$ function cannot be backpropagated [35]. The first term is used to update the register codebook $\mathbf{E}$, and the second term is used to update the parameters of the linear layer that learns $\mathbf{x}_{\mathrm{e}}$.
In this way, the vectors in the register codebook $\mathbf{E}$ cluster the embeddings of different data and learn the domain-specific centers for pre-trained datasets, which can represent domain-specific information. As a vector in the register codebook $\mathbf{E}$, $\mathbf{e}_{\delta}$ represents the domain-specific information for input $\mathbf{x}$. $\mathbf{e}_{\delta}$ is invariant under small perturbations in $\mathbf{x}_{\mathrm{e}}$ that represents $\mathbf{x}$, which promotes better representation of domain-specific information and robustness of the vectors in the register. This also avoids their over-reliance on detailed information about specific datasets [36].
The cluster center vector $\mathbf{e}_{\delta}$ is then patched into $\mathbf{X}_{\mathrm{u}}\in\mathbb{R}^{N_{\text{r}}\times D}$, where $N_{\text{r}}$ is the number of the register tokens and $D$ is the dimensionality of transformer latent space. $\mathbf{X}_{\mathrm{u}}$ is called register tokens, which are used as the prefix of the patch tokens $\mathbf{X}_{\mathrm{p}}\in\mathbb{R}^{P\times D}$ and input for the transformer encoder to provide domain-specific information.
Fine-tuning stage. As shown in Figure 2(c), after obtaining a register codebook $\mathbf{E}$ that contains domain-specific information through pre-training, we freeze the register parameters to adaptively use this domain-specific information in the downstream task.
Since the target domain may not strictly fit one of the upstream domains, learning the embedding of the downstream data employs a top-k strategy that selects top-k similar vectors in the register codebook. As shown in Equation 8, the embedding of input time series $\mathbf{x}_{\mathrm{e}}$ picks the $k$ nearest vectors in the register codebook $\mathbf{E}$, and uses their average as $\bar{\mathbf{e}}_{k}$ to represent the domain-specific information from the pre-train stage. $\bar{\mathbf{e}}_{k}$ is also patched into $\mathbf{X}_{\mathrm{d}}\in\mathbb{R}^{N_{\text{r}}\times D}$ and is used as the register tokens.
$\bar{\mathbf{e}}_{k}=\frac{1}{k}\sum_{i=1}^{k}\mathbf{e}_{\delta_{i}},\ \ \left\{\delta_{1},\cdots,\delta_{k}\right\}=\underset{j=1:H}{\arg\operatorname{Topk}}(\frac{1}{\left\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{j}\right\|_{2}}).$ (8)
Since the downstream data has its own specific information at the dataset level in addition to the domain level, this may not be fully represented by the domain information obtained from the pre-trained dataset alone. Therefore, we additionally set a matrix $\mathbf{A}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ to adjust $\mathbf{X}_{\mathrm{d}}$ to complement the specific information of downstream data. Since the pre-trained model has a very low intrinsic dimension [37], in order to get better fine-tuning results, $\mathbf{A}$ is set as a low-rank matrix:
$\mathbf{A}=\mathbf{u}\times\mathbf{v^{\mathrm{T}}},$ (9)
where $\mathbf{u}\in\mathbb{R}^{N_{\mathrm{r}}}$ and $\mathbf{v}\in\mathbb{R}^{D}$, and only the vectors $\mathbf{u}$ and $\mathbf{v}$ need to be retrained in the fine-tuning step. As illustrated in Equation 10, the register token $\mathbf{X}_{\mathrm{r}}$ of the downstream scenario is obtained by doing the Hadamard product of $\mathbf{X}_{\mathrm{d}}$, which represents the domain-specific information obtained at the pre-train stage, and $\mathbf{A}$, which represents the downstream dataset-specific information.
$\mathbf{X}_{\mathrm{r}}=\mathbf{X}_{\mathrm{d}}\odot\mathbf{A}.$ (10)
3.4 Training
To improve the model’s prediction ability when only few samples are available downstream, we co-train supervised prediction with self-supervised reconstruction that uses multi-frequency mask to learn unified features that are more applicable to the downstream prediction task.
[…]
$\mathcal{L}_{\text{prediction}}=\sum_{F\in\{96,192,336,720\}}||{\mathbf{Y}_{F}-{\mathbf{\hat{Y}}_{F}}}||_{2}^{2}.$ (11)
Pre-training. The reconstruction task learns generalized features through the transformer encoder and reconstruction decoder. To utilize these features for the prediction task, the parameters of the reconstruction decoder are copied to the prediction decoder during forward propagation. To avoid prediction training affecting the generalization performance of the model, the gradients of the prediction heads are skipped at back-propagation. The overall pre-training loss is shown in Equation 12.
$\mathcal{L_{\text{pre-train}}}=\mathcal{L}_{\text{reconstruction}}+\mathcal{L}_{\text{prediction}}+\mathcal{L}_{\text{register}}.$ (12)
Fine-tuning. We only perform a prediction task in fine-tuning. Patch tokens $\mathbf{X}_{\text{p}}$ are concatenated with the adjusted register tokens $\mathbf{X}_{\mathrm{r}}$. For a downstream task with a fixed prediction length, we use the corresponding pre-trained prediction head to fine-tune the model.
4 Experiments
Pre-training datasets. Large-scale datasets are crucial for pre-training a general time series forecasting model. In light of this, we gather a considerable amount of publicly available datasets from various domains, including energy, nature, health, transport, web, and economics, etc. The details of these datasets are shown in the Appendix A.1. To enhance data utilization, we downsampled fine-grained datasets to coarser granularity, resulting in approximately 886 million time points.
Evaluation datasets. To conduct comprehensive and fair comparisons for different models, we conducted experiments on eight real-world benchmarks as the target datasets, including Weather, Solar, Traffic, Electricity, and ETT (4 subsets), which cover multiple domains.
Baselines. We select the state-of-the-art models as our baselines, including Transformer-based models: Timer [2], iTransformer [20], and PatchTST [6]; CNN-based model: TimesNet [18]; MLP-based models: FITS [38], TIDE [39], and DLinear [40]. It is worth noting that Timer is a recent general time series model with a GPT-style architecture. Since the size of the look-back window can affect the performance of different models, we choose the look-back window size in $\{96,336,512,720\}$ for each baseline that achieves its best results for fair comparisons.
Setup. Consistent with previous works, we adopted Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics. All methods predict the future values with lengths $F=\{96,192,336,720\}$. More implementation details are presented in the Appendix A.3.
4.1 Main results
Full-shot setting. We first pre-train ROSE on our pre-training datasets and then fine-tune with full downstream data (full-shot setting). Similarly, all baselines are also run in the same full-shot setting. As shown in Table 1, we also present the results of the ROSE in 10% few-shot setting, where the model uses the same validation and test data but is only fine-tuned with 10% training data of each target dataset. Key observations are summarized as follows. First, as a general forecasting model, ROSE achieves superior performance compared to the seven state-of-the-art baselines with full-data training, especially Timer, which shows that our decomposed frequency learning and register help to learn generalized representations from large-scale datasets and adaptively transfer the multi-domain information to specific downstream scenarios. Second, we observe that ROSE achieves competitive performance in 10% few-shot setting, even compared to baselines trained with full data. This observation validates the transferability of ROSE pre-trained with large multi-source data.
Figure 4: The forecasting results of ROSE obtained by training from scratch and fine-tuning from the pre-trained model. The right, upper corner is the best case.
Few-shot setting. We also compared all models under the 10% few-shot setting in Table 9 in Appendix H.2. We can see that ROSE performs the best on all eight datasets in the scenarios where the training data are rare in the target domain, surpassing extensive advanced models. Figure 4 shows the results of pre-trained ROSE and ROSE trained from scratch on the Solar and ETTh1 with different fine-tuning data percentages, noting the best baselines in full-shot setting. The pre-trained ROSE shows stable, superior performance even with limited fine-tuned samples. Specifically, the pre-trained ROSE achieves performance that surpasses the SOTA with only 5% for Solar and 1% data for ETTh1. Moreover, compared to the ROSE trained from scratch, the pre-trained ROSE exhibits a slower decline in predictive performance with the reduction of fine-tuning data, demonstrating the impressive generalization ability of ROSE through large-scale pre-training.
Table 1: The results for ROSE in full-shot setting and 10% few-shot setting, compared with other methods in full-shot setting. The average results of all predicted lengths are listed here.
Models ROSE ROSE 10% Timer ITransformer PatchTST FITS TimesNet TiDE DLinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 0.341 0.367 0.349 0.372 0.436 0.456 0.407 0.410 0.353 0.382 0.355 0.379 0.400 0.406 0.355 0.378 0.357 0.379
ETTm2 0.246 0.305 0.249 0.308 0.308 0.351 0.288 0.332 0.256 0.317 0.249 0.312 0.291 0.333 0.249 0.312 0.267 0.332
ETTh1 0.391 0.414 0.397 0.419 0.483 0.459 0.454 0.447 0.413 0.434 0.407 0.429 0.458 0.450 0.419 0.430 0.423 0.437
ETTh2 0.330 0.374 0.335 0.380 0.366 0.406 0.383 0.407 0.331 0.381 0.333 0.382 0.414 0.427 0.345 0.394 0.431 0.447
Traffic 0.390 0.264 0.418 0.278 0.413 0.284 0.428 0.282 0.391 0.264 0.410 0.282 0.62 0.336 0.356 0.261 0.434 0.295
Weather 0.217 0.251 0.224 0.252 0.277 0.306 0.246 0.298 0.226 0.264 0.218 0.260 0.259 0.287 0.236 0.282 0.246 0.300
Solar 0.188 0.215 0.192 0.216 0.251 0.262 0.233 0.262 0.207 0.294 0.223 0.264 0.301 0.319 0.223 0.288 0.230 0.295
Electricity 0.155 0.248 0.164 0.253 0.181 0.274 0.178 0.270 0.159 0.253 0.163 0.257 0.192 0.295 0.159 0.257 0.166 0.264
4.2 Ablation studies
Model architecture. To validate effectiveness of our model design, we perform ablation studies on TS-Register, prediction tasks, and reconstruction task in 10% few-shot setting. Table 2 shows the impact of each module. The TS-Register leverages multi-domain information during pre-training, aiding adaptive transfer to downstream datasets, as further discussed in section 4.4. The prediction tasks enhance performance in data-scarce situations. Without it, performance significantly drops on ETTh1 and ETTh2 with limited samples. Without the reconstruction task, our model shows negative transfer effects on ETTm1 and ETTm2, likely due to the prediction task making the model more susceptible to pre-training data biases.
Mask method.
To further validate the effectiveness of decomposed frequency learning, we replace the multi-frequency masking with different masking methods, including two mainstream time-domain methods: patch masking [6] and multi-patch masking [5], as well as a single frequency masking, which masks time series in the frequency domain only once. The results show that frequency domain modeling yields superior performance improvements compared to time-domain modeling. Additionally, we observe that applying multi-frequency masking results in enhanced outcomes, thus affirming that incorporating a multi-frequency perspective significantly aids the model in more accurately comprehending temporal patterns.
Table 2: Ablations on ROSE. We replace different masking methods, in addition to component removal. The average results of all predicted lengths are listed here.
Design ETTm1 ETTm2 ETTh1 ETTh2
MSE MAE MSE MAE MSE MAE MSE MAE
ROSE 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.380
Replace Multi-Freq Masking Single Freq Masking 0.354 0.379 0.257 0.313 0.403 0.423 0.347 0.387
Multi Patch Masking 0.356 0.379 0.259 0.316 0.404 0.426 0.349 0.389
Patch Masking 0.378 0.4 0.261 0.319 0.408 0.432 0.375 0.407
w/o TS-Register 0.354 0.378 0.256 0.312 0.418 0.427 0.355 0.39
Predict Task 0.36 0.384 0.257 0.314 0.422 0.438 0.372 0.41
Reconstruct Task 0.387 0.403 0.269 0.327 0.412 0.428 0.361 0.399
From Scratch 0.370 0.393 0.261 0.318 0.470 0.480 0.401 0.425
4.3 Scalability
Scalability is crucial for a general model, enabling significant performance improvements by expanding pre-training data and model sizes. To investigate the scalability of ROSE, we increased both the model size and dataset size and evaluated its predictive performance on four ETT datasets.
Figure 5: (a)/(b): Larger ROSE demonstrates better performance on downstream forecasting. (c)/(d): ROSE pre-trained on larger datasets demonstrates better performance on downstream forecasting.
Model size. Constrained by computational resources, we use 40% pre-training datasets. The results are shown in Figure 5(a) and (b). When maintaining the model dimension, we increased the model layers, increasing model parameters from 2M to 4.5M. This led to 10.37% and 9.34% improvements in the few-shot scenario with 5% and 10% downstream data, respectively.
Data size. When keeping the model size, we increase the size of the pre-training datasets from 178M to 886M. The results are shown in Figure 5(c) and (d). The performance of our model steadily improves with the increase in dataset size and achieves improvements of 7.4% and 4.8% respectively.
4.4 Model Analysis
Zero-shot and inference efficiency. By adopting the dual pre-train task, our model possesses strong generalization ability and can forecast without requiring any downstream data for fine-tuning.
Figure 6: Model performance in zero-shot setting comparison and inference efficiency comparison.
To validate ROSE’s zero-shot capabilities and assess its inference efficiency, we compared it with other models that have recently demonstrated similar abilities, namely Timer [2], MOIRAI [8], PatchTST [6], and GPT4TS [10]. Specific implementation details and results can be found in Appendix B. As shown in Figure 6, ROSE exhibits superior performance in zero-shot setting, with significantly fewer model parameters and inference time compared to other models. This underscores the robust generalization and scalability of ROSE.
Visualization of TS-Register. To validate the TS-Register’s capability to transfer domain-specific information adaptively from pre-training datasets to target datasets, we visualize the cosine similarity of register vector selections from datasets across different domains. As shown in Figure 7(a), the cosine similarity is higher for datasets within the same domain and lower between different domains. We also visualize the register vector selections from different datasets in Figures 7(b) and (c), where datasets from the same domain show similar visualizations. This confirms the TS-Register’s capability of adaptive transfer from multi-source to target datasets across various domains.
Figure 7: Visualization of TS-Register.
Sensitivity. The sensitivity analyses for the upper bound $a$ of the thresholds, the number of masked series $K_{\mathrm{f}}$, the number of register tokens $N_{\text{r}}$ and the size of codebook $H$ are presented in Appendix C.
5 Conclusion and future work
In this work, we propose ROSE, a novel general model, addressing the challenges of leveraging multi-domain datasets for enhancing downstream prediction task performance. ROSE utilizes decomposed frequency learning and TS-Register to capture generalized and domain-specific representations, enabling improved fine-tuning results, especially in data-scarce scenarios. Our experiments demonstrate ROSE’s superior performance over baselines with both full-data and few-data fine-tuning, as well as its impressive zero-shot capabilities. Future efforts will concentrate on expanding pre-training datasets and extending ROSE’s applicability across diverse time series analysis tasks, e.g., classification.
Appendix A IMPLEMENTATION DETAILS
A.1 Pre-training Datasets
We use multi-source datasets in pre-training which contain subsets of Monash [41], UEA [42] and UCR [43] time series regression datasets, as well as some other time series classical datasets [44, 45, 46, 47, 48]. The final list of all pre-training datasets is shown in Table 4. There is no overlap between the pre-training datasets and the target datasets. It is worth noting that the dataset weather in the pre-training dataset is a univariate dataset, which is different to the multivariate dataset weather in the target task. The pre-trained datasets can be categorized into 6 different domains according to their sources: Energy, Nature, Health, Transport, and Web. The sampling frequencies of the datasets show a remarkable diversity, ranging from millisecond samples to monthly samples, which reflects the diverse application scenarios and complexity of the real world. For all pre-training datasets, we split them into univariate sequences and train them in a channel-independent manner.
A.2 Evaluation Datasets
We use the following 8 multivariate time-series datasets for downstream fine-tuning and forecasting: ETT datasets contain 7 variates collected from two different electric transformers from July 2016 to July 2018. ETT datasets [^1] https://github.com/zhouhaoyi/ETDataset consist of four subsets, of which ETTh1/ETTh2 are recorded hourly and ETTm1/ETTm2 are recorded every 15 minutes. Traffic [^2] https://pems.dot.ca.gov/ contains road occupancy rates measured by 862 sensors on freeways in the San Francisco Bay Area from 2015 to 2016, recorded hourly. Weather [^3] https://www.bgc-jena.mpg.de/wetter/ collects 21 meteorological indicators, such as temperature and barometric pressure, for Germany in 2020, recorded every 10 minutes. Solar [^4] https://www.nrel.gov/grid/solar-power-data.html records solar power generation from 137 PV plants in 2006, every 10 minutes. Electricity [^5] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 contains the electricity consumption of 321 customers from July 2016 to July 2019, recorded hourly. We split each evaluation dataset into train-validation-test sets and detailed statistics of evaluation datasets are shown in Table 3.
Table 3: The statistics of evaluation datasets
Dataset ETTm1 ETTm2 ETTh1 ETTh2 Traffic Weather Solar Electricity
Variables 7 7 7 7 862 21 137 321
Timestamps 69680 69680 17420 17420 17544 52696 52560 26304
Split Ratio 6:2:2 6:2:2 6:2:2 6:2:2 7:1:2 7:1:2 7:1:2 7:1:2
Table 4: List of pretraining datasets.
Domain Dataset Frequency Time Pionts Source
Energy Aus. Electricity Demand Half Hourly 1155264 Monash[41]
Wind 4 Seconds 7397147 Monash[41]
Wind Farms Minutely 172178060 Monash[41]
London Smart Meters Half Hourly 166527216 Monash[41]
Nature Phoneme - 2160640 UCR[43]
EigenWorms - 27947136 UEA[42]
PRSA Hourly 4628448 [44]
Temperature Rain Daily 23252200 Monash[41]
StarLightCurves - 9457664 UCR[43]
Worms 0.033 Seconds 232200 UCR[43]
Saugeen River Flow Daily 23741 Monash[41]
Sunspot Daily 73924 Monash[41]
Weather Daily 43032000 Monash[41]
KDD Cup 2018 Daily 2942364 Monash[41]
US Births Daily 7305 Monash[41]
Health MotorImagery 0.001 Seconds 72576000 UEA[42]
SelfRegulationSCP1 0.004 Seconds 3015936 UEA[42]
SelfRegulationSCP2 0.004 Seconds 3064320 UEA[42]
AtrialFibrillation 0.008 Seconds 38400 UEA[42]
PigArtPressure - 624000 UCR[43]
PIGCVP - 624000 UCR[43]
TDbrain 0.002 Seconds 79232703 [45]
Transport Pems03 5 Minute 9382464 [46]
Pems04 5 Minute 5216544 [46]
Pems07 5 Minute 24921792 [46]
Pems08 5 Minute 3035520 [46]
Pems-bay 5 Minute 16937700 [46]
Pedestrian_Counts Hourly 3132346 Monash[41]
Web Web Traffic Daily 116485589 Monash[41]
Economic FRED_MD Monthly 77896 [47]
Bitcoin Daily 75364 Monash[41]
NN5 Daily 87801 [48]
A.3 Setting
We implemented ROSE in PyTorch [49] and all the experiments were conducted on 8 NVIDIA A800 80GB GPU. We used ADAM [50] with an initial learning rate of $5\times 10^{-4}$ and implemented learning rate decay using the StepLR method to implement learning rate decaying pre-training. By default, ROSE contains 3 encoder layers and 3 decoder layers with head number $H=16$ and the dimension of latent space $D=256$.
Pre-training. We use $N_{\mathrm{r}}=3$ as the number of register tokens and $P=8$ as the path tokens. We set the input length to 512 for the supervised prediction task with target lengths of 96, 192, 336, and 720. We also set the input length to 512 for self-supervised multi-frequency mask-reconstruction task, where mask number $K_{\mathbf{f}}=4$. The batch size is set to 8192 in pre-training.
Fine-tuning. We fix the lookback window to 512, and perform predictions with target lengths of 96, 192, 336, and 720, respectively. The number of register tokens $N_{\mathrm{r}}$ and patch tokens $P$ is the same as in pre-training, and the parameter $k=3$ in topk is set when selection vectors are performed in the register codebook. The batch size is set to 64 in fine-tuning.
A.4 Baselines
We select the state-of-the-art methods as our baselines to compare the performance of different models with full-data fine-tuning and 10% few-shot setting, including Transformer-based methods: Timer [2], iTransformer [20], and PatchTST [6]; CNN-based backbones method: TimesNet [18]; MLP-based methods: FITS [38], TIDE [39], and DLinear [40]. It is worth noting that Timer is a recent general time series model with a GPT-style architecture. Since the size of the look-back window can affect the performance of different models, we choose the look-back window size in $\{96,336,512,720\}$ for each baseline that achieves its best results for fair comparisons. The specific code base for these models is listed in Table 5:
Table 5: Code repositories for baselines
Baselines code repositories
Timer https://github.com/thuml/Large-Time-Series-Model
iTransformer https://github.com/thuml/iTransformer
PatchTST https://github.com/yuqinie98/PatchTST
FITS https://anonymous.4open.science/r/FITS/README.md
TimesNet https://github.com/thuml/TimesNet
TIDE https://github.com/google-research/google-research/tree/master/tide
Dlinear https://github.com/cure-lab/LTSF-Linear
Appendix B Zero shot and inference efficiency
To explore the inference efficiency and zero-shot capability of ROSE, we selected recent models with zero-shot capability, including Timer [2] and MOIRAI [8], as well as models with the potential for zero-shot learning, including PatchTST [6] and GPT4TS [10], for comparison on the ETT dataset. And we evaluate their inference efficiency by recording inference time per batch on the ETTh2 dataset with a batch size of 32.
Pre-train: We select the models with the largest parameter weights for Timer and MOIRAI as baselines. GPT4TS and PatchTST were pre-trained on our pre-training dataset with prediction tasks to acquire zero-shot capability. The specific settings are shown in the table 6.
Detailed results: For fare comparison, we set the input length as 96, 336, 512, and 720 for all zero-shot baseline models. We do not consider longer input length, as many real-world scenarios could offer very limited samples, and thus the zero-shot setting with only a few samples is crucial for them. As the baselines do not contain results with this setting in their original papers, we rerun them and report their best results in Table 11. ROSE consistently surpasses the baselines despite having significantly fewer parameters and inference time.
Table 6: Pre-train setting and inference efficiency
Model Parameters Pre-train data size Inference time per batch code repositories
ROSE 6.4M 0.89B 4.3 ms -
MOIRAI_large 311M 27B 90.3ms https://github.com/redoules/moirai
GPT4TS 81.9M 0.89B 45.6 ms https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All
Timer 67.4M 1B 68 ms https://github.com/thuml/Large-Time-Series-Model
PatchTST 5.9M 0.89B 4.2 ms https://github.com/PatchTST/PatchTST
Appendix C Sensitivity
Number of masked series. As described in Section 3.2, we propose decomposed frequency learning, which employs multiple thresholds to randomly mask high and low frequencies in the frequency domain, thereby decomposing the original time series into multiple frequency components. This allows the model to understand the time series from multiple frequency perspectives. In this experiment, we study the influence of the number of masked series $K_{m}athrm{f}$ on downstream performance. We train ROSE with 1, 2, 3, 4, 5, or 6 mask series. We report the results of this analysis in Figure 8(a). We find that as the number of masked sequences increases, the downstream performance gradually improves. This is because the model can better understand the time series from the decomposed frequency components, which enhances the model’s generalization ability. However, more masked series do not bring better downstream performance. This could be due to an excessive number of masked sequences leading to information redundancy. In all our experiments, we kept 4 mask series.
Number of register tokens. The TS-register module presented in Section 3.3 supports the configuration of an arbitrary number of register tokens. In Figure 8(b), we visualize the relationship between the performance on the ETT datasets under a 10% few-shot setting and the number of register tokens. It is observed that when the number of register tokens ranges from 1 to 6, the model’s performance remains relatively stable, with an optimal outcome achieved when the number is set to 3. This phenomenon may be because when the number of register tokens is too small, they contain insufficient domain-specific information, which limits their effectiveness in enhancing the model’s performance. Conversely, an excess of register tokens may introduce redundant information, hindering the accurate representation of domain-specific information. Additionally, we compared the results without the adjustment of a low-rank matrix on the register tokens and found that the incorporation of a low-rank matrix adjustment led to improvements across all quantities of register tokens. This finding underscores the significance of utilizing a low-rank matrix to supplement the register tokens with downstream data-specific information.
Figure 8: (a): Analysis of the number of masked series. (b): Analysis of the number of register tokens
Thresholds upper bound. Figure 9(a) illustrates the relationship between threshold upper bound and model performance. We have observed that the upper bound of the threshold has a minimal impact on the model’s performance. Generally, the information density is higher in low-frequency components compared to high-frequency ones. Therefore, the upper bound of the threshold should be biased towards the low-frequency range to balance the information content between low-frequency and high-frequency components. However, this bias should not be excessive. Our experiments indicate that an upper bound of $L/10$ performs worse than $L/5$ as an overly left-skewed threshold results in insufficient information in the low-frequency range, making the reconstruction task either too difficult or too simple. Based on our findings, we recommend using $L/5$ as the upper bound for the threshold.
Register codebook size. Figure 9(b) illustrates the relationship between register codebook size and model performance. The codebook size determines the upper limit of domain-specific information that the register can store. We can observe that there is a significant improvement in the model effect when the codebook size is increased from 32 to 128. When the codebook size exceeds 128, the improvement of the model effect with the increase of codebook size is no longer obvious. Therefore, we believe that 128 is an appropriate codebook size for the current pre-training datasets.
Figure 9: (a): Analysis of the threshold upper bound. (b): Analysis of the size of codebook
Appendix D Model generality
We evaluate the effectiveness of our proposed Multiple Frequency Mask (MFM) on Transformer-based models and CNN-based models, whose results are shown in Table 7. It is notable that MFM consistently improves these forecasting models. Specifically, it achieves average improvements of 6.3%, 3.7%, 1.5% in Autoformer [19], TimesNet [18], and PatchTST [6], respectively. This indicates that MFM can be widely utilized across various time series forecasting models to learn generalized time series representations and improve prediction accuracy.
Table 7: Performance Multiple Frequency Masking.
Datasets ETTm1 ETTm2 ETTh1 ETTh2
Metric MSE MAE MSE MAE MSE MAE MSE MAE
Autoformer +Multi Freq Mask 0.600 0.521 0.328 0.365 0.493 0.487 0.452 0.458
0.549 0.488 0.306 0.349 0.474 0.478 0.406 0.425
TimesNet +Multi Freq Mask 0.4 0.406 0.291 0.333 0.458 0.45 0.414 0.427
0.386 0.398 0.282 0.324 0.446 0.438 0.386 0.403
PatchTST +Multi Freq Mask 0.353 0.382 0.256 0.317 0.413 0.434 0.331 0.381
0.347 0.372 0.252 0.308 0.405 0.424 0.337 0.379
Appendix E Visualization of forecasting results
To provide a distinct comparison among different models, we present visualizations of the forecasting results on the ETTh2 dataset and the weather dataset in different settings, as shown in Figures 10 to Figures 13, given by the following models: DLinear [40], TimesNet [18], iTransfomrer [20], PatchTST [6], and TiDE [39]. Among the methods, ROSE demonstrates the most accurate prediction ability.
Figure 10: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset with full-sample.
Figure 11: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset with 10% samples.
Figure 12: Visualization of input-512 and predict-336 forecasting results on the weather dataset with full-sample.
Figure 13: Visualization of input-512 and predict-336 forecasting results on the weather dataset with 10%-sample.
Appendix F Limitations
Our pre-training datasets still have room for expansion. While ROSE demonstrates strong performance, its effectiveness could be enhanced with a larger and more diverse set of pre-training datasets. By collecting additional real-world scenario data for pre-training, we aim to further improve the robustness and generalizability of ROSE across various domains.
Appendix G Broader Impacts
This paper presents ROSE, a novel general time series forecasting model that demonstrates promising performance. ROSE performs well in few-shot scenarios and shows notable zero-shot capabilities, highlighting its scalability and efficiency. We believe it can be a valuable addition to the pre-training research community. To support future research, we will also release the codebase for time-series pre-training.
This paper only focuses on the algorithm design. Using all the codes and datasets strictly follows the corresponding licenses A.1. There is no potential ethical risk or negative social impact.
Appendix H Full Results
H.1 Full-shot results
Table 8: Full results in full-shot setting.
Models ROSE Timer ITransformer PatchTST FITS Timesnet TiDE Dlinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.275 0.328 0.339 0.383 0.334 0.368 0.293 0.346 0.303 0.352 0.338 0.375 0.306 0.349 0.299 0.343
192 0.324 0.358 0.396 0.430 0.377 0.391 0.333 0.370 0.337 0.368 0.374 0.387 0.335 0.366 0.335 0.365
336 0.354 0.377 0.451 0.471 0.426 0.420 0.369 0.392 0.366 0.385 0.410 0.411 0.364 0.384 0.369 0.386
720 0.411 0.407 0.558 0.540 0.491 0.459 0.416 0.420 0.415 0.411 0.478 0.450 0.413 0.413 0.425 0.421
avg 0.341 0.367 0.436 0.456 0.407 0.410 0.353 0.382 0.355 0.379 0.400 0.406 0.355 0.378 0.357 0.379
ETTm2 96 0.157 0.243 0.184 0.263 0.180 0.264 0.166 0.256 0.162 0.253 0.187 0.267 0.161 0.251 0.167 0.260
192 0.213 0.283 0.278 0.334 0.250 0.309 0.223 0.296 0.216 0.291 0.249 0.309 0.215 0.289 0.224 0.303
336 0.266 0.319 0.334 0.375 0.311 0.348 0.274 0.329 0.268 0.326 0.321 0.351 0.267 0.326 0.281 0.342
720 0.347 0.373 0.435 0.431 0.412 0.407 0.362 0.385 0.348 0.378 0.408 0.403 0.352 0.383 0.397 0.421
avg 0.246 0.305 0.308 0.351 0.288 0.332 0.256 0.317 0.249 0.312 0.291 0.333 0.249 0.312 0.267 0.332
ETTh1 96 0.354 0.385 0.358 0.345 0.386 0.405 0.370 0.400 0.372 0.402 0.384 0.402 0.375 0.398 0.375 0.399
192 0.389 0.407 0.438 0.448 0.441 0.436 0.413 0.429 0.404 0.422 0.436 0.429 0.412 0.422 0.405 0.416
336 0.406 0.422 0.491 0.484 0.487 0.458 0.422 0.440 0.427 0.438 0.491 0.469 0.435 0.433 0.439 0.443
720 0.413 0.443 0.644 0.559 0.503 0.491 0.447 0.468 0.424 0.454 0.521 0.500 0.454 0.465 0.472 0.490
avg 0.391 0.414 0.483 0.459 0.454 0.447 0.413 0.434 0.407 0.429 0.458 0.450 0.419 0.430 0.423 0.437
ETTh2 96 0.265 0.320 0.288 0.346 0.297 0.349 0.274 0.337 0.271 0.336 0.340 0.374 0.270 0.336 0.289 0.353
192 0.328 0.369 0.380 0.409 0.380 0.400 0.341 0.382 0.331 0.374 0.402 0.414 0.332 0.380 0.383 0.418
336 0.353 0.391 0.372 0.413 0.428 0.432 0.329 0.384 0.354 0.395 0.452 0.452 0.360 0.407 0.448 0.465
720 0.376 0.417 0.425 0.456 0.427 0.445 0.379 0.422 0.377 0.423 0.462 0.468 0.419 0.451 0.605 0.551
avg 0.330 0.374 0.366 0.406 0.383 0.407 0.331 0.381 0.333 0.382 0.414 0.427 0.345 0.394 0.431 0.447
Traffic 96 0.354 0.252 0.351 0.255 0.395 0.268 0.360 0.249 0.385 0.271 0.593 0.321 0.336 0.253 0.410 0.282
192 0.377 0.257 0.387 0.269 0.417 0.276 0.379 0.256 0.397 0.275 0.617 0.336 0.346 0.257 0.423 0.287
336 0.396 0.262 0.426 0.288 0.433 0.283 0.392 0.264 0.410 0.280 0.629 0.336 0.355 0.260 0.436 0.296
720 0.434 0.283 0.489 0.323 0.467 0.302 0.432 0.286 0.448 0.301 0.640 0.350 0.386 0.273 0.466 0.315
avg 0.390 0.264 0.413 0.284 0.428 0.282 0.391 0.264 0.410 0.282 0.620 0.336 0.356 0.261 0.434 0.295
Weather 96 0.145 0.182 0.154 0.208 0.174 0.214 0.149 0.198 0.143 0.195 0.172 0.220 0.166 0.222 0.176 0.237
192 0.183 0.226 0.243 0.287 0.221 0.254 0.194 0.241 0.186 0.238 0.219 0.261 0.209 0.263 0.220 0.282
336 0.232 0.267 0.329 0.345 0.278 0.296 0.245 0.282 0.236 0.277 0.280 0.306 0.254 0.301 0.265 0.319
720 0.309 0.327 0.384 0.384 0.358 0.347 0.314 0.334 0.307 0.330 0.365 0.359 0.313 0.340 0.323 0.362
avg 0.217 0.251 0.277 0.306 0.258 0.278 0.226 0.264 0.218 0.260 0.259 0.287 0.236 0.282 0.246 0.300
Solar 96 0.171 0.199 0.230 0.245 0.203 0.237 0.190 0.273 0.198 0.253 0.250 0.292 0.200 0.270 0.206 0.281
192 0.185 0.215 0.249 0.261 0.233 0.261 0.204 0.302 0.218 0.262 0.296 0.318 0.219 0.285 0.225 0.291
336 0.193 0.218 0.255 0.265 0.248 0.273 0.212 0.293 0.235 0.270 0.319 0.330 0.234 0.296 0.240 0.300
720 0.198 0.224 0.272 0.276 0.249 0.275 0.221 0.310 0.241 0.273 0.338 0.337 0.242 0.303 0.248 0.307
avg 0.187 0.214 0.251 0.262 0.233 0.262 0.207 0.294 0.223 0.264 0.301 0.319 0.223 0.288 0.230 0.295
Electricity 96 0.125 0.220 0.136 0.230 0.148 0.240 0.129 0.222 0.134 0.231 0.168 0.272 0.132 0.229 0.140 0.237
192 0.142 0.235 0.168 0.260 0.162 0.253 0.147 0.240 0.149 0.244 0.184 0.289 0.147 0.243 0.153 0.249
336 0.162 0.252 0.192 0.285 0.178 0.269 0.163 0.259 0.165 0.260 0.198 0.300 0.161 0.261 0.169 0.267
720 0.191 0.284 0.230 0.322 0.225 0.317 0.197 0.290 0.203 0.293 0.220 0.320 0.196 0.294 0.203 0.301
avg 0.155 0.248 0.181 0.274 0.178 0.270 0.159 0.253 0.163 0.257 0.193 0.295 0.159 0.257 0.166 0.264
H.2 Few-shot results
Table 9: Full results in 10% few-shot setting
Models ROSE Timer ITransformer PatchTST FITS Timesnet TiDE Dlinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.287 0.336 0.351 0.398 0.353 0.392 0.317 0.363 0.311 0.354 0.481 0.446 0.373 0.388 0.454 0.475
192 0.331 0.362 0.429 0.448 0.385 0.410 0.351 0.382 0.351 0.378 0.621 0.491 0.406 0.403 0.575 0.548
336 0.362 0.379 0.496 0.495 0.422 0.432 0.376 0.398 0.377 0.396 0.521 0.479 0.436 0.425 0.773 0.631
720 0.416 0.412 0.573 0.573 0.494 0.472 0.435 0.430 0.441 0.436 0.571 0.508 0.499 0.457 0.943 0.716
avg 0.349 0.372 0.462 0.478 0.414 0.426 0.370 0.393 0.370 0.391 0.549 0.481 0.429 0.418 0.686 0.593
ETTm2 96 0.159 0.247 0.196 0.273 0.183 0.279 0.170 0.259 0.166 0.258 0.212 0.292 0.195 0.280 0.493 0.476
192 0.217 0.287 0.290 0.338 0.247 0.320 0.226 0.297 0.222 0.295 0.297 0.353 0.257 0.317 0.923 0.658
336 0.268 0.322 0.350 0.386 0.3 0.353 0.284 0.333 0.273 0.328 0.328 0.364 0.320 0.354 1.407 0.822
720 0.355 0.377 0.436 0.435 0.385 0.408 0.363 0.382 0.370 0.385 0.456 0.440 0.421 0.410 1.626 0.905
avg 0.249 0.308 0.318 0.358 0.279 0.340 0.261 0.318 0.258 0.317 0.323 0.362 0.298 0.340 1.112 0.715
ETTh1 96 0.367 0.395 0.361 0.409 0.442 0.464 0.458 0.463 0.438 0.449 0.579 0.522 0.708 0.569 1.355 0.816
192 0.399 0.416 0.398 0.427 0.476 0.475 0.481 0.490 0.512 0.489 0.641 0.553 0.744 0.584 1.210 0.825
336 0.405 0.423 0.405 0.436 0.486 0.482 0.465 0.475 0.584 0.532 0.721 0.582 0.768 0.599 1.487 0.914
720 0.416 0.443 0.492 0.503 0.509 0.506 0.478 0.492 0.681 0.590 0.630 0.574 0.770 0.606 1.369 0.826
avg 0.397 0.419 0.414 0.444 0.478 0.482 0.470 0.480 0.554 0.515 0.643 0.558 0.748 0.589 1.355 0.845
ETTh2 96 0.273 0.332 0.298 0.356 0.333 0.385 0.350 0.389 0.297 0.360 0.378 0.413 0.434 0.432 1.628 0.724
192 0.334 0.376 0.374 0.404 0.402 0.428 0.416 0.426 0.363 0.399 0.463 0.460 0.528 0.480 1.388 0.713
336 0.358 0.397 0.385 0.425 0.438 0.452 0.401 0.429 0.388 0.424 0.507 0.495 0.473 0.457 1.595 0.772
720 0.376 0.417 0.468 0.487 0.466 0.477 0.436 0.457 0.423 0.456 0.516 0.501 0.567 0.513 1.664 0.857
avg 0.335 0.380 0.381 0.418 0.410 0.436 0.401 0.425 0.368 0.410 0.466 0.467 0.501 0.471 1.569 0.766
Traffic 96 0.398 0.270 0.353 0.277 0.458 0.314 0.421 0.299 0.406 0.301 0.705 0.386 0.681 0.412 0.616 0.385
192 0.405 0.270 0.405 0.288 0.473 0.319 0.439 0.313 0.523 0.424 0.710 0.393 0.624 0.385 0.710 0.480
336 0.417 0.277 0.428 0.300 0.491 0.329 0.448 0.318 0.624 0.479 0.863 0.456 0.613 0.368 0.723 0.481
720 0.452 0.294 0.487 0.332 0.536 0.361 0.478 0.320 0.937 0.629 0.928 0.485 0.653 0.387 0.673 0.436
avg 0.418 0.278 0.418 0.299 0.490 0.331 0.447 0.312 0.623 0.458 0.801 0.430 0.643 0.388 0.680 0.446
Weather 96 0.145 0.184 0.180 0.228 0.189 0.229 0.166 0.217 0.174 0.222 0.199 0.248 0.208 0.240 0.230 0.318
192 0.190 0.227 0.256 0.296 0.239 0.269 0.211 0.257 0.226 0.265 0.249 0.285 0.249 0.274 0.357 0.425
336 0.245 0.269 0.321 0.344 0.294 0.308 0.261 0.296 0.279 0.301 0.297 0.316 0.283 0.306 0.464 0.493
720 0.317 0.328 0.420 0.398 0.366 0.356 0.328 0.342 0.361 0.354 0.367 0.361 0.365 0.354 0.515 0.532
avg 0.224 0.252 0.294 0.317 0.272 0.291 0.242 0.278 0.260 0.285 0.278 0.303 0.276 0.293 0.391 0.442
Solar 96 0.177 0.203 0.230 0.284 0.292 0.310 0.202 0.245 0.242 0.306 0.306 0.323 0.359 0.382 0.748 0.757
192 0.185 0.214 0.246 0.300 0.313 0.321 0.222 0.256 0.262 0.315 0.356 0.354 0.395 0.400 0.752 0.759
336 0.193 0.218 0.278 0.319 0.332 0.335 0.252 0.272 0.276 0.321 0.387 0.362 0.429 0.410 0.754 0.760
720 0.208 0.225 0.255 0.300 0.355 0.338 0.264 0.273 0.280 0.329 0.396 0.361 0.410 0.378 0.755 0.761
avg 0.191 0.215 0.252 0.301 0.323 0.326 0.235 0.262 0.265 0.318 0.361 0.350 0.398 0.392 0.752 0.759
Electricity 96 0.135 0.226 0.156 0.254 0.184 0.276 0.161 0.256 0.144 0.242 0.279 0.359 0.271 0.359 0.227 0.334
192 0.150 0.240 0.254 0.275 0.192 0.284 0.163 0.257 0.174 0.283 0.282 0.363 0.216 0.308 0.265 0.366
336 0.166 0.258 0.212 0.306 0.216 0.308 0.173 0.266 0.225 0.341 0.289 0.367 0.237 0.324 0.339 0.417
720 0.205 0.290 0.271 0.352 0.265 0.347 0.221 0.313 0.408 0.495 0.333 0.399 0.258 0.330 0.482 0.478
avg 0.164 0.253 0.223 0.297 0.214 0.304 0.180 0.273 0.238 0.341 0.296 0.372 0.245 0.330 0.328 0.399
H.3 Ablation study results
Table 10: Full results of ablation study
Design Pred_len ETTm1 ETTm2 ETTh1 ETTh2
MSE MAE MSE MAE MSE MAE MSE MAE
ROSE 96 0.287 0.336 0.159 0.247 0.367 0.395 0.273 0.332
192 0.331 0.362 0.217 0.287 0.399 0.416 0.334 0.376
336 0.362 0.379 0.269 0.322 0.405 0.423 0.358 0.397
720 0.416 0.412 0.357 0.377 0.416 0.443 0.376 0.417
avg 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.38
Replace Multi Freq Masking Single Freq Masking 96 0.296 0.346 0.163 0.253 0.371 0.398 0.272 0.335
192 0.336 0.368 0.221 0.291 0.406 0.420 0.340 0.376
336 0.362 0.385 0.276 0.327 0.411 0.426 0.379 0.406
720 0.424 0.418 0.366 0.381 0.422 0.447 0.399 0.431
avg 0.354 0.379 0.257 0.313 0.403 0.423 0.347 0.387
Multi Patch Masking 96 0.302 0.348 0.168 0.257 0.377 0.408 0.282 0.343
192 0.336 0.367 0.228 0.297 0.404 0.423 0.343 0.379
336 0.364 0.385 0.277 0.328 0.405 0.420 0.374 0.403
720 0.423 0.416 0.364 0.381 0.431 0.455 0.396 0.430
avg 0.356 0.379 0.259 0.316 0.404 0.426 0.349 0.389
Patch Masking 96 0.318 0.366 0.168 0.259 0.388 0.412 0.303 0.359
192 0.355 0.388 0.228 0.298 0.402 0.422 0.370 0.399
336 0.388 0.406 0.279 0.331 0.411 0.435 0.413 0.428
720 0.450 0.438 0.370 0.388 0.431 0.459 0.413 0.443
avg 0.378 0.4 0.261 0.319 0.408 0.432 0.375 0.407
w/o Aqp 96 0.297 0.345 0.164 0.252 0.379 0.399 0.276 0.336
192 0.334 0.367 0.221 0.290 0.419 0.420 0.350 0.380
336 0.360 0.384 0.275 0.325 0.438 0.442 0.393 0.411
720 0.424 0.416 0.364 0.379 0.435 0.448 0.400 0.432
avg 0.354 0.378 0.256 0.312 0.418 0.427 0.355 0.39
Predict Task 96 0.301 0.348 0.166 0.255 0.380 0.407 0.295 0.359
192 0.343 0.374 0.221 0.291 0.410 0.426 0.372 0.406
336 0.374 0.393 0.275 0.327 0.440 0.443 0.403 0.429
720 0.424 0.420 0.366 0.384 0.458 0.476 0.418 0.446
avg 0.36 0.384 0.257 0.314 0.422 0.438 0.372 0.41
Reconstruct Task 96 0.329 0.371 0.175 0.265 0.374 0.399 0.296 0.355
192 0.363 0.391 0.233 0.304 0.407 0.422 0.354 0.389
336 0.394 0.407 0.287 0.340 0.437 0.440 0.385 0.413
720 0.461 0.442 0.379 0.396 0.430 0.453 0.408 0.438
avg 0.387 0.403 0.269 0.327 0.412 0.428 0.361 0.399
H.4 Zero-shot results
Table 11: Full results in zero-shot setting
Model ROSE MOIRAI Timer PatchTST GPT4TS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.512 0.460 0.660 0.476 0.596 0.519 0.906 0.578 0.743 0.545
192 0.512 0.462 0.707 0.500 0.668 0.560 0.892 0.573 0.759 0.563
336 0.523 0.470 0.730 0.515 1.580 0.871 0.876 0.567 0.748 0.570
720 0.552 0.490 0.758 0.536 1.587 0.882 0.836 0.554 0.780 0.576
avg 0.525 0.471 0.714 0.507 1.108 0.708 0.877 0.568 0.758 0.564
ETTm2 96 0.224 0.309 0.216 0.282 0.221 0.300 0.301 0.346 0.275 0.343
192 0.266 0.333 0.294 0.330 0.303 0.349 0.337 0.367 0.298 0.356
336 0.310 0.358 0.368 0.373 0.411 0.423 0.376 0.389 0.347 0.386
720 0.395 0.407 0.494 0.439 0.505 0.473 0.456 0.434 0.426 0.417
avg 0.299 0.352 0.343 0.356 0.360 0.386 0.367 0.384 0.337 0.376
ETTh1 96 0.382 0.408 0.405 0.397 0.414 0.439 0.583 0.502 0.573 0.521
192 0.400 0.420 0.458 0.428 0.458 0.465 0.600 0.516 0.594 0.529
336 0.404 0.426 0.509 0.454 0.475 0.479 0.594 0.525 0.682 0.567
720 0.420 0.447 0.529 0.494 0.521 0.511 0.569 0.544 0.543 0.524
avg 0.401 0.425 0.475 0.443 0.467 0.474 0.586 0.522 0.598 0.535
ETTh2 96 0.298 0.362 0.303 0.338 0.313 0.366 0.401 0.418 0.369 0.411
192 0.336 0.385 0.369 0.384 0.394 0.415 0.412 0.425 0.382 0.423
336 0.353 0.399 0.397 0.410 0.420 0.439 0.410 0.427 0.401 0.442
720 0.395 0.432 0.447 0.450 0.440 0.459 0.415 0.437 0.424 0.439
avg 0.346 0.394 0.379 0.396 0.392 0.420 0.409 0.427 0.394 0.429

### v2

@@
-ROSE: Register Assisted General Time Series Forecasting with Decomposed Frequency Learning
+ROSE: Register-assisted General Time Series Forecasting with Decomposed Frequency Learning
 Abstract
@@
 To address these challenges, we propose a Register Assisted General Time Series Forecasting Model with Decomposed Frequency Learning (ROSE), a novel pre-trained model for time series forecasting.
-ROSE employs Decomposed Frequency Learning for the pre-training task, which decomposes coupled semantic and periodic information in time series with frequency-based masking and reconstruction to obtain unified representations across domains.
-We also equip ROSE with a Time Series Register, which learns to generate a register codebook to capture domain-specific representations during pre-training and enhances domain-adaptive transfer by selecting related register tokens on downstream tasks.
-After pre-training on large-scale time series data, ROSE achieves state-of-the-art forecasting performance on 8 real-world benchmarks.
-Remarkably, even in few-shot scenarios, it demonstrates competitive or superior performance compared to existing methods trained with full data.
+ROSE employs Decomposed Frequency Learning for the pre-training task, which decomposes coupled semantic information in time series with frequency-based masking and reconstruction to obtain unified representations across domains.
+We also equip ROSE with a Time Series Register, which learns to generate a register to capture domain-specific representations during pre-training and enhances domain-adaptive transfer by selecting related register tokens on downstream tasks.
+After pre-training on large-scale time series data, ROSE achieves state-of-the-art forecasting performance on 7 real-world benchmarks.
+Remarkably, it demonstrates competitive or superior few-shot and zero-shot abilities.
 1 Introduction
-Time Series Forecasting plays a pivotal role across numerous domains, such as energy, smart transportation, weather, and economics [1].
-However, training specific models in deep learning for each dataset is costly and requires tailored parameter tuning, whose prediction accuracy may be limited due to data scarcity [2].
-One solution is to pre-train a general model on diverse time series datasets and fine-tune it with a few data for different downstream scenarios.
-However, applying this paradigm to heterogeneous time series data presents challenges.
+Time Series Forecasting plays a pivotal role across numerous domains, such as energy, smart transportation, weather, and economics (Qiu et al., 2024).
+However, training specific models in deep learning for each dataset is costly and requires tailored parameter tuning, whose prediction accuracy may be limited due to data scarcity (Liu et al., 2024).
+One solution is to pre-train a general model on diverse time series datasets and fine-tune it with a few data for different downstream scenarios or direct predict without fine-tuning.
+Following this idea, foundation models for time series forecasting have recently raised growing attention, continually scaling up the pre-training datasets and model sizes to improve the generalization performance (Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024).
+However, an excessively large scale also causes increasing costs in training and inference, which may go against the original intention of a general model, especially in resource-constrained situations.
+In addition to scaling up, general time series forecasting models can also be designed from the perspectives of pre-training tasks and downstream transfer adaptation.
+From these two angles, we identify the following challenges.
 Obtaining a unified representation from time series data across various domains is challenging.
-Time series from each domain involve complex temporal patterns, composed of multiple frequency components combined with each other [3], which is frequency superposition.
+Time series from each domain involve complex temporal patterns, composed of multiple frequency components combined with each other (Zhou et al., 2022), which is frequency superposition.
 As shown in Figure 1(a), different frequency components contain distinct semantic information.
-For example, low and high-frequency components represent long-term trends and rapid variations, respectively [4].
-Additionally, different frequency components also contain cross-period information, revealing the interactions between short and long-term patterns in the time series.
-Thus, the frequency superposition leads to coupled semantic and periodic information in the time domain.
-Furthermore, large-scale time series data from different domains introduce even more complex temporal patterns and frequency diversity.
-Existing pre-training frameworks [5, 6, 7], such as mask modeling and contrastive learning, were proposed to learn a unified representation from time series data.
+For example, low and high-frequency components represent long-term trends and rapid variations, respectively (Zhang et al., 2022).
+Furthermore, different datasets exhibit diverse frequency distributions, and the significance of low-frequency and high-frequency components for time series modeling varies across domains(Zhang et al., 2024).
+As a result, large-scale time series data from different domains introduce even more complex temporal patterns and frequency diversity.
+Existing pre-training frameworks (Dong et al., 2024; Nie et al., 2022; Lee et al., 2023), such as masked modeling and contrastive learning, were proposed to learn a unified representation from time domain.
 However, these methods overlook the frequency diversity and complexity exhibited in heterogeneous time series that come from various domains, making it difficult to capture intricate patterns, thus limiting their generalization capabilities.
 Adaptive transferring information from multi-domain time series to specific downstream scenarios presents a challenge.
-Multi-source time series data originate from various domains [8], whose data exhibit domain-specific information [2].
-Information from the same or similar domain as the target domain is useful for improving the model’s effectiveness in the target task [9].
-However, as shown in Figure 1(a), existing time series pre-training frameworks [8, 2, 10] focus mainly on capturing generalized features during pre-training and overlook domain-specific features.
-Thus, they only transfer the same generalized representation to different target domains, called direct transfer, which limits the model’s effectiveness in specific downstream tasks, as shown in Figure 1(b).
+Multi-source time series data originate from various domains (Woo et al., 2024), whose data exhibit domain-specific information (Liu et al., 2024).
+Information from the same or similar domain as the target domain is useful for improving the model’s effectiveness in the target task (Chen et al., 2023a).
+However, as shown in Figure 1(a), existing time series pre-training frameworks (Woo et al., 2024; Liu et al., 2024; Zhou et al., 2024) focus mainly on learning generalized representation during pre-training and overlook domain-specific representation.
+Thus, they only transfer the same generalized representation to different target domains, called direct transfer, which limits the model’s effectiveness in specific downstream tasks.
 Therefore, it is necessary to learn domain-specific information during pre-training and adaptively transfer the specific representations to each target domain, called adaptive transfer.
@@
 To address these challenges, we propose a register assisted general time series forecasting model with decomposed frequency learning (ROSE).
-First, we propose Decomposed Frequency Learning that learns generalized representations to solve the issue with coupled semantic and periodic information.
-We decompose individual time series using the Fourier transform with a novel frequency-based mask method, and then convert it back to the time domain to obtain decoupled time series for reconstruction.
+First, we propose Decomposed Frequency Learning that learns generalized representations to solve the issue with coupled semantic information.
+We decompose individual time series using the Fourier transform with a novel frequency-based masking method, and then convert it back to the time domain to obtain decoupled time series for reconstruction.
 It makes complex temporal patterns disentangled, thus benefiting the model to learn generalized representations.
 Second, we introduce Time Series Register (TS-Register) to learn domain-specific information in multi-domain data.
-By setting up a register codebook, we generate register tokens to learn each domain-specific information during pre-training.
-In a downstream scenario, the model adaptively selects vectors from the register codebook that are close to the target domain of interest.
-During fine-tuning, we incorporate learnable vectors into the selected register tokens to complement target specific information to perform more flexible adaptive transfer.
+By setting up a register, we generate register tokens to learn each domain-specific information during pre-training.
+In a downstream scenario, the model adaptively selects Top-K vectors from the register that are close to the target domain of interest.
+During fine-tuning, we adjust the selected register tokens with a novel learnable low-rank matrix, which complements target-specific information to perform more flexible adaptive transfer.
 As shown in Figure 1(b), adaptive transfer successfully utilizes domain-specific information in multi-domain time series, which contributes to the model’s performance in target tasks.
 The contributions are summarized as follows:
-  • We propose ROSE, a novel general time series forecasting model using multi-domain datasets for pre-training and improving downstream fine-tuning performance and efficiency.
-  • We propose a novel Decomposed Frequency Learning that employs multi-frequency masking to decouple complex temporal patterns, empowering the model’s generalization capability.
-  • We propose TS-Register to capture domain-specific information in pre-training and enable adaptive transfer of target-oriented specific information for downstream tasks in fine-tuning.
-  • Our experiments with 8 real-world benchmarks demonstrate that ROSE achieves state-of-the-art performance in full-shot setting and achieves competitive or superior results in few-shot setting, along with impressive transferability in zero-shot setting.
+  • We propose ROSE, a novel light weight general time series forecasting model using multi-domain datasets for pre-training and improving downstream fine-tuning performance and efficiency.
+  • We propose a novel Decomposed Frequency Learning that employs multi-frequency masking to learn complex general temporal patterns from multi-domain data, empowering the model’s generalization capability.
+  • We propose a novel TS-Register to capture domain-specific information in pre-training and enable adaptive transfer of target-oriented specific information for downstream tasks.
+  • Our experiments with 7 real-world benchmarks demonstrate that ROSE achieves state-of-the-art performance in full-shot setting and achieves competitive or superior results in few-shot setting, along with impressive transferability in zero-shot setting.
 2 Related work
-2.1 Time series forecasting
-Traditional time series prediction models like ARIMA [11], despite their theoretical support, are limited in modeling nonlinearity.
-With the rise of deep learning, many RNN-based models [12, 13, 14] have been proposed, modeling the sequential data with an autoregressive process.
-CNN-based models [15, 16] have also received widespread attention due to their ability to capture local features.
-MICN [17] utilizes TCN to capture both local and global features, while TimesNet [18] focuses on modeling 2D temporal variations.
+2.1 Traditional time series forecasting
+The statistical time series forecasting models like ARIMA (Box and Jenkins, 1968), despite their theoretical support, are limited in modeling nonlinearity.
+With the rise of deep learning, many RNN-based models (Cirstea et al., 2019; Wen et al., 2017; Salinas et al., 2020) have been proposed, modeling the sequential data with an autoregressive process.
+CNN-based models (Luo and Wang, 2024; Liu et al., 2022a) have also received widespread attention due to their ability to capture local features.
+MICN (Wang et al., 2022) utilizes TCN to capture both local and global features, while TimesNet (Wu et al., 2022) focuses on modeling 2D temporal variations.
 However, both RNNs and CNNs struggle to capture long-term dependencies.
-Transformer-based models [3, 6, 19, 20, 21], with their attention mechanism, can capture long dependencies and extract global information, leading to widespread applications in long-time series prediction.
+Transformer-based models (Zhou et al., 2022; Nie et al., 2022; Wu et al., 2021; Liu et al., 2023; Chen et al., 2024), with their attention mechanism, can capture long dependencies and extract global information, leading to widespread applications in long-time series prediction.
 However, this case-by-case paradigm requires meticulous hyperparameter design for different datasets, and its predictive performance can also be affected by data scarcity.
-2.2 Time series pre-training
-Masked modeling [5, 22, 23] and contrastive learning [24, 25, 26] have attracted significant attention in the time series domain.
-PatchTST [6] proposes to capture local semantic information by predicting masked subsequence patches.
-LaST [27] explores using variational inference to decompose time series into representations of seasonality and periodicity within a latent space.
-PITS [7] combines masked modeling and contrastive learning to obtain better generalization representations.
-However, the above methods are limited to one-to-one pre-training and fine-tuning, overlooking the frequency diversity and complexity within multi-domain data, thus making them challenging to directly apply to pre-training for multi-domain time series.
-Multi-source time series pre-training has recently garnered widespread attention [28, 29, 30].
-MOMENT [31] and MOIRAI [8] adopt a BERT-style pre-training approach, while Timer [2] and PreDcT [32] use a GPT-style pre-training approach, giving rise to improved performance in time series prediction.
+2.2 Pre-trained models for times seires forecasting
+LLMs for time series forecasting: Recent studies have shown that Large Language Models (LLMs) can enhance time series forecasting with limited fine-tuning (Zhou et al., 2024), prompting (Jin et al., 2024; Cao et al., 2024), and modality alignment (Jin et al., 2024).
+GPT4TS (Zhou et al., 2024) fine-tunes a subset of LLM parameters, and achieves competitive performance by leveraging pre-trained text knowledge.
+TimeLLM (Jin et al., 2024) transforms time series into text to align with LLM representations.
+Despite their positive impact, LLMs’ high inference costs limit their practicality.
+Time series foundation model: Pre-training with multiple sources time series has recently received widespread attention (Rasul et al., 2023; Dooley et al., 2024; Garza and Mergenthaler-Canseco, 2023; Kamarthi and Prakash, 2023).
+MOMENT (Goswami et al., 2024) and MOIRAI (Woo et al., 2024) adopt a BERT-style pre-training approach, while Timer (Liu et al., 2024), Chronos (Ansari et al., 2024) and TimsFM (Das et al., 2023a) use a GPT-style pre-training approach, giving rise to improved performance in time series prediction.
 However, the above methods overlook domain-specific information from multi-source data, thus limiting the performance of the models.
@@
 The time series forecasting paradigm of ROSE contains two steps: pre-training and fine-tuning.
-First, the model is pre-trained on large-scale datasets from various domains, with two pre-training tasks: reconstruction and prediction.
-We set up the self-supervised reconstruction task to help the model understand time series comprehensively and set the supervised prediction task to enhance the model’s few-shot learning ability.
-Then, ROSE fine-tunes with a target dataset in the downstream scenario.
+First, ROSE is pre-trained on large-scale datasets from various domains, with two pre-training tasks: reconstruction and prediction.
+We set up the reconstruction task to help the model understand time series comprehensively and set the prediction task to enhance the model’s few-shot and zero-shot abilities.
+Then, ROSE is fine-tuned with a target dataset in the downstream scenario.
 As shown in Figure 2, ROSE employs the encoder-decoder architecture to model time series.
-The backbone consists of multiple Transformer layers to process sequential information and effectively capture temporal dependencies [33].
-ROSE is pre-trained in a channel-independent way, which is widely used in time series forecasting [6].
+The backbone consists of multiple Transformer layers to process sequential information and effectively capture temporal dependencies (Vaswani et al., 2017).
+The reconstruction decoder and prediction decoder use the same structure as the Transformer encoder.
+They are used for reconstruction and prediction tasks respectively.
+ROSE is pre-trained in a channel-independent way, which is widely used in time series forecasting (Nie et al., 2022).
 Input Representations.
 To enhance the generalization of ROSE for adaptive transferring from multi-domains to different target domains, we model the inputs $\mathbf{x}$ with patch tokens and register tokens.
-Patch tokens are obtained by partitioning the time series using patching layers [6], to preserve local temporal information.
-Register tokens are obtained by linear mapping and clustering each of the entire time series input into discrete embedding, to capture domain-specific information, which will be introduced in Section 3.3.
+Patch tokens are obtained by partitioning the time series using patching layers (Nie et al., 2022), to preserve local temporal information.
+Register tokens that capture domain-specific information will be introduced in Section 3.3.
 3.2 Decomposed frequency learning
-As shown in Figure 1, time series data are composed of multiple combined frequency components, resulting in overlapping of different temporal variations.
+As shown in Figure 1, time series data are composed of multiple superimposed frequency components, resulting in the overlap of different temporal changes.
 Furthermore, low-frequency components typically contain information about overall trends and longer-scale variations, and high-frequency components usually contain information about short-term fluctuations and shorter-scale variations, therefore, understanding time series from low and high frequencies separately benefits general time series representation learning.
-Based on the observations above, we propose decomposed frequency learning based on multi-freq masking to understand the time series from multiple-frequency perspectives, which enhances the model’s ability to learn generalized representations.
+Based on the observations above, we propose a novel frequency-based masked modeling that randomly mask either high-frequency or low-frequency components of a time series multiple times as the key to enable learning of common time series patterns, such as trends and various long and short term fluctuations.
+Finally, reconstruction task assists the model in comprehending the data from various frequency perspectives, enabling it to learn generalized representations.
+In contrast, existing frequency masking methods (Zhang et al., 2022), which randomly mask frequencies of a single time series once, show limited forecasting effectiveness due to the lack of common pattern learning from heterogeneous time series that come from various domains.
 Figure 3: An illustration of decomposed frequency learning.
-Based on the pre-defined thresholds, we randomly apply low/high-frequency masking to the time series in the frequency domain and then transform it back to the time domain for reconstruction.
-Multi-freq masking.
-Given a time series $\mathbf{x}\in\mathbb{R}^{L}$, we utilize the Real Fast Fourier Transform (rFFT) [34] to transform it into the frequency domain, giving rise to $\mathbf{x}_{\mathrm{freq}}\in\mathbb{R}^{L/2}$.
+Based on the sampled thresholds, we randomly apply low/high-frequency masking to the time series in the frequency domain and then transform it back to the time domain for reconstruction.
+Multi-frequency masking.
+As shown in the green part of Figure 3, given a time series $\mathbf{x}\in\mathbb{R}^{L}$, we utilize the Real Fast Fourier Transform (rFFT) (Brigham and Morrow, 1967) to transform it into the frequency domain, giving rise to $\mathbf{x}_{\mathrm{freq}}\in\mathbb{C}^{L/2+1}$.
 $\mathbf{x}_{\mathrm{freq}}=\mathrm{rFFT}(\mathbf{x}).$ (1)
-To separately model high-frequency and low-frequency information in time series, we sample $K_{\mathrm{f}}$ thresholds ${\tau_{1},\tau_{2},\tau_{3},...,\tau_{K_{\mathrm{f}}}}$ and $K_{\mathrm{f}}$ random numbers ${\mu_{1},\mu_{2},\mu_{3},...,\mu_{K_{\mathrm{f}}}}$ for multi-frequency masks, where $\tau\in\mathrm{Uniform}(0,a)$, $a<L/2$ , and $\mu\in\mathrm{Bernoulli}(1,p)$.
+To separately model high-frequency and low-frequency information in time series, we sample $K_{\mathrm{f}}$ thresholds ${\tau_{1},\tau_{2},\tau_{3},...,\tau_{K_{\mathrm{f}}}}$ and $K_{\mathrm{f}}$ random numbers ${\mu_{1},\mu_{2},\mu_{3},...,\mu_{K_{\mathrm{f}}}}$ for multi-frequency masks, where $\tau\in\mathrm{Uniform}(0,a)$, $a<L/2+1$ , and $\mu\in\mathrm{Bernoulli}(p)$.
 Each pair of $\tau_{i}$ and $\mu_{i}$ corresponds to the $i_{th}$ frequency mask.
-This generates a mask matrix $\mathbf{M}\in\{0,1\}^{K_{\mathrm{f}}\times L/2}$, where each row corresponds to the $i_{th}$ frequency mask, each column corresponds to the $j_{th}$ frequency, and each element $m_{ij}$ is 0 or 1, meaning that the $j_{th}$ frequency is masked with the $i_{th}$ frequency mask or not.
-$m_{ij}=\left\{\begin{aligned} &\quad\mu_{i}&,if\,\ \mathbf{x}_{\mathrm{freq}_{j}}<\tau_{i}\\ &(1-\mu_{i})&,if\,\ \mathbf{x}_{\mathrm{freq}_{j}}>\tau_{i}\end{aligned}\right.,$ (2)
-where $\tau_{i}$ and $\mu_{i}$ denote the threshold and random number for the $i_{th}$ frequency domain mask, $\mathbf{x}_{\mathrm{freq}_{j}}$ represents the $j_{th}$ frequency of $\mathbf{x}_{\mathrm{freq}}$.
-If $\mu_{i}=1$, it indicates masking of the high frequencies, whereas if $\mu_{i}=0$, it signifies the masking of the low frequencies.
-After obtaining the mask matrix $\mathbf{M}$, we replicate $\mathbf{x}_{\mathrm{freq}}$ $K_{\mathrm{f}}$ times to get the $\mathbf{X}_{\mathrm{freq}}\in\mathbb{R}^{K_{\mathrm{f}}\times L/2}$ and perform element-wise Hadamard product with the mask matrix $\mathbf{M}$ to get masked frequency of time series.
+This generates a mask matrix $\mathbf{M}\in\{0,1\}^{K_{\mathrm{f}}\times(L/2+1)}$, where each row corresponds to the $i_{th}$ frequency mask, each column corresponds to the $j_{th}$ frequency, and each element $m_{ij}$ is 0 or 1, meaning that the $j_{th}$ frequency is masked with the $i_{th}$ frequency mask or not.
+$m_{ij}=\left\{\begin{aligned} &\quad\mu_{i}&,if\,\ j<\tau_{i}\\ &(1-\mu_{i})&,if\,\ j>\tau_{i}\end{aligned}\right.,$ (2)
+where $\tau_{i}$ and $\mu_{i}$ denote the threshold and random number for the $i_{th}$ frequency domain mask.
+If $\mu_{i}=1$, it means that frequency components above $\tau_{i}$ will be masked, indicating to mask high frequency , as shown by the threshold $\tau_{1}$ in Figure 3.
+Conversely, if $\mu_{i}=0$, it signifies that frequency components below $\tau_{i}$ will be masked, indicating to mask low frequency, exemplified by threshold $\tau_{2}$ in Figure 3.
+After obtaining the mask matrix $\mathbf{M}$, we replicate $\mathbf{x}_{\mathrm{freq}}$ $K_{\mathrm{f}}$ times to get the $\mathbf{X}_{\mathrm{freq}}\in\mathbb{C}^{K_{\mathrm{f}}\times L/2+1}$ and perform element-wise Hadamard product with the mask matrix $\mathbf{M}$ to get masked frequency of time series.
 Then, we use the inverse Real Fast Fourier Transform (irFFT) to convert the results from the frequency domain back to the time domain and get $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}=\{\mathbf{x}_{\mathrm{mask}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$, where each $\mathbf{x}_{\mathrm{mask}}^{i}\in\mathbb{R}^{L}$ corresponding to masking with a different threshold $\tau_{i}$.
@@
 Representation learning.
-After obtaining the $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}$, we divide each sequence $\mathbf{x}_{\mathrm{mask}}^{i}$ into $P$ non-overlapping patches, and use a linear layer to transforming them into $P$ patch tokens, and thus we get $\mathcal{X}_{\mathrm{mp}}=\{\mathbf{X}_{\mathrm{mp}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$ to capture general information, where each $\mathbf{X}_{\mathrm{mp}}^{i}\in\mathbb{R}^{P\times D}$, and $D$ is the dimension for each patch token.
+As shown in the yellow part of Figure 3, after obtaining the $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}$, we divide each sequence $\mathbf{x}_{\mathrm{mask}}^{i}$ into $P$ non-overlapping patches, and use a linear layer to transforming them into $P$ patch tokens, and thus we get $\mathcal{X}_{\mathrm{mp}}=\{\mathbf{X}_{\mathrm{mp}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$ to capture general information, where each $\mathbf{X}_{\mathrm{mp}}^{i}\in\mathbb{R}^{P\times D}$, and $D$ is the dimension for each patch token.
 We replicate the register tokens $\mathbf{X}_{\mathrm{u}}$ $K_{\mathrm{f}}$ times to get $\mathcal{X}_{\mathrm{u}}\in\mathbb{R}^{K_{\mathrm{f}}\times N_{\mathrm{r}}\times D}$, where $\mathbf{X}_{\mathrm{u}}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ is obtained by inputting the original sequence into the TS-Register, as detailed in Section 3.3.
 Then, we concatenate the patch tokens $\mathcal{X}_{\mathrm{mp}}$ with the register tokens $\mathcal{X}_{\mathrm{u}}$, and feed them into the Transformer encoder to obtain the representation of each masked series.
-These representations are then aggregated to yield a unified representation $\mathbf{S}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$:
-$\mathbf{S}=\mathrm{Aggregator(Encoder(Concatenate(\mathcal{X}_{mp},\mathcal{X}_{u})))}.$ (4)
+These representations are then aggregated to yield a unified representation $\mathbf{S}_{\mathrm{m}}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$.
+The aggregator is the averaging operation.
+$\mathbf{S}_{\mathrm{m}}=\mathrm{Aggregator(Encoder(Concatenate(\mathcal{X}_{mp},\mathcal{X}_{u})))}.$ (4)
 Reconstruction task.
-After obtaining the representation $\mathbf{S}$, we feed it into the reconstruction decoder and ultimately reconstruct the original sequence $\mathbf{\hat{x}}\in\mathbb{R}^{L}$ through the reconstruction head.
+After obtaining the representation $\mathbf{S}_{\mathrm{m}}$, we feed it into the reconstruction decoder, which shares same stucture as the Tranformer encoder, and ultimately reconstruct the original sequence $\mathbf{\hat{x}}\in\mathbb{R}^{L}$ through the reconstruction head, which is a linear layer.
 As frequency domain masking affects the overall time series, we compute the Mean Squared Error (MSE) reconstruction loss for the entire time series.
@@
 3.3 Time series register
-By decomposed frequency learning, we can obtain the domain-general representations, and additionally, we propose the TS-Register that learns register tokens as the domain-specific information from the multi-domain datasets for adaptive transfer.
-It clusters domain-specific information from the multi-domain datasets into register tokens and stores such domain-specific information in the register codebook during pre-training.
-Then, it adaptively selects similar information from the register codebook to enhance the target domain during fine-tuning.
-We set up a randomly initialized register codebook $\mathbf{E}\in\mathbb{R}^{H\times D_{\mathrm{r}}}$ with $H$ cluster center vectors $\mathbf{e}_{i}\in\mathbb{R}^{D_{\mathrm{r}}},i\in\{1,2,…,H\}$.
+By decomposed frequency learning, we can obtain the general representations, and additionally, we propose the TS-Register that learns register tokens as the domain-specific information from the multi-domain datasets for adaptive transfer.
+It clusters domain-specific information from the multi-domain datasets into register tokens and stores such domain-specific information in the register during pre-training.
+Then, it adaptively selects domain-specific information from the register via a Top-K selection strategy to enhance the performance in the target domain.
+A novel learnable low-rank matrix is proposed to set to complement the downstream dataset-specific information through fine-tuning.
+We set up a randomly initialized register $\mathbf{E}\in\mathbb{R}^{H\times D_{\mathrm{r}}}$ with $H$ cluster center vectors $\mathbf{e}_{i}\in\mathbb{R}^{D_{\mathrm{r}}},i\in\{1,2,…,H\}$.
 Each of input time series $\mathbf{x}\in\mathbb{R}^{L}$ is projected into a data-dependent embedding $\mathbf{x}_{\mathrm{e}}\in\mathbb{R}^{D_{\mathrm{r}}}$ through a linear layer.
 Pre-training stage.
-As shown in Figure 2(b), we use the register codebook to cluster these data-dependent embeddings, which generate domain-specific information, and store them in pre-training.
-Specifically, We find a cluster center vector $\mathbf{e}_{\delta}$ from the register codebook $\mathbf{E}$ where we use $\delta$ to denote the cluster that the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ belongs to, as shown in Equation 6.
-$\delta=\underset{j=1:H}{\arg\operatorname{min}}\left\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{j}\right\|_{2}.$ (6)
-$\mathcal{L}_{\text{register}}=\|\text{sg}(\mathbf{x}_{\mathrm{e}})-\mathbf{e}_{\delta}\|^{2}_{2}+\|\mathbf{x}_{\mathrm{e}}-\text{sg}(\mathbf{e}_{\delta})\|^{2}_{2}.$ (7)
-To update the cluster center vectors in the register codebook $\mathbf{E}$ that represent the domain information of the pre-trained datasets, we set the loss function shown in Equation 7 that minimizes the distance between the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ and the cluster center $\mathbf{e}_{\delta}$.
-$\text{sg}()$ refers to the stop gradient operation, and is used to solve the problem that the gradient of the $\arg\operatorname{min}$ function cannot be backpropagated [35].
-The first term is used to update the register codebook $\mathbf{E}$, and the second term is used to update the parameters of the linear layer that learns $\mathbf{x}_{\mathrm{e}}$.
-In this way, the vectors in the register codebook $\mathbf{E}$ cluster the embeddings of different data and learn the domain-specific centers for pre-trained datasets, which can represent domain-specific information.
-As a vector in the register codebook $\mathbf{E}$, $\mathbf{e}_{\delta}$ represents the domain-specific information for input $\mathbf{x}$.
+As shown in Figure 2(b), we use the register to cluster these data-dependent embeddings, which generate domain-specific information, and store them in pre-training.
+Specifically, We find a cluster center vector $\mathbf{e}_{\delta}$ from the register $\mathbf{E}$ where we use $\delta$ to denote the cluster that the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ belongs to.
+$\mathcal{L}_{\text{register}}=\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{\delta}\|^{2}_{2},\ \ \delta=\underset{j=1:H}{\arg\operatorname{min}}\left\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{j}\right\|_{2}.$ (6)
+To update the cluster center vectors in the register $\mathbf{E}$ that represents the domain information of the pre-trained datasets, we set the loss function shown in Equation 6 that minimizes the distance between the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ and the cluster center $\mathbf{e}_{\delta}$.
+To solve the problem that the gradient of the $\arg\operatorname{min}$ function cannot be backpropagated, we use the stop gradient operation to pass the gradient of $\mathbf{e}_{\delta}$ directly to $\mathbf{x}_{\mathrm{e}}$.
+In this way, the vectors in the register $\mathbf{E}$ cluster the embeddings of different data and learn the domain-specific centers for pre-trained datasets, which can represent domain-specific information.
+As a vector in the register $\mathbf{E}$, $\mathbf{e}_{\delta}$ represents the domain-specific information for input $\mathbf{x}$.
 $\mathbf{e}_{\delta}$ is invariant under small perturbations in $\mathbf{x}_{\mathrm{e}}$ that represents $\mathbf{x}$, which promotes better representation of domain-specific information and robustness of the vectors in the register.
-This also avoids their over-reliance on detailed information about specific datasets [36].
+This also avoids their over-reliance on detailed information about specific datasets.
 The cluster center vector $\mathbf{e}_{\delta}$ is then patched into $\mathbf{X}_{\mathrm{u}}\in\mathbb{R}^{N_{\text{r}}\times D}$, where $N_{\text{r}}$ is the number of the register tokens and $D$ is the dimensionality of Transformer latent space.
@@
 Fine-tuning stage.
-As shown in Figure 2(c), after obtaining a register codebook $\mathbf{E}$ that contains domain-specific information through pre-training, we freeze the register parameters to adaptively use this domain-specific information in the downstream task.
-Since the target domain may not strictly fit one of the upstream domains, learning the embedding of the downstream data employs a top-k strategy that selects top-k similar vectors in the register codebook.
-As shown in Equation 8, the embedding of input time series $\mathbf{x}_{\mathrm{e}}$ picks the $k$ nearest vectors in the register codebook $\mathbf{E}$, and uses their average as $\bar{\mathbf{e}}_{k}$ to represent the domain-specific information from the pre-train stage.
-$\bar{\mathbf{e}}_{k}$ is also patched into $\mathbf{X}_{\mathrm{d}}\in\mathbb{R}^{N_{\text{r}}\times D}$ and is used as the register tokens.
+As shown in Figure 2(c), after obtaining a register $\mathbf{E}$ that contains domain-specific information through pre-training, we freeze the register parameters to adaptively use this domain-specific information in the downstream tasks.
+Since the target domain may not strictly fit one of the upstream domains, we propose a novel embedding learning of the downstream data by employing a Top-K strategy that selects $k$ similar vectors in the register.
+As shown in Equation 7, the embedding of input time series $\mathbf{x}_{\mathrm{e}}$ picks the $k$ nearest vectors in the register $\mathbf{E}$, and uses their average as $\bar{\mathbf{e}}_{k}$ to represent the domain-specific information from the pre-train stage.
+$\bar{\mathbf{e}}_{k}$ is also patched into $\mathbf{X}_{\mathrm{d}}\in\mathbb{R}^{N_{\text{r}}\times D}$ and is used as the domain specific register tokens.
 $\bar{\mathbf{e}}_{k}=\frac{1}{k}\sum_{i=1}^{k}\mathbf{e}_{\delta_{i}},\ \ \left\{\delta_{1},\cdots,\delta_{k}\right\}=\underset{j=1:H}{\arg\operatorname{Topk}}(\frac{1}{\left\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{j}\right\|_{2}}).$ (7)
 Since the downstream data has its own specific information at the dataset level in addition to the domain level, this may not be fully represented by the domain information obtained from the pre-trained dataset alone.
-Therefore, we additionally set a matrix $\mathbf{A}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ to adjust $\mathbf{X}_{\mathrm{d}}$ to complement the specific information of downstream data.
-Since the pre-trained model has a very low intrinsic dimension [37], in order to get better fine-tuning results, $\mathbf{A}$ is set as a low-rank matrix:
+Therefore, we innovatively set a learnable matrix $\mathbf{A}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ to adjust $\mathbf{X}_{\mathrm{d}}$ to complement the specific information of downstream data.
+Since the pre-trained model has a very low intrinsic dimension (Aghajanyan et al., 2020), in order to get better fine-tuning results, $\mathbf{A}$ is set as a low-rank matrix:
 $\mathbf{A}=\mathbf{u}\times\mathbf{v^{\mathrm{T}}},$ (8)
@@
 3.4 Training
-To improve the model’s prediction ability when only few samples are available downstream, we co-train supervised prediction with self-supervised reconstruction that uses multi-frequency mask to learn unified features that are more applicable to the downstream prediction task.
-[…]
+To improve the model’s prediction ability in zero-shot and few-shot setting, we co-train supervised prediction with self-supervised reconstruction that uses multi-frequency masking to learn unified features that are more applicable to the downstream prediction task.
+We normalize time series by employing the REVIN (Kim et al., 2021) that is commonly used by the state-of-the-art time series models, which first normalizes each input sample and subsequently applies inverse normalization to recover the model’s output.
+Prediction task.
+The input time series $\mathbf{x}\in\mathbb{R}^{L}$ is sliced into $P$ non-overlapping patches and then mapped to $\mathbf{X}_{\mathrm{p}}\in\mathbb{R}^{P\times D}$.
+Based on common forecasted needs (Qiu et al., 2024), we set up four prediction heads mapping to prediction lengths of $\{96,192,336,720\}$ to accomplish the prediction task.
+Patch tokens $\mathbf{X}_{\mathrm{p}}$ are concatenated with the register tokens $\mathbf{X}_{\mathrm{u}}$ and then successively fed into the Transformer encoder to yield the representation $\mathbf{S}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$:
+$\mathbf{S}=\mathrm{Encoder(Concatenate(\mathcal{X}_{p},\mathcal{X}_{u}))}.$ (10)
+We feed the representation $\mathbf{S}$ into the prediction decoder and prediction heads to obtain four prediction results $\mathbf{\hat{Y}}_{F}$, where $F\in\{96,192,336,720\}$.
+With the ground truth $\mathbf{Y}_{F}$, the prediction loss $\mathcal{L}_{\text{prediction}}$ is shown in Equation 11.
 $\mathcal{L}_{\text{prediction}}=\sum_{F\in\{96,192,336,720\}}||{\mathbf{Y}_{F}-{\mathbf{\hat{Y}}_{F}}}||_{2}^{2}.$ (11)
@@
 To avoid prediction training affecting the generalization performance of the model, the gradients of the prediction heads are skipped at back-propagation.
-The overall pre-training loss is shown in Equation 12.
+The overall loss of ROSE in pre-training stage is shown in Equation 12.
 $\mathcal{L_{\text{pre-train}}}=\mathcal{L}_{\text{reconstruction}}+\mathcal{L}_{\text{prediction}}+\mathcal{L}_{\text{register}}.$ (12)
@@
 Pre-training datasets.
-Large-scale datasets are crucial for pre-training a general time series forecasting model.
+The datasets are crucial for pre-training a general time series forecasting model.
 In light of this, we gather a considerable amount of publicly available datasets from various domains, including energy, nature, health, transport, web, and economics, etc.
-The details of these datasets are shown in the Appendix A.1.
-To enhance data utilization, we downsampled fine-grained datasets to coarser granularity, resulting in approximately 886 million time points.
+The details of these datasets are shown in the Appendix A.1.1.
+To enhance data utilization, we downsample fine-grained datasets to coarser granularity, resulting in approximately 887 million time points.
 Evaluation datasets.
-To conduct comprehensive and fair comparisons for different models, we conducted experiments on eight real-world benchmarks as the target datasets, including Weather, Solar, Traffic, Electricity, and ETT (4 subsets), which cover multiple domains.
+To conduct comprehensive and fair comparisons for different models, we conducted experiments on seven well-known forecasting benchmarks as the target datasets, including Weather, Traffic, Electricity, and ETT (4 subsets), which cover multiple domains.
 Baselines.
-We select the state-of-the-art models as our baselines, including Transformer-based models: Timer [2], iTransformer [20], and PatchTST [6]; CNN-based model: TimesNet [18]; MLP-based models: FITS [38], TIDE [39], and DLinear [40].
-It is worth noting that Timer is a recent general time series model with a GPT-style architecture.
-Since the size of the look-back window can affect the performance of different models, we choose the look-back window size in $\{96,336,512,720\}$ for each baseline that achieves its best results for fair comparisons.
+We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023), and two LLM-based models: GPT4TS (Zhou et al., 2024) and $\textbf{S}^{2}$-LLM (Pan et al., 2024).
+In addition, we select five foundation models for comparison in zero-shot setting, including Timer (Liu et al., 2024), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., 2024), TimesFM (Das et al., 2023b), and Moment (Goswami et al., 2024).
 Setup.
 Consistent with previous works, we adopted Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics.
-All methods predict the future values with lengths $F=\{96,192,336,720\}$.
-More implementation details are presented in the Appendix A.3.
-4.1 Main results
-Full-shot setting.
-We first pre-train ROSE on our pre-training datasets and then fine-tune with full downstream data (full-shot setting).
-Similarly, all baselines are also run in the same full-shot setting.
-As shown in Table 1, we also present the results of the ROSE in 10% few-shot setting, where the model uses the same validation and test data but is only fine-tuned with 10% training data of each target dataset.
+For fair comparison, all methods fix the look-back window $L$ = 512 and predict the future values with lengths $F=\{96,192,336,720\}$.
+More implementation details are presented in the Appendix A.1.3.
+4.1 In distribution forecasting
+Setting. In full-shot setting, we fine-tune pre-trained ROSE and the baselines with the full downstream data. In few-shot setting, we fine-tune all models with only 10% train data.
+Full-shot results.
+As shown in Table 1, we also present the results of the ROSE in 10% few-shot setting.
 Key observations are summarized as follows.
-First, as a general forecasting model, ROSE achieves superior performance compared to the seven state-of-the-art baselines with full-data training, especially Timer, which shows that our decomposed frequency learning and register help to learn generalized representations from large-scale datasets and adaptively transfer the multi-domain information to specific downstream scenarios.
-Second, we observe that ROSE achieves competitive performance in 10% few-shot setting, even compared to baselines trained with full data.
+First, as a general forecasting model, ROSE achieves superior performance compared to the six state-of-the-art baselines with full-data training, achieving an average MSE reduction of 15%, which shows that our decomposed frequency learning and register help to learn generalized representations from large-scale datasets and adaptively transfer the multi-domain information to specific downstream scenarios.
+Second, we observe that ROSE in 10% few-shot setting shockingly improves a large margin as MSE reduction in average exceeding 12% over the baselines trained with full data.
 This observation validates the transferability of ROSE pre-trained with large multi-source data.
+Table 1: The results for ROSE in full-shot setting and 10% few-shot setting, compared with other methods in full-shot setting. The average results of all predicted lengths are listed here.
+Models ROSE ROSE (10%) ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$-LLM
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.391 0.414 0.397 0.419 0.439 0.448 0.413 0.434 0.582 0.533 0.416 0.436 0.427 0.426 0.406 0.427
+ETTh2 0.331 0.374 0.335 0.380 0.374 0.406 0.331 0.381 0.409 0.438 0.508 0.485 0.354 0.394 0.347 0.391
+ETTm1 0.341 0.367 0.349 0.372 0.362 0.391 0.353 0.382 0.490 0.464 0.356 0.378 0.352 0.383 0.343 0.379
+ETTm2 0.246 0.305 0.250 0.308 0.269 0.329 0.256 0.317 0.317 0.358 0.259 0.325 0.266 0.326 0.257 0.319
+Weather 0.217 0.251 0.224 0.252 0.233 0.271 0.226 0.264 0.329 0.336 0.239 0.289 0.237 0.270 0.222 0.259
+Electricity 0.155 0.248 0.164 0.253 0.164 0.261 0.159 0.253 0.195 0.296 0.166 0.267 0.167 0.263 0.161 0.257
+Traffic 0.390 0.264 0.418 0.278 0.397 0.282 0.391 0.264 0.623 0.333 0.433 0.305 0.414 0.294 0.405 0.286
+Few-shot results. The results under the 10% few-shot setting are presented in Table 13 in Appendix A.10.2. ROSE outperforms advanced models when training data is scarce in the target domain.
 Figure 4: The forecasting results of ROSE obtained by training from scratch and fine-tuning from the pre-trained model. The right, upper corner is the best case.
-Few-shot setting.
-We also compared all models under the 10% few-shot setting in Table 9 in Appendix H.2.
-We can see that ROSE performs the best on all eight datasets in the scenarios where the training data are rare in the target domain, surpassing extensive advanced models.
-Figure 4 shows the results of pre-trained ROSE and ROSE trained from scratch on the Solar and ETTh1 with different fine-tuning data percentages, noting the best baselines in full-shot setting.
-The pre-trained ROSE shows stable, superior performance even with limited fine-tuned samples.
-Specifically, the pre-trained ROSE achieves performance that surpasses the SOTA with only 5% for Solar and 1% data for ETTh1.
-Moreover, compared to the ROSE trained from scratch, the pre-trained ROSE exhibits a slower decline in predictive performance with the reduction of fine-tuning data, demonstrating the impressive generalization ability of ROSE through large-scale pre-training.
-Table 1: The results for ROSE in full-shot setting and 10% few-shot setting, compared with other methods in full-shot setting. The average results of all predicted lengths are listed here.
-Models ROSE ROSE 10% Timer ITransformer PatchTST FITS TimesNet TiDE DLinear
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1 0.341 0.367 0.349 0.372 0.436 0.456 0.407 0.410 0.353 0.382 0.355 0.379 0.400 0.406 0.355 0.378 0.357 0.379
-ETTm2 0.246 0.305 0.249 0.308 0.308 0.351 0.288 0.332 0.256 0.317 0.249 0.312 0.291 0.333 0.249 0.312 0.267 0.332
-ETTh1 0.391 0.414 0.397 0.419 0.483 0.459 0.454 0.447 0.413 0.434 0.407 0.429 0.458 0.450 0.419 0.430 0.423 0.437
-ETTh2 0.330 0.374 0.335 0.380 0.366 0.406 0.383 0.407 0.331 0.381 0.333 0.382 0.414 0.427 0.345 0.394 0.431 0.447
-Traffic 0.390 0.264 0.418 0.278 0.413 0.284 0.428 0.282 0.391 0.264 0.410 0.282 0.62 0.336 0.356 0.261 0.434 0.295
-Weather 0.217 0.251 0.224 0.252 0.277 0.306 0.246 0.298 0.226 0.264 0.218 0.260 0.259 0.287 0.236 0.282 0.246 0.300
-Solar 0.188 0.215 0.192 0.216 0.251 0.262 0.233 0.262 0.207 0.294 0.223 0.264 0.301 0.319 0.223 0.288 0.230 0.295
-Electricity 0.155 0.248 0.164 0.253 0.181 0.274 0.178 0.270 0.159 0.253 0.163 0.257 0.192 0.295 0.159 0.257 0.166 0.264
+Figure 4 shows the performance of pre-trained ROSE and ROSE trained from scratch on ETTh1 and ETTm2 with different fine-tuning data percentages, noting the best baselines in full-shot setting.
+The pre-trained ROSE shows stable, superior performance even with limited fine-tuning samples.
+Specifically, the pre-trained ROSE exceeds SOTA performance with only 1% train data for ETTh1 and 2% for ETTm2.
+Moreover, compared to the ROSE trained from scratch, the pre-trained ROSE exhibits a slower decline in predictive performance with the reduction of fine-tuning data, demonstrating the impressive generalization ability of ROSE through pre-training.
+4.2 Zero-shot forecasting
+Setting.
+In this section, to ensure a fair comparison, we conduct zero-shot predictions for each foundational model on downstream datasets not included in their pre-training data.
+It is worth noting that, unlike a few foundation models (Woo et al., 2024) that require much longer inputs to achieve better predictive performance, we fix the input length of all baselines to 512 without considering longer input lengths, as many real-world scenarios could offer very limited samples.
+Results.
+As shown in Table 2, ROSE significantly outperforms across the majority of datasets, achieving an average reduction of 15% in Mean Squared Error (MSE).
+In comparison to Timer and Moirai, ROSE achieves average MSE reductions of 9% and 6%, respectively, and demonstrates a remarkable 43% relative improvement over Moment.
+Notably, ROSE stands out not only for its superior performance but also for its exceptional lightweight and efficient design, which sets it apart from other foundational models.
+Detailed analysis of these aspects will be presented in Section 4.3.
+Table 2: The results for ROSE and other foundation models in zero-shot setting.
+The average results of all predicted lengths are listed here.
+We use ’-’ to indicate that the dataset has been involved in the model’s pre-training, and thus not used for testing.
+Models ROSE Timer MOIRAI Chronos TimesFM Moment
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.401 0.425 0.451 0.463 0.475 0.443 0.560 0.452 0.489 0.444 0.708 0.580
+ETTh2 0.346 0.394 0.366 0.408 0.379 0.396 0.392 0.397 0.396 0.405 0.392 0.430
+ETTm1 0.525 0.471 0.544 0.476 0.714 0.507 0.636 0.495 0.434 0.419 0.697 0.555
+ETTm2 0.299 0.352 0.360 0.386 0.343 0.356 0.313 0.363 0.320 0.353 0.319 0.360
+Weather 0.265 0.305 0.292 0.312 0.267 0.300 0.288 0.310 - - 0.291 0.323
+Electricity 0.234 0.320 0.297 0.375 0.241 0.328 0.245 0.312 - - 0.861 0.766
+Traffic 0.588 0.412 0.613 0.407 - - 0.371 0.370 - - 1.411 0.804
+4.3 Model Analysis
+Efficiency analysis.
+To exhibit the performance and efficiency advantages of ROSE, we compare its parameter count to other foundation models and evaluate their performance and testing time averaged on ETTh1 and ETTh2 datasets in zero-setting.
+Similarly, for each specific model, we evaluate its parameter count as well as its performance in full-shot setting and training-to-testing time averaged on the same datasets.
+Specific implementation details and results can be found in Appendix A.2.
+Figure 5: Model performance, number of parameter and efficiency comparison.
+As shown in Figure 5, ROSE is a lightweight general model with 7.4M parameters and short inference time, which are only about one-tenth of the second fastest/smallest foundation model (Timer).
+Importantly, ROSE uses the least number of parameters among foundation models, with its parameter count approaching that of specific models, while exhibiting superior zero-shot performance.
+This is attributed to our proposed decomposed frequency learning that enhances the model’s comprehension of time series.
+Concurrently, the TS-Register achieves the adaptive transfer thus efficiently adapting to downstream tasks without the need of scaling up to achieve strong generalizability.
+Compared to foundation models with large scale, ROSE may better meet the need for general models in real scenarios that require high computational and parameter efficiency as well as high prediction accuracy with scarce downstream data.
+Visualization of TS-Register.
+To validate the TS-Register’s capability to transfer domain-specific information adaptively from pre-training datasets to target datasets, we visualize the cosine similarity of register vector selections from datasets across different domains.
+As shown in Figure 6(a), the cosine similarity is higher for datasets within the same domain and lower between different domains.
+We also visualize the register vector selections from different datasets in Figures 6(b) and (c), where datasets from the same domain show similar visualizations.
+This confirms the TS-Register’s capability of adaptive transfer from multi-source to target datasets across various domains.
+Figure 6: Visualization of TS-Register. The calculation of cosine similarity is in the Appendix A.3.
+Scalability and Sensitivity.
+The scalability analysis of ROSE’s model size and pre-training data size are presented in Appendix A.4.
+The sensitivity analyses for the upper bound $a$ of the thresholds, the number of masked series $K_{\mathrm{f}}$, the number of register tokens $N_{\text{r}}$, the size of register $H$ and number of selections $k$ in Top-K strategy are presented in Appendix A.5.
 4.4 Ablation studies
@@
 Without the reconstruction task, our model shows negative transfer effects on ETTm1 and ETTm2, likely due to the prediction task making the model more susceptible to pre-training data biases.
-Mask method.
-To further validate the effectiveness of decomposed frequency learning, we replace the multi-frequency masking with different masking methods, including two mainstream time-domain methods: patch masking [6] and multi-patch masking [5], as well as a single frequency masking, which masks time series in the frequency domain only once.
-The results show that frequency domain modeling yields superior performance improvements compared to time-domain modeling.
-Additionally, we observe that applying multi-frequency masking results in enhanced outcomes, thus affirming that incorporating a multi-frequency perspective significantly aids the model in more accurately comprehending temporal patterns.
-Table 2: Ablations on ROSE. We replace different masking methods, in addition to component removal. The average results of all predicted lengths are listed here.
+Masking method.
+To further validate the effectiveness of decomposed frequency learning, we replace the multi-frequency masking with different masking methods, including two mainstream time-domain methods: patch masking (Nie et al., 2022) and multi-patch masking (Dong et al., 2024), as well as random frequency masking (Chen et al., 2023b).
+The results in Table 4 show that random frequency masking and patch masking led to negative transfer on ETTm1 and ETTm2, likely due to significant disruption of the original time series, causing overfitting.
+In contrast, multi-patch masking and multi-frequency masking resulted in positive transfer across all datasets by preventing excessive disruption.
+Multi-frequency masking achieved better results, demonstrating its ability to help the model understand temporal patterns from a multi-frequency perspective.
+We also compare with some other pre-training tasks in Table 14 in Appendix A.10.3.
+Table 3: Ablations on key components of model architecture, including TS-register, prediction task and reconstruction task. The average results of all predicted lengths are listed here.
 Design ETTm1 ETTm2 ETTh1 ETTh2
@@
 ROSE 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.380
-Replace Multi-Freq Masking Single Freq Masking 0.354 0.379 0.257 0.313 0.403 0.423 0.347 0.387
-Multi Patch Masking 0.356 0.379 0.259 0.316 0.404 0.426 0.349 0.389
-Patch Masking 0.378 0.4 0.261 0.319 0.408 0.432 0.375 0.407
-w/o TS-Register 0.354 0.378 0.256 0.312 0.418 0.427 0.355 0.39
-Predict Task 0.36 0.384 0.257 0.314 0.422 0.438 0.372 0.41
-Reconstruct Task 0.387 0.403 0.269 0.327 0.412 0.428 0.361 0.399
-From Scratch 0.370 0.393 0.261 0.318 0.470 0.480 0.401 0.425
-4.3 Scalability
+w/o TS-Register 0.354 0.378 0.256 0.312 0.418 0.427 0.355 0.390
+Prediction Task 0.360 0.384 0.257 0.314 0.422 0.438 0.372 0.410
+Reconstruction Task 0.387 0.403 0.269 0.327 0.412 0.428 0.361 0.399
+From scratch 0.371 0.391 0.261 0.318 0.470 0.480 0.400 0.425
+Table 4: Ablations on decomposed frequency learning, where we replace Multi-freq masking with other masking methods. The average results of all predicted lengths are listed here.
+Design ETTm1 ETTm2 ETTh1 ETTh2
+MSE MAE MSE MAE MSE MAE MSE MAE
+ROSE 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.380
+Replace Multi-Frequency Masking Random Freq Masking 0.381 0.397 0.261 0.324 0.410 0.427 0.374 0.405
+Multi-Patch Masking 0.356 0.379 0.259 0.316 0.404 0.426 0.349 0.389
+Patch Masking 0.378 0.400 0.261 0.319 0.408 0.432 0.375 0.407
+From scratch 0.371 0.391 0.261 0.318 0.470 0.480 0.400 0.425
+5 Conclusion and future work
+In this work, we propose ROSE, a novel general model, addressing the challenges of leveraging multi-domain datasets for enhancing downstream prediction task performance.
+ROSE utilizes decomposed frequency learning and TS-Register to capture generalized and domain-specific representations, enabling improved fine-tuning results, especially in data-scarce scenarios.
+Our experiments demonstrate ROSE’s superior performance over baselines with both full-data and few-data fine-tuning, as well as its impressive zero-shot capabilities.
+Future efforts will concentrate on expanding pre-training datasets and extending ROSE’s applicability across diverse time series analysis tasks, e.g., classification.
+We provide our code at https://anonymous.4open.science/r/ROSE-A235.
+6 Reproducibility
+Our work meets reproducibility requirements.
+Specifically, you can download our evaluation datasets in a standardized format from the public link (Wang et al., 2024a): https://drive.google.com/drive/folders/13Cg1KYOlzM5C7K8gK8NfC-F3EYxkM3D2, and we provide our code at https://anonymous.4open.science/r/ROSE-A235.
+Appendix A Appendix
+A.1 IMPLEMENTATION DETAILS
+A.1.1 Pre-training Datasets
+We use multi-source datasets in pre-training which contain subsets of Monash [Godahewa et al., 2021], UEA [Bagnall et al., 2018] and UCR [Dau et al., 2019] time series datasets, as well as some other time series classical datasets [Zhang et al., 2017, Wang et al., 2024b, Liu et al., 2022b, McCracken and Ng, 2016, Taieb et al., 2012].
+The final list of all pre-training datasets is shown in Table 6.
+There is no overlap between the pre-training datasets and the target datasets.
+It is worth noting that the dataset weather in the pre-training dataset is a univariate dataset, which is different to the multivariate dataset weather in the target task.
+The pre-trained datasets can be categorized into 6 different domains according to their sources: Energy, Nature, Health, Transport, and Web.
+The sampling frequencies of the datasets show a remarkable diversity, ranging from millisecond samples to monthly samples, which reflects the diverse application scenarios and complexity of the real world.
+For all pre-training datasets, we split them into univariate sequences and train them in a channel-independent manner.
+A.1.2 Evaluation Datasets
+We use the following 7 multivariate time-series datasets for downstream fine-tuning and forecasting: ETT datasets [^1] https://github.com/zhouhaoyi/ETDataset contain 7 variates collected from two different electric transformers from July 2016 to July 2018.
+It consists of four subsets, of which ETTh1/ETTh2 are recorded hourly and ETTm1/ETTm2 are recorded every 15 minutes.
+Traffic [^2] https://pems.dot.ca.gov/ contains road occupancy rates measured by 862 sensors on freeways in the San Francisco Bay Area from 2015 to 2016, recorded hourly.
+Weather [^3] https://www.bgc-jena.mpg.de/wetter/ collects 21 meteorological indicators, such as temperature and barometric pressure, for Germany in 2020, recorded every 10 minutes.
+Electricity [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 contains the electricity consumption of 321 customers from July 2016 to July 2019, recorded hourly.
+We split each evaluation dataset into train-validation-test sets and detailed statistics of evaluation datasets are shown in Table 5.
+Table 5: The statistics of evaluation datasets.
+Dataset ETTm1 ETTm2 ETTh1 ETTh2 Traffic Weather Electricity
+Variables 7 7 7 7 862 21 321
+Timestamps 69680 69680 17420 17420 17544 52696 26304
+Split Ratio 6:2:2 6:2:2 6:2:2 6:2:2 7:1:2 7:1:2 7:1:2
+Table 6: List of pretraining datasets.
+Domain Dataset Frequency Time Pionts Source
+Energy Aus. Electricity Demand Half Hourly 1155264 Monash[Godahewa et al., 2021]
+Wind 4 Seconds 7397147 Monash[Godahewa et al., 2021]
+Wind Farms Minutely 172178060 Monash[Godahewa et al., 2021]
+Solar 10 Minutes 7200720 Monash[Godahewa et al., 2021]
+Solar Power 4 Seconds 7397222 Monash[Godahewa et al., 2021]
+London Smart Meters Half Hourly 166527216 Monash[Godahewa et al., 2021]
+Nature Phoneme - 2160640 UCRDau et al. [2019]
+EigenWorms - 27947136 UEA[Bagnall et al., 2018]
+PRSA Hourly 4628448 [Zhang et al., 2017]
+Temperature Rain Daily 23252200 Monash[Godahewa et al., 2021]
+StarLightCurves - 9457664 UCR[Dau et al., 2019]
+Worms 0.033 Seconds 232200 UCR[Dau et al., 2019]
+Saugeen River Flow Daily 23741 Monash[Godahewa et al., 2021]
+Sunspot Daily 73924 Monash[Godahewa et al., 2021]
+Weather Daily 43032000 Monash[Godahewa et al., 2021]
+KDD Cup 2018 Daily 2942364 MonashGodahewa et al. [2021]
+US Births Daily 7305 Monash[Godahewa et al., 2021]
+Health MotorImagery 0.001 Seconds 72576000 UEA[Bagnall et al., 2018]
+SelfRegulationSCP1 0.004 Seconds 3015936 UEA[Bagnall et al., 2018]
+SelfRegulationSCP2 0.004 Seconds 3064320 UEA[Bagnall et al., 2018]
+AtrialFibrillation 0.008 Seconds 38400 UEA[Bagnall et al., 2018]
+PigArtPressure - 624000 UCR[Dau et al., 2019]
+PIGCVP - 624000 UCR[Dau et al., 2019]
+TDbrain 0.002 Seconds 79232703 [Wang et al., 2024b]
+Transport Pems03 5 Minute 9382464 [Liu et al., 2022b]
+Pems04 5 Minute 5216544 [Liu et al., 2022b]
+Pems07 5 Minute 24921792 [Liu et al., 2022b]
+Pems08 5 Minute 3035520 [Liu et al., 2022b]
+Pems-bay 5 Minute 16937700 [Liu et al., 2022b]
+Pedestrian_Counts Hourly 3132346 Monash[Godahewa et al., 2021]
+Web Web Traffic Daily 116485589 Monash[Godahewa et al., 2021]
+Economic FRED_MD Monthly 77896 [McCracken and Ng, 2016]
+Bitcoin Daily 75364 Monash[Godahewa et al., 2021]
+NN5 Daily 87801 [Taieb et al., 2012]
+A.1.3 Setting
+We implemented ROSE in PyTorch [Paszke et al., 2019] and all the experiments were conducted on 8 NVIDIA A800 80GB GPU.
+We used ADAM [Kingma and Ba, 2014] with an initial learning rate of $5\times 10^{-4}$ and implemented learning rate decay using the StepLR method to implement learning rate decaying pre-training.
+By default, ROSE contains 3 encoder layers and 3 decoder layers with head number of 16 and the dimension of latent space $D=256$.
+The patch size for patching is set to 64.
+Pre-training.
+We use $N_{\mathrm{r}}=3$ as the number of register tokens and $P=8$ as the path tokens.
+We set the input length to 512 for the supervised prediction task with target lengths of 96, 192, 336, and 720.
+We also set the input length to 512 and mask number $K_{\mathbf{f}}=4$.
+The batch size is set to 8192 in pre-training.
+Fine-tuning.
+We fix the lookback window to 512, and perform predictions with target lengths of 96, 192, 336, and 720, respectively.
+The number of register tokens $N_{\mathrm{r}}$ and patch tokens $P$ is the same as in pre-training, and the parameter $k=3$ in TopK is set when selection vectors are performed in the register.
+A.1.4 Baselines
+We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer [Liu et al., 2023], PatchTST [Nie et al., 2022], TimesNet [Wu et al., 2022], and DLinear [Zeng et al., 2023], and two LLM-based models: GPT4TS [Zhou et al., 2024] and $\textbf{S}^{2}$-LLM [Pan et al., 2024].
+In addition, we selected five foundation models for comparison in zero-shot setting, including Timer [Liu et al., 2024], MOIRAI [Woo et al., 2024], Chronos [Ansari et al., 2024], TimesFM [Das et al., 2023b] and Moment [Goswami et al., 2024].
+The specific code base for these models is listed in Table 7:
+Table 7: Code repositories for baselines.
+Model Types Models Code Repositories
+Small Model iTransformer https://github.com/thuml/iTransformer
+PatchTST https://github.com/yuqinie98/PatchTST
+TimesNet https://github.com/thuml/TimesNet
+Dlinear https://github.com/cure-lab/LTSF-Linear
+Foundation Model Timer https://github.com/thuml/Large-Time-Series-Model
+MOIRAI https://github.com/redoules/moirai
+Chronos https://github.com/amazon-science/chronos-forecasting
+TimesFM https://github.com/google-research/timesfm/
+Moment https://anonymous.4open.science/r/BETT-773F/README.md
+LLM-based Model GPT4TS https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All
+S2P-LLM https://github.com/panzijie825/S2IP-LLM
+A.2 Efficiency analysis
+As an important aspect of foundation models, inference efficiency is crucial.
+Therefore, we evaluate the testing time of ROSE and five foundation models in the ETTh1 and ETTh2 dataset in zero-shot setting.
+Similarly, we evaluate the time of the entire process of training, validation, and testing for four specific models in the same datasets in full-shot setting.
+The above experiments all set the batch size to 32.
+The specific results are shown in Table 8 and Figure 5.
+We observe that ROSE maintains its advantage in zero-shot performance while also being significantly faster compared to the baselines, even being approximately ten times faster than the second-fastest foundation model, Timer.
+This raises our reflection on whether time-series foundation models require extremely large parameter sizes and whether existing time-series foundation models have validated their architectures’ scaling laws on time-series data.
+Table 8: Efficiency analysis.
+Model Parameters Pre-train datasize Averaged time
+ROSE 7.4M 0.89B 0.652s
+MOIRAI 311M 27B 7.920s
+Timer 67.4M 1B 5.989s
+Chronos 46M 84B 176s
+TimesFM 200M 100B 167.2s
+Moment 385M 1.13B 88s
+Itransformer 3.8M - 34.18s
+PatchTST 3.2M - 35.47s
+TimesNet 1.8M - 146s
+Dlinear 0.018M - 24.06s
+A.3 Calculation of cosine similarity
+In Figure 6, we visualize the cosine similarity of register vector selections.
+For each sample in a dataset, during the inference process, $k$ vectors are selected from the register based on the Top-K strategy.
+We iterate through all samples in the dataset and count the number of times each vector is selected, which allows us to obtain a record vector of length equivalent to the size of the register for each dataset.
+The $i_{th}$ position in the record vector represents the number of times the $i_{th}$ vector in the register has been selected by samples in the dataset.
+For a pair of datasets, we can obtain a unique record vector for each dataset, and then we are able to calculate the cosine similarity of two vectors.
+A.4 Scalability
 Scalability is crucial for a general model, enabling significant performance improvements by expanding pre-training data and model sizes.
@@
 Data size.
-When keeping the model size, we increase the size of the pre-training datasets from 178M to 886M.
+When keeping the model size, we increase the size of the pre-training datasets from 178M to 887M.
 The results are shown in Figure 7(c) and (d).
 The performance of our model steadily improves with the increase in dataset size and achieves improvements of 7.4% and 4.8% respectively.
-4.4 Model Analysis
-Zero-shot and inference efficiency. By adopting the dual pre-train task, our model possesses strong generalization ability and can forecast without requiring any downstream data for fine-tuning.
-Figure 6: Model performance in zero-shot setting comparison and inference efficiency comparison.
-To validate ROSE’s zero-shot capabilities and assess its inference efficiency, we compared it with other models that have recently demonstrated similar abilities, namely Timer [2], MOIRAI [8], PatchTST [6], and GPT4TS [10].
-Specific implementation details and results can be found in Appendix B.
-As shown in Figure 6, ROSE exhibits superior performance in zero-shot setting, with significantly fewer model parameters and inference time compared to other models.
-This underscores the robust generalization and scalability of ROSE.
-Visualization of TS-Register.
-To validate the TS-Register’s capability to transfer domain-specific information adaptively from pre-training datasets to target datasets, we visualize the cosine similarity of register vector selections from datasets across different domains.
-As shown in Figure 7(a), the cosine similarity is higher for datasets within the same domain and lower between different domains.
-We also visualize the register vector selections from different datasets in Figures 7(b) and (c), where datasets from the same domain show similar visualizations.
-This confirms the TS-Register’s capability of adaptive transfer from multi-source to target datasets across various domains.
-Figure 7: Visualization of TS-Register.
-Sensitivity.
-The sensitivity analyses for the upper bound $a$ of the thresholds, the number of masked series $K_{\mathrm{f}}$, the number of register tokens $N_{\text{r}}$ and the size of codebook $H$ are presented in Appendix C.
-5 Conclusion and future work
-In this work, we propose ROSE, a novel general model, addressing the challenges of leveraging multi-domain datasets for enhancing downstream prediction task performance.
-ROSE utilizes decomposed frequency learning and TS-Register to capture generalized and domain-specific representations, enabling improved fine-tuning results, especially in data-scarce scenarios.
-Our experiments demonstrate ROSE’s superior performance over baselines with both full-data and few-data fine-tuning, as well as its impressive zero-shot capabilities.
-Future efforts will concentrate on expanding pre-training datasets and extending ROSE’s applicability across diverse time series analysis tasks, e.g., classification.
-Appendix A IMPLEMENTATION DETAILS
-A.1 Pre-training Datasets
-We use multi-source datasets in pre-training which contain subsets of Monash [41], UEA [42] and UCR [43] time series regression datasets, as well as some other time series classical datasets [44, 45, 46, 47, 48].
-The final list of all pre-training datasets is shown in Table 4.
-There is no overlap between the pre-training datasets and the target datasets.
-It is worth noting that the dataset weather in the pre-training dataset is a univariate dataset, which is different to the multivariate dataset weather in the target task.
-The pre-trained datasets can be categorized into 6 different domains according to their sources: Energy, Nature, Health, Transport, and Web.
-The sampling frequencies of the datasets show a remarkable diversity, ranging from millisecond samples to monthly samples, which reflects the diverse application scenarios and complexity of the real world.
-For all pre-training datasets, we split them into univariate sequences and train them in a channel-independent manner.
-A.2 Evaluation Datasets
-We use the following 8 multivariate time-series datasets for downstream fine-tuning and forecasting: ETT datasets contain 7 variates collected from two different electric transformers from July 2016 to July 2018.
-ETT datasets [^1] https://github.com/zhouhaoyi/ETDataset consist of four subsets, of which ETTh1/ETTh2 are recorded hourly and ETTm1/ETTm2 are recorded every 15 minutes.
-Traffic [^2] https://pems.dot.ca.gov/ contains road occupancy rates measured by 862 sensors on freeways in the San Francisco Bay Area from 2015 to 2016, recorded hourly.
-Weather [^3] https://www.bgc-jena.mpg.de/wetter/ collects 21 meteorological indicators, such as temperature and barometric pressure, for Germany in 2020, recorded every 10 minutes.
-Solar [^4] https://www.nrel.gov/grid/solar-power-data.html records solar power generation from 137 PV plants in 2006, every 10 minutes.
-Electricity [^5] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 contains the electricity consumption of 321 customers from July 2016 to July 2019, recorded hourly.
-We split each evaluation dataset into train-validation-test sets and detailed statistics of evaluation datasets are shown in Table 3.
-Table 3: The statistics of evaluation datasets
-Dataset ETTm1 ETTm2 ETTh1 ETTh2 Traffic Weather Solar Electricity
-Variables 7 7 7 7 862 21 137 321
-Timestamps 69680 69680 17420 17420 17544 52696 52560 26304
-Split Ratio 6:2:2 6:2:2 6:2:2 6:2:2 7:1:2 7:1:2 7:1:2 7:1:2
-Table 4: List of pretraining datasets.
-Domain Dataset Frequency Time Pionts Source
-Energy Aus. Electricity Demand Half Hourly 1155264 Monash[41]
-Wind 4 Seconds 7397147 Monash[41]
-Wind Farms Minutely 172178060 Monash[41]
-London Smart Meters Half Hourly 166527216 Monash[41]
-Nature Phoneme - 2160640 UCR[43]
-EigenWorms - 27947136 UEA[42]
-PRSA Hourly 4628448 [44]
-Temperature Rain Daily 23252200 Monash[41]
-StarLightCurves - 9457664 UCR[43]
-Worms 0.033 Seconds 232200 UCR[43]
-Saugeen River Flow Daily 23741 Monash[41]
-Sunspot Daily 73924 Monash[41]
-Weather Daily 43032000 Monash[41]
-KDD Cup 2018 Daily 2942364 Monash[41]
-US Births Daily 7305 Monash[41]
-Health MotorImagery 0.001 Seconds 72576000 UEA[42]
-SelfRegulationSCP1 0.004 Seconds 3015936 UEA[42]
-SelfRegulationSCP2 0.004 Seconds 3064320 UEA[42]
-AtrialFibrillation 0.008 Seconds 38400 UEA[42]
-PigArtPressure - 624000 UCR[43]
-PIGCVP - 624000 UCR[43]
-TDbrain 0.002 Seconds 79232703 [45]
-Transport Pems03 5 Minute 9382464 [46]
-Pems04 5 Minute 5216544 [46]
-Pems07 5 Minute 24921792 [46]
-Pems08 5 Minute 3035520 [46]
-Pems-bay 5 Minute 16937700 [46]
-Pedestrian_Counts Hourly 3132346 Monash[41]
-Web Web Traffic Daily 116485589 Monash[41]
-Economic FRED_MD Monthly 77896 [47]
-Bitcoin Daily 75364 Monash[41]
-NN5 Daily 87801 [48]
-A.3 Setting
-We implemented ROSE in PyTorch [49] and all the experiments were conducted on 8 NVIDIA A800 80GB GPU.
-We used ADAM [50] with an initial learning rate of $5\times 10^{-4}$ and implemented learning rate decay using the StepLR method to implement learning rate decaying pre-training.
-By default, ROSE contains 3 encoder layers and 3 decoder layers with head number $H=16$ and the dimension of latent space $D=256$.
-Pre-training.
-We use $N_{\mathrm{r}}=3$ as the number of register tokens and $P=8$ as the path tokens.
-We set the input length to 512 for the supervised prediction task with target lengths of 96, 192, 336, and 720.
-We also set the input length to 512 for self-supervised multi-frequency mask-reconstruction task, where mask number $K_{\mathbf{f}}=4$.
-The batch size is set to 8192 in pre-training.
-Fine-tuning.
-We fix the lookback window to 512, and perform predictions with target lengths of 96, 192, 336, and 720, respectively.
-The number of register tokens $N_{\mathrm{r}}$ and patch tokens $P$ is the same as in pre-training, and the parameter $k=3$ in topk is set when selection vectors are performed in the register codebook.
-The batch size is set to 64 in fine-tuning.
-A.4 Baselines
-We select the state-of-the-art methods as our baselines to compare the performance of different models with full-data fine-tuning and 10% few-shot setting, including Transformer-based methods: Timer [2], iTransformer [20], and PatchTST [6]; CNN-based backbones method: TimesNet [18]; MLP-based methods: FITS [38], TIDE [39], and DLinear [40].
-It is worth noting that Timer is a recent general time series model with a GPT-style architecture.
-Since the size of the look-back window can affect the performance of different models, we choose the look-back window size in $\{96,336,512,720\}$ for each baseline that achieves its best results for fair comparisons.
-The specific code base for these models is listed in Table 5:
-Table 5: Code repositories for baselines
-Baselines code repositories
-Timer https://github.com/thuml/Large-Time-Series-Model
-iTransformer https://github.com/thuml/iTransformer
-PatchTST https://github.com/yuqinie98/PatchTST
-FITS https://anonymous.4open.science/r/FITS/README.md
-TimesNet https://github.com/thuml/TimesNet
-TIDE https://github.com/google-research/google-research/tree/master/tide
-Dlinear https://github.com/cure-lab/LTSF-Linear
-Appendix B Zero shot and inference efficiency
-To explore the inference efficiency and zero-shot capability of ROSE, we selected recent models with zero-shot capability, including Timer [2] and MOIRAI [8], as well as models with the potential for zero-shot learning, including PatchTST [6] and GPT4TS [10], for comparison on the ETT dataset.
-And we evaluate their inference efficiency by recording inference time per batch on the ETTh2 dataset with a batch size of 32.
-Pre-train: We select the models with the largest parameter weights for Timer and MOIRAI as baselines.
-GPT4TS and PatchTST were pre-trained on our pre-training dataset with prediction tasks to acquire zero-shot capability.
-The specific settings are shown in the table 6.
-Detailed results: For fare comparison, we set the input length as 96, 336, 512, and 720 for all zero-shot baseline models.
-We do not consider longer input length, as many real-world scenarios could offer very limited samples, and thus the zero-shot setting with only a few samples is crucial for them.
-As the baselines do not contain results with this setting in their original papers, we rerun them and report their best results in Table 11.
-ROSE consistently surpasses the baselines despite having significantly fewer parameters and inference time.
-Table 6: Pre-train setting and inference efficiency
-Model Parameters Pre-train data size Inference time per batch code repositories
-ROSE 6.4M 0.89B 4.3 ms -
-MOIRAI_large 311M 27B 90.3ms https://github.com/redoules/moirai
-GPT4TS 81.9M 0.89B 45.6 ms https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All
-Timer 67.4M 1B 68 ms https://github.com/thuml/Large-Time-Series-Model
-PatchTST 5.9M 0.89B 4.2 ms https://github.com/PatchTST/PatchTST
-Appendix C Sensitivity
+A.5 Sensitivity
+We perform the sensitivity analyses for the upper bound $a$ of the thresholds, the number of masked series $K_{\mathrm{f}}$, the number of register tokens $N_{\text{r}}$, the size of register $H$ and the number of selections $k$ in Top-K strategy.
+All the sensitivity experiments present the average results on the four ETT datasets: ETTh1, ETTh2, ETTm1 and ETTm2 under 10% few-shot setting.
 Number of masked series.
@@
 This allows the model to understand the time series from multiple frequency perspectives.
-In this experiment, we study the influence of the number of masked series $K_{m}athrm{f}$ on downstream performance.
+In this experiment, we study the influence of the number of masked series $K_{\mathrm{f}}$ on downstream performance.
 We train ROSE with 1, 2, 3, 4, 5, or 6 mask series.
@@
 This could be due to an excessive number of masked sequences leading to information redundancy.
-In all our experiments, we kept 4 mask series.
+In all our experiments, we keep 4 mask series.
 Number of register tokens.
@@
 This finding underscores the significance of utilizing a low-rank matrix to supplement the register tokens with downstream data-specific information.
-Figure 8: (a): Analysis of the number of masked series. (b): Analysis of the number of register tokens
+Figure 8: (a): Analysis of the number of masked series. (b): Analysis of the number of register tokens.
 Thresholds upper bound.
@@
 Based on our findings, we recommend using $L/5$ as the upper bound for the threshold.
-Register codebook size.
-Figure 9(b) illustrates the relationship between register codebook size and model performance.
-The codebook size determines the upper limit of domain-specific information that the register can store.
-We can observe that there is a significant improvement in the model effect when the codebook size is increased from 32 to 128.
-When the codebook size exceeds 128, the improvement of the model effect with the increase of codebook size is no longer obvious.
-Therefore, we believe that 128 is an appropriate codebook size for the current pre-training datasets.
-Figure 9: (a): Analysis of the threshold upper bound. (b): Analysis of the size of codebook
-Appendix D Model generality
-We evaluate the effectiveness of our proposed Multiple Frequency Mask (MFM) on Transformer-based models and CNN-based models, whose results are shown in Table 7.
-It is notable that MFM consistently improves these forecasting models.
-Specifically, it achieves average improvements of 6.3%, 3.7%, 1.5% in Autoformer [19], TimesNet [18], and PatchTST [6], respectively.
-This indicates that MFM can be widely utilized across various time series forecasting models to learn generalized time series representations and improve prediction accuracy.
-Table 7: Performance Multiple Frequency Masking.
+Register size.
+Figure 9(b) illustrates the relationship between register size and model performance.
+The register size determines the upper limit of domain-specific information that the register can store.
+We can observe that there is a significant improvement in the model effect when the register size is increased from 32 to 128.
+When the register size exceeds 128, the improvement of the model effect with the increase of register size is no longer obvious.
+Therefore, we believe that 128 is an appropriate register size for the current pre-training datasets.
+Number of selections in Top-K strategy.
+Figure 9(c) illustrates the relationship between the number of selections $k$ in Top-K strategy and model performance when we use the register to realize adaptive transfer of domain-specific information in downstream tasks.
+It can be seen that the model effect performance peaks at 3 tokens at $k$ = 3, which has some advantages over selecting once ($k$ = 1), indicating that the TopK strategy can compensate for the problem of incomplete matching of upstream and downstream domains to some extent.
+However, too large $k$ will also introduce redundant information and limit the accuracy of domain-specific information transfer.
+Figure 9: (a): Analysis of the threshold upper bound. (b): Analysis of the size of register. (c): Analysis of the number of selections in Top-K strategy.
+A.6 Short-term forecasting
+We also try to apply ROSE to short-term forecasting on the M4 [Makridakis, 2018] dataset, which contains the yearly, quarterly and monthly collected univariate marketing data.
+We follow TimesNet’s [Wu et al., 2022] setting and metrics (SMAPE, MASE and OWA) for testing.
+As shown in Table 9, ROSE also exhibits competitive performance on the M4 dataset compared to the baselines.
+Table 9: Full results on Short-term forecasting.
+Metric ROSE iTransformer PatchTST Timesnet Dlinear GPT4TS S2IP-LLM
+SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA SMAPE MASE OWA
+Yearly 13.302 3.014 0.833 13.238 2.952 0.823 16.766 4.331 1.018 13.387 2.996 0.786 16.965 4.283 1.058 13.531 3.015 0.793 13.413 3.024 0.792
+Quarterly 9.998 1.165 0.885 10.001 1.278 0.949 12.132 1.513 0.966 10.100 1.182 0.890 12.145 1.520 1.106 10.100 1.194 0.898 10.352 1.228 0.922
+Monthly 12.650 0.915 0.866 13.399 1.031 0.949 13.428 0.997 0.948 12.670 0.933 0.933 13.514 1.037 0.956 12.894 0.956 0.897 12.995 0.970 0.910
+Others 4.668 3.126 1.020 6.558 4.511 1.401 6.667 4.834 1.417 4.891 3.302 1.035 6.709 4.953 1.487 4.940 3.228 1.029 4.805 3.247 1.071
+A.7 Model generality
+We evaluate the effectiveness of our proposed multi-frequency masking on Transformer-based models and CNN-based models, whose results are shown in Table 10.
+It is notable that multi-frequency masking consistently improves these forecasting models.
+Specifically, it achieves average improvements of 6.3%, 3.7%, 1.5% in Autoformer [Wu et al., 2021], TimesNet [Wu et al., 2022], and PatchTST [Nie et al., 2022], respectively.
+This indicates that multi-frequency Masking can be widely utilized across various time series forecasting models to learn generalized time series representations and improve prediction accuracy.
+Table 10: Performance of multi-frequency masking.
 Datasets ETTm1 ETTm2 ETTh1 ETTh2
 Metric MSE MAE MSE MAE MSE MAE MSE MAE
-Autoformer +Multi Freq Mask 0.600 0.521 0.328 0.365 0.493 0.487 0.452 0.458
+Autoformer +Multi-frequency Masking 0.600 0.521 0.328 0.365 0.493 0.487 0.452 0.458
 0.549 0.488 0.306 0.349 0.474 0.478 0.406 0.425
-TimesNet +Multi Freq Mask 0.4 0.406 0.291 0.333 0.458 0.45 0.414 0.427
+TimesNet +Multi-frequency Masking 0.400 0.406 0.291 0.333 0.458 0.450 0.414 0.427
 0.386 0.398 0.282 0.324 0.446 0.438 0.386 0.403
-PatchTST +Multi Freq Mask 0.353 0.382 0.256 0.317 0.413 0.434 0.331 0.381
+PatchTST +Multi-frequency Masking 0.353 0.382 0.256 0.317 0.413 0.434 0.331 0.381
 0.347 0.372 0.252 0.308 0.405 0.424 0.337 0.379
-Appendix E Visualization of forecasting results
-To provide a distinct comparison among different models, we present visualizations of the forecasting results on the ETTh2 dataset and the weather dataset in different settings, as shown in Figures 10 to Figures 13, given by the following models: DLinear [40], TimesNet [18], iTransfomrer [20], PatchTST [6], and TiDE [39].
+A.8 Results deviation
+We have conducted ROSE three times with different random seeds and have recorded the standard deviations for both the full-shot setting and the 10% few-shot setting, as illustrated in Table 11.
+As the baselines didn’t report deviations in the original paper, we only reported the deviations of the PatchTST in the full-shot setting as a comparison.
+It can be observed that ROSE exhibits stable performance.
+Table 11: Results deviation.
+Models ROSE ROSE (10%) PatchTST confidence interval
+Metric MSE MAE MSE MAE MSE MAE -
+ETTm1 0.342$\pm$0.003 0.367$\pm$0.002 0.349$\pm$0.003 0.372$\pm$0.002 0.349$\pm$ 0.004 0.383$\pm$0.003 99%
+ETTm2 0.246$\pm$0.002 0.303$\pm$0.004 0.249$\pm$0.002 0.308$\pm$0.002 0.255$\pm$0.002 0.314$\pm$0.003 99%
+ETTh1 0.392$\pm$0.004 0.413$\pm$0.004 0.397$\pm$0.003 0.419$\pm$0.003 0.411$\pm$0.003 0.432$\pm$0.005 99%
+ETTh2 0.330$\pm$0.003 0.374$\pm$0.002 0.335$\pm$0.004 0.380$\pm$0.003 0.348$\pm$0.004 0.390$\pm$0.004 99%
+Traffic 0.391$\pm$0.008 0.266$\pm$0.005 0.418$\pm$0.011 0.278$\pm$0.006 0.404$\pm$0.009 0.283$\pm$0.002 99%
+Weather 0.217$\pm$0.008 0.250$\pm$0.007 0.224$\pm$0.007 0.252$\pm$0.009 0.223$\pm$0.011 0.263$\pm$0.014 99%
+Electricity 0.156$\pm$0.007 0.249$\pm$0.009 0.164$\pm$0.004 0.253$\pm$0.004 0.163$\pm$0.009 0.261$\pm$0.013 99%
+A.9 Visualization
+A.9.1 Visualization analysis
+To showcase the benefits of cross-domain pre-training, we performed visualizations in both the zero-shot setting and full-shot setting.
+Zero-shot: We pre-train the baselines iTransformer and PatchTST on the energy domain dataset ETTm1 and test their zero-shot performance on two different domains (weather, traffic) .
+ROSE, without fine-tuning, is evaluated to the same two test-sets.
+As shown in the Figure 10, we find that the baselines generally perform worse during domain shift due to their poor generalization.
+However, ROSE excels in scenarios across all domains, which demonstrates the benefits of cross-domain pre-training for improving generalization.
+Full-shot: We train the baselines on the train-set of downstream dataset ETTh2 and fine-tune ROSE on the same train-set.
+As shown in the Figure 11, We find that the baselines is limited by data diversity, leading to poor performance on patterns which rarely appear.
+However, ROSE excels in these cases, as the cross-domain pre-training allows ROSE to learn diverse temporal patterns, and helps ROSE to predict the patterns which rarely appear in the downstream train-set well.
+Figure 10: Visualization comparison of ROSE with cross-domain pre-training and other SOTA baselines in the zero-shot setting for three domain datasets.
+Figure 11: Visualization comparison of ROSE with cross-domain pre-training and other SOTA baselines in the full-shot setting for rare and common patterns.
+A.9.2 Visualization showcase
+To provide a distinct comparison among different models, we present visualizations of the forecasting results on the ETTh2 dataset and the weather dataset in different settings, as shown in Figures 12 to Figures 15, given by the following models: DLinear [Zeng et al., 2023], TimesNet [Wu et al., 2022], iTransfomrer [Liu et al., 2023], and PatchTST [Nie et al., 2022].
 Among the methods, ROSE demonstrates the most accurate prediction ability.
-Figure 10: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset with full-sample.
-Figure 11: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset with 10% samples.
-Figure 12: Visualization of input-512 and predict-336 forecasting results on the weather dataset with full-sample.
-Figure 13: Visualization of input-512 and predict-336 forecasting results on the weather dataset with 10%-sample.
-Appendix F Limitations
-Our pre-training datasets still have room for expansion.
-While ROSE demonstrates strong performance, its effectiveness could be enhanced with a larger and more diverse set of pre-training datasets.
-By collecting additional real-world scenario data for pre-training, we aim to further improve the robustness and generalizability of ROSE across various domains.
-Appendix G Broader Impacts
-This paper presents ROSE, a novel general time series forecasting model that demonstrates promising performance.
-ROSE performs well in few-shot scenarios and shows notable zero-shot capabilities, highlighting its scalability and efficiency.
-We believe it can be a valuable addition to the pre-training research community.
-To support future research, we will also release the codebase for time-series pre-training.
-This paper only focuses on the algorithm design. Using all the codes and datasets strictly follows the corresponding licenses A.1. There is no potential ethical risk or negative social impact.
-Appendix H Full Results
-H.1 Full-shot results
+Figure 12: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset in full-shot setting.
+Figure 13: Visualization of input-512 and predict-336 forecasting results on the ETTh2 dataset in 10% few-shot setting.
+Figure 14: Visualization of input-512 and predict-336 forecasting results on the weather dataset in full-shot setting.
+Figure 15: Visualization of input-512 and predict-336 forecasting results on the weather dataset in 10% few-shot setting.
+A.10 Full Results
+A.10.1 Full-shot results
 Table 12: Full results in full-shot setting.
-Models ROSE Timer ITransformer PatchTST FITS Timesnet TiDE Dlinear
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1 96 0.275 0.328 0.339 0.383 0.334 0.368 0.293 0.346 0.303 0.352 0.338 0.375 0.306 0.349 0.299 0.343
-192 0.324 0.358 0.396 0.430 0.377 0.391 0.333 0.370 0.337 0.368 0.374 0.387 0.335 0.366 0.335 0.365
-336 0.354 0.377 0.451 0.471 0.426 0.420 0.369 0.392 0.366 0.385 0.410 0.411 0.364 0.384 0.369 0.386
-720 0.411 0.407 0.558 0.540 0.491 0.459 0.416 0.420 0.415 0.411 0.478 0.450 0.413 0.413 0.425 0.421
-avg 0.341 0.367 0.436 0.456 0.407 0.410 0.353 0.382 0.355 0.379 0.400 0.406 0.355 0.378 0.357 0.379
-ETTm2 96 0.157 0.243 0.184 0.263 0.180 0.264 0.166 0.256 0.162 0.253 0.187 0.267 0.161 0.251 0.167 0.260
-192 0.213 0.283 0.278 0.334 0.250 0.309 0.223 0.296 0.216 0.291 0.249 0.309 0.215 0.289 0.224 0.303
-336 0.266 0.319 0.334 0.375 0.311 0.348 0.274 0.329 0.268 0.326 0.321 0.351 0.267 0.326 0.281 0.342
-720 0.347 0.373 0.435 0.431 0.412 0.407 0.362 0.385 0.348 0.378 0.408 0.403 0.352 0.383 0.397 0.421
-avg 0.246 0.305 0.308 0.351 0.288 0.332 0.256 0.317 0.249 0.312 0.291 0.333 0.249 0.312 0.267 0.332
-ETTh1 96 0.354 0.385 0.358 0.345 0.386 0.405 0.370 0.400 0.372 0.402 0.384 0.402 0.375 0.398 0.375 0.399
-192 0.389 0.407 0.438 0.448 0.441 0.436 0.413 0.429 0.404 0.422 0.436 0.429 0.412 0.422 0.405 0.416
-336 0.406 0.422 0.491 0.484 0.487 0.458 0.422 0.440 0.427 0.438 0.491 0.469 0.435 0.433 0.439 0.443
-720 0.413 0.443 0.644 0.559 0.503 0.491 0.447 0.468 0.424 0.454 0.521 0.500 0.454 0.465 0.472 0.490
-avg 0.391 0.414 0.483 0.459 0.454 0.447 0.413 0.434 0.407 0.429 0.458 0.450 0.419 0.430 0.423 0.437
-ETTh2 96 0.265 0.320 0.288 0.346 0.297 0.349 0.274 0.337 0.271 0.336 0.340 0.374 0.270 0.336 0.289 0.353
-192 0.328 0.369 0.380 0.409 0.380 0.400 0.341 0.382 0.331 0.374 0.402 0.414 0.332 0.380 0.383 0.418
-336 0.353 0.391 0.372 0.413 0.428 0.432 0.329 0.384 0.354 0.395 0.452 0.452 0.360 0.407 0.448 0.465
-720 0.376 0.417 0.425 0.456 0.427 0.445 0.379 0.422 0.377 0.423 0.462 0.468 0.419 0.451 0.605 0.551
-avg 0.330 0.374 0.366 0.406 0.383 0.407 0.331 0.381 0.333 0.382 0.414 0.427 0.345 0.394 0.431 0.447
-Traffic 96 0.354 0.252 0.351 0.255 0.395 0.268 0.360 0.249 0.385 0.271 0.593 0.321 0.336 0.253 0.410 0.282
-192 0.377 0.257 0.387 0.269 0.417 0.276 0.379 0.256 0.397 0.275 0.617 0.336 0.346 0.257 0.423 0.287
-336 0.396 0.262 0.426 0.288 0.433 0.283 0.392 0.264 0.410 0.280 0.629 0.336 0.355 0.260 0.436 0.296
-720 0.434 0.283 0.489 0.323 0.467 0.302 0.432 0.286 0.448 0.301 0.640 0.350 0.386 0.273 0.466 0.315
-avg 0.390 0.264 0.413 0.284 0.428 0.282 0.391 0.264 0.410 0.282 0.620 0.336 0.356 0.261 0.434 0.295
-Weather 96 0.145 0.182 0.154 0.208 0.174 0.214 0.149 0.198 0.143 0.195 0.172 0.220 0.166 0.222 0.176 0.237
-192 0.183 0.226 0.243 0.287 0.221 0.254 0.194 0.241 0.186 0.238 0.219 0.261 0.209 0.263 0.220 0.282
-336 0.232 0.267 0.329 0.345 0.278 0.296 0.245 0.282 0.236 0.277 0.280 0.306 0.254 0.301 0.265 0.319
-720 0.309 0.327 0.384 0.384 0.358 0.347 0.314 0.334 0.307 0.330 0.365 0.359 0.313 0.340 0.323 0.362
-avg 0.217 0.251 0.277 0.306 0.258 0.278 0.226 0.264 0.218 0.260 0.259 0.287 0.236 0.282 0.246 0.300
-Solar 96 0.171 0.199 0.230 0.245 0.203 0.237 0.190 0.273 0.198 0.253 0.250 0.292 0.200 0.270 0.206 0.281
-192 0.185 0.215 0.249 0.261 0.233 0.261 0.204 0.302 0.218 0.262 0.296 0.318 0.219 0.285 0.225 0.291
-336 0.193 0.218 0.255 0.265 0.248 0.273 0.212 0.293 0.235 0.270 0.319 0.330 0.234 0.296 0.240 0.300
-720 0.198 0.224 0.272 0.276 0.249 0.275 0.221 0.310 0.241 0.273 0.338 0.337 0.242 0.303 0.248 0.307
-avg 0.187 0.214 0.251 0.262 0.233 0.262 0.207 0.294 0.223 0.264 0.301 0.319 0.223 0.288 0.230 0.295
-Electricity 96 0.125 0.220 0.136 0.230 0.148 0.240 0.129 0.222 0.134 0.231 0.168 0.272 0.132 0.229 0.140 0.237
-192 0.142 0.235 0.168 0.260 0.162 0.253 0.147 0.240 0.149 0.244 0.184 0.289 0.147 0.243 0.153 0.249
-336 0.162 0.252 0.192 0.285 0.178 0.269 0.163 0.259 0.165 0.260 0.198 0.300 0.161 0.261 0.169 0.267
-720 0.191 0.284 0.230 0.322 0.225 0.317 0.197 0.290 0.203 0.293 0.220 0.320 0.196 0.294 0.203 0.301
-avg 0.155 0.248 0.181 0.274 0.178 0.270 0.159 0.253 0.163 0.257 0.193 0.295 0.159 0.257 0.166 0.264
-H.2 Few-shot results
+Models ROSE ITransformer PatchTST Timesnet Dlinear GPT4TS S2P-LLM
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 96 0.354 0.385 0.386 0.405 0.370 0.400 0.470 0.470 0.367 0.396 0.376 0.397 0.366 0.396
+192 0.389 0.407 0.424 0.440 0.413 0.429 0.568 0.523 0.400 0.417 0.416 0.418 0.401 0.420
+336 0.406 0.422 0.449 0.460 0.422 0.440 0.595 0.547 0.428 0.439 0.442 0.433 0.412 0.431
+720 0.413 0.443 0.495 0.487 0.447 0.468 0.694 0.591 0.468 0.491 0.477 0.456 0.440 0.458
+avg 0.391 0.414 0.439 0.448 0.413 0.434 0.582 0.533 0.416 0.436 0.427 0.426 0.406 0.427
+ETTh2 96 0.265 0.320 0.297 0.348 0.274 0.337 0.351 0.399 0.302 0.368 0.285 0.342 0.278 0.340
+192 0.328 0.369 0.371 0.403 0.341 0.382 0.394 0.429 0.404 0.433 0.354 0.389 0.246 0.385
+336 0.353 0.391 0.404 0.428 0.329 0.384 0.415 0.443 0.511 0.498 0.373 0.407 0.367 0.406
+720 0.376 0.417 0.424 0.444 0.379 0.422 0.477 0.481 0.815 0.640 0.406 0.441 0.400 0.436
+avg 0.331 0.374 0.374 0.406 0.331 0.381 0.409 0.438 0.508 0.485 0.354 0.394 0.347 0.391
+ETTm1 96 0.275 0.328 0.300 0.353 0.293 0.346 0.405 0.421 0.303 0.346 0.292 0.262 0.288 0.346
+192 0.324 0.358 0.345 0.382 0.333 0.370 0.508 0.473 0.335 0.365 0.332 0.301 0.323 0.365
+336 0.354 0.377 0.374 0.398 0.369 0.392 0.523 0.479 0.365 0.384 0.366 0.341 0.359 0.390
+720 0.411 0.407 0.429 0.430 0.416 0.420 0.523 0.484 0.418 0.415 0.417 0.401 0.403 0.418
+avg 0.341 0.367 0.362 0.391 0.353 0.382 0.490 0.464 0.356 0.378 0.352 0.383 0.343 0.379
+ETTm2 96 0.157 0.243 0.175 0.266 0.166 0.256 0.233 0.305 0.164 0.255 0.173 0.262 0.165 0.257
+192 0.213 0.283 0.242 0.312 0.223 0.296 0.265 0.328 0.224 0.304 0.229 0.301 0.222 0.299
+336 0.266 0.319 0.282 0.340 0.274 0.329 0.379 0.392 0.277 0.339 0.286 0.341 0.277 0.330
+720 0.347 0.373 0.378 0.398 0.362 0.385 0.390 0.407 0.371 0.401 0.378 0.401 0.363 0.390
+avg 0.246 0.305 0.269 0.329 0.256 0.317 0.317 0.358 0.259 0.325 0.266 0.326 0.257 0.319
+Weather 96 0.145 0.182 0.159 0.208 0.149 0.198 0.193 0.244 0.170 0.230 0.162 0.212 0.145 0.195
+192 0.183 0.226 0.200 0.248 0.194 0.241 0.320 0.329 0.212 0.267 0.204 0.248 0.190 0.235
+336 0.232 0.267 0.253 0.289 0.245 0.282 0.363 0.366 0.257 0.305 0.254 0.286 0.243 0.280
+720 0.309 0.327 0.321 0.338 0.314 0.334 0.440 0.404 0.318 0.356 0.326 0.337 0.312 0.326
+avg 0.217 0.251 0.233 0.271 0.226 0.264 0.329 0.336 0.239 0.289 0.237 0.270 0.222 0.259
+Electricity 96 0.125 0.220 0.138 0.237 0.129 0.222 0.182 0.287 0.141 0.241 0.139 0.238 0.135 0.230
+192 0.142 0.235 0.157 0.256 0.147 0.240 0.193 0.293 0.154 0.254 0.153 0.251 0.149 0.247
+336 0.162 0.252 0.167 0.264 0.163 0.259 0.196 0.298 0.168 0.271 0.169 0.266 0.167 0.266
+720 0.191 0.284 0.194 0.286 0.197 0.290 0.209 0.307 0.203 0.303 0.206 0.297 0.200 0.287
+avg 0.155 0.248 0.164 0.261 0.159 0.253 0.195 0.296 0.166 0.267 0.167 0.263 0.161 0.257
+Traffic 96 0.354 0.252 0.363 0.265 0.360 0.249 0.611 0.323 0.411 0.294 0.388 0.282 0.379 0.274
+192 0.377 0.257 0.385 0.273 0.379 0.256 0.609 0.327 0.421 0.298 0.407 0.290 0.397 0.282
+336 0.396 0.262 0.396 0.277 0.392 0.264 0.616 0.335 0.431 0.304 0.412 0.294 0.407 0.289
+720 0.434 0.283 0.445 0.312 0.432 0.286 0.656 0.349 0.468 0.325 0.450 0.312 0.440 0.301
+avg 0.390 0.264 0.397 0.282 0.391 0.264 0.623 0.333 0.433 0.305 0.414 0.294 0.405 0.286
+A.10.2 Few-shot results
 Table 13: Full results in 10% few-shot setting
-Models ROSE Timer ITransformer PatchTST FITS Timesnet TiDE Dlinear
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1 96 0.287 0.336 0.351 0.398 0.353 0.392 0.317 0.363 0.311 0.354 0.481 0.446 0.373 0.388 0.454 0.475
-192 0.331 0.362 0.429 0.448 0.385 0.410 0.351 0.382 0.351 0.378 0.621 0.491 0.406 0.403 0.575 0.548
-336 0.362 0.379 0.496 0.495 0.422 0.432 0.376 0.398 0.377 0.396 0.521 0.479 0.436 0.425 0.773 0.631
-720 0.416 0.412 0.573 0.573 0.494 0.472 0.435 0.430 0.441 0.436 0.571 0.508 0.499 0.457 0.943 0.716
-avg 0.349 0.372 0.462 0.478 0.414 0.426 0.370 0.393 0.370 0.391 0.549 0.481 0.429 0.418 0.686 0.593
-ETTm2 96 0.159 0.247 0.196 0.273 0.183 0.279 0.170 0.259 0.166 0.258 0.212 0.292 0.195 0.280 0.493 0.476
-192 0.217 0.287 0.290 0.338 0.247 0.320 0.226 0.297 0.222 0.295 0.297 0.353 0.257 0.317 0.923 0.658
-336 0.268 0.322 0.350 0.386 0.3 0.353 0.284 0.333 0.273 0.328 0.328 0.364 0.320 0.354 1.407 0.822
-720 0.355 0.377 0.436 0.435 0.385 0.408 0.363 0.382 0.370 0.385 0.456 0.440 0.421 0.410 1.626 0.905
-avg 0.249 0.308 0.318 0.358 0.279 0.340 0.261 0.318 0.258 0.317 0.323 0.362 0.298 0.340 1.112 0.715
-ETTh1 96 0.367 0.395 0.361 0.409 0.442 0.464 0.458 0.463 0.438 0.449 0.579 0.522 0.708 0.569 1.355 0.816
-192 0.399 0.416 0.398 0.427 0.476 0.475 0.481 0.490 0.512 0.489 0.641 0.553 0.744 0.584 1.210 0.825
-336 0.405 0.423 0.405 0.436 0.486 0.482 0.465 0.475 0.584 0.532 0.721 0.582 0.768 0.599 1.487 0.914
-720 0.416 0.443 0.492 0.503 0.509 0.506 0.478 0.492 0.681 0.590 0.630 0.574 0.770 0.606 1.369 0.826
-avg 0.397 0.419 0.414 0.444 0.478 0.482 0.470 0.480 0.554 0.515 0.643 0.558 0.748 0.589 1.355 0.845
-ETTh2 96 0.273 0.332 0.298 0.356 0.333 0.385 0.350 0.389 0.297 0.360 0.378 0.413 0.434 0.432 1.628 0.724
-192 0.334 0.376 0.374 0.404 0.402 0.428 0.416 0.426 0.363 0.399 0.463 0.460 0.528 0.480 1.388 0.713
-336 0.358 0.397 0.385 0.425 0.438 0.452 0.401 0.429 0.388 0.424 0.507 0.495 0.473 0.457 1.595 0.772
-720 0.376 0.417 0.468 0.487 0.466 0.477 0.436 0.457 0.423 0.456 0.516 0.501 0.567 0.513 1.664 0.857
-avg 0.335 0.380 0.381 0.418 0.410 0.436 0.401 0.425 0.368 0.410 0.466 0.467 0.501 0.471 1.569 0.766
-Traffic 96 0.398 0.270 0.353 0.277 0.458 0.314 0.421 0.299 0.406 0.301 0.705 0.386 0.681 0.412 0.616 0.385
-192 0.405 0.270 0.405 0.288 0.473 0.319 0.439 0.313 0.523 0.424 0.710 0.393 0.624 0.385 0.710 0.480
-336 0.417 0.277 0.428 0.300 0.491 0.329 0.448 0.318 0.624 0.479 0.863 0.456 0.613 0.368 0.723 0.481
-720 0.452 0.294 0.487 0.332 0.536 0.361 0.478 0.320 0.937 0.629 0.928 0.485 0.653 0.387 0.673 0.436
-avg 0.418 0.278 0.418 0.299 0.490 0.331 0.447 0.312 0.623 0.458 0.801 0.430 0.643 0.388 0.680 0.446
-Weather 96 0.145 0.184 0.180 0.228 0.189 0.229 0.166 0.217 0.174 0.222 0.199 0.248 0.208 0.240 0.230 0.318
-192 0.190 0.227 0.256 0.296 0.239 0.269 0.211 0.257 0.226 0.265 0.249 0.285 0.249 0.274 0.357 0.425
-336 0.245 0.269 0.321 0.344 0.294 0.308 0.261 0.296 0.279 0.301 0.297 0.316 0.283 0.306 0.464 0.493
-720 0.317 0.328 0.420 0.398 0.366 0.356 0.328 0.342 0.361 0.354 0.367 0.361 0.365 0.354 0.515 0.532
-avg 0.224 0.252 0.294 0.317 0.272 0.291 0.242 0.278 0.260 0.285 0.278 0.303 0.276 0.293 0.391 0.442
-Solar 96 0.177 0.203 0.230 0.284 0.292 0.310 0.202 0.245 0.242 0.306 0.306 0.323 0.359 0.382 0.748 0.757
-192 0.185 0.214 0.246 0.300 0.313 0.321 0.222 0.256 0.262 0.315 0.356 0.354 0.395 0.400 0.752 0.759
-336 0.193 0.218 0.278 0.319 0.332 0.335 0.252 0.272 0.276 0.321 0.387 0.362 0.429 0.410 0.754 0.760
-720 0.208 0.225 0.255 0.300 0.355 0.338 0.264 0.273 0.280 0.329 0.396 0.361 0.410 0.378 0.755 0.761
-avg 0.191 0.215 0.252 0.301 0.323 0.326 0.235 0.262 0.265 0.318 0.361 0.350 0.398 0.392 0.752 0.759
-Electricity 96 0.135 0.226 0.156 0.254 0.184 0.276 0.161 0.256 0.144 0.242 0.279 0.359 0.271 0.359 0.227 0.334
-192 0.150 0.240 0.254 0.275 0.192 0.284 0.163 0.257 0.174 0.283 0.282 0.363 0.216 0.308 0.265 0.366
-336 0.166 0.258 0.212 0.306 0.216 0.308 0.173 0.266 0.225 0.341 0.289 0.367 0.237 0.324 0.339 0.417
-720 0.205 0.290 0.271 0.352 0.265 0.347 0.221 0.313 0.408 0.495 0.333 0.399 0.258 0.330 0.482 0.478
-avg 0.164 0.253 0.223 0.297 0.214 0.304 0.180 0.273 0.238 0.341 0.296 0.372 0.245 0.330 0.328 0.399
-H.3 Ablation study results
+Models ROSE ITransformer PatchTST Timesnet Dlinear GPT4TS S2P-LLM
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 96 0.367 0.395 0.442 0.464 0.458 0.463 0.579 0.522 1.355 0.816 0.458 0.456 0.481 0.474
+192 0.399 0.416 0.476 0.475 0.481 0.490 0.641 0.553 1.210 0.825 0.570 0.516 0.518 0.491
+336 0.405 0.423 0.486 0.482 0.465 0.475 0.721 0.582 1.487 0.914 0.608 0.535 0.664 0.570
+720 0.416 0.443 0.509 0.506 0.478 0.492 0.630 0.574 1.369 0.826 0.725 0.591 0.711 0.584
+avg 0.397 0.419 0.478 0.482 0.470 0.480 0.643 0.558 1.355 0.845 0.590 0.525 0.593 0.529
+ETTh2 96 0.273 0.332 0.333 0.385 0.350 0.389 0.378 0.413 1.628 0.724 0.331 0.374 0.354 0.400
+192 0.334 0.376 0.402 0.428 0.416 0.426 0.463 0.460 1.388 0.713 0.402 0.411 0.400 0.423
+336 0.358 0.397 0.438 0.452 0.401 0.429 0.507 0.495 1.595 0.772 0.406 0.433 0.442 0.450
+720 0.376 0.417 0.466 0.477 0.436 0.457 0.516 0.501 1.664 0.857 0.449 0.464 0.480 0.486
+avg 0.335 0.380 0.410 0.436 0.401 0.425 0.466 0.467 1.569 0.766 0.397 0.421 0.419 0.439
+ETTm1 96 0.287 0.336 0.353 0.392 0.317 0.363 0.481 0.446 0.454 0.475 0.390 0.404 0.388 0.401
+192 0.331 0.362 0.385 0.410 0.351 0.382 0.621 0.491 0.575 0.548 0.429 0.423 0.422 0.421
+336 0.362 0.379 0.422 0.432 0.376 0.398 0.521 0.479 0.773 0.631 0.469 0.439 0.456 0.430
+720 0.416 0.412 0.494 0.472 0.435 0.430 0.571 0.508 0.943 0.716 0.569 0.498 0.554 0.490
+avg 0.349 0.372 0.414 0.426 0.370 0.393 0.549 0.481 0.686 0.593 0.464 0.441 0.455 0.435
+ETTm2 96 0.159 0.247 0.183 0.279 0.170 0.259 0.212 0.292 0.493 0.476 0.188 0.269 0.192 0.274
+192 0.217 0.287 0.247 0.320 0.226 0.297 0.297 0.353 0.923 0.658 0.251 0.309 0.246 0.313
+336 0.269 0.322 0.300 0.353 0.284 0.333 0.328 0.364 1.407 0.822 0.307 0.346 0.301 0.340
+720 0.357 0.377 0.385 0.408 0.363 0.382 0.456 0.440 1.626 0.905 0.426 0.417 0.400 0.403
+avg 0.250 0.308 0.279 0.340 0.261 0.318 0.323 0.362 1.112 0.715 0.293 0.335 0.284 0.332
+Weather 96 0.145 0.184 0.189 0.229 0.166 0.217 0.199 0.248 0.230 0.318 0.163 0.215 0.159 0.210
+192 0.190 0.227 0.239 0.269 0.211 0.257 0.249 0.285 0.357 0.425 0.210 0.254 0.200 0.251
+336 0.245 0.269 0.294 0.308 0.261 0.296 0.297 0.316 0.464 0.493 0.256 0.292 0.257 0.293
+720 0.317 0.328 0.366 0.356 0.328 0.342 0.367 0.361 0.515 0.532 0.321 0.339 0.317 0.335
+avg 0.224 0.252 0.272 0.291 0.242 0.278 0.278 0.303 0.391 0.442 0.238 0.275 0.233 0.272
+Electricity 96 0.135 0.226 0.184 0.276 0.161 0.256 0.279 0.359 0.227 0.334 0.139 0.237 0.143 0.243
+192 0.150 0.240 0.192 0.284 0.163 0.257 0.282 0.363 0.265 0.366 0.156 0.252 0.159 0.258
+336 0.166 0.258 0.216 0.308 0.173 0.266 0.289 0.367 0.339 0.417 0.175 0.270 0.170 0.269
+720 0.205 0.290 0.265 0.347 0.221 0.313 0.333 0.399 0.482 0.478 0.233 0.317 0.230 0.315
+avg 0.164 0.253 0.214 0.304 0.180 0.273 0.296 0.372 0.328 0.399 0.176 0.269 0.175 0.271
+Traffic 96 0.398 0.270 0.458 0.314 0.421 0.299 0.705 0.386 0.616 0.385 0.414 0.297 0.403 0.293
+192 0.405 0.270 0.473 0.319 0.439 0.313 0.710 0.393 0.710 0.480 0.426 0.301 0.412 0.295
+336 0.417 0.277 0.491 0.329 0.448 0.318 0.863 0.456 0.723 0.481 0.434 0.303 0.427 0.316
+720 0.452 0.294 0.536 0.361 0.478 0.320 0.928 0.485 0.673 0.436 0.487 0.337 0.469 0.325
+avg 0.418 0.278 0.490 0.331 0.447 0.312 0.801 0.430 0.680 0.446 0.440 0.310 0.427 0.307
+A.10.3 Ablation study results
+Novelty of decomposed frequency learning.
+Frequency masking is not a new concept, but past approaches randomly mask frequencies of a single time series once  [Chen et al., 2023b, Zhang et al., 2023], which show limited forecasting effectiveness due to the lack of common pattern learning from heterogeneous time series that come from various domains.
+While the multi-frequency masking we proposed randomly mask either high-frequency or low-frequency components of a time series multiple times as the key to enable learning of common time series patterns, such as trends and various long and short term fluctuations.
+Moreover, different from utilizing frequency masking as a way of data augmentation to enhance the diversity of input data  [Chen et al., 2023b, Zhang et al., 2023], we combine multi-frequency masking with reconstruction task as a novel pre-training framework, that learns a universal and unified feature representation by comprehending the data from various frequency perspectives, thereby enabling it to learn generalized representations.
+Difference between frequency-domain masking and time-domain noise addition.
+Multi-frequency masking and reconstruction are not equivalent to the pre-training methods of adding noise and denoising (Noise).
+Due to the sparsity of time series, the process of adding noise and denoising may potentially disrupt the information of original time series [Dong et al., 2024].
+In contrast, multi-frequency masking not only preserves the series from such disruption but also helps the model understand temporal patterns from a multi-frequency perspective, thereby helping the model to learn general features better.
+Other pre-training tasks.
+Based on the two points above, we conduct experiments to compare two other pre-training tasks: 1) using frequency-domain augmentation only for data expansion without reconstruction task (Aug); 2) replacing multi-frequency masking and reconstruction task with adding time-domian noise and denoise task (Noise).As shown in Table 14, we find that ROSE is significantly more effective than Aug and Noise, which demonstrates the effectiveness of multi-frequency masking and reconstruction task in learning generalized features.
 Table 14: Full results of ablation study
@@
 720 0.416 0.412 0.357 0.377 0.416 0.443 0.376 0.417
-avg 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.38
-Replace Multi Freq Masking Single Freq Masking 96 0.296 0.346 0.163 0.253 0.371 0.398 0.272 0.335
-192 0.336 0.368 0.221 0.291 0.406 0.420 0.340 0.376
-336 0.362 0.385 0.276 0.327 0.411 0.426 0.379 0.406
-720 0.424 0.418 0.366 0.381 0.422 0.447 0.399 0.431
-avg 0.354 0.379 0.257 0.313 0.403 0.423 0.347 0.387
-Multi Patch Masking 96 0.302 0.348 0.168 0.257 0.377 0.408 0.282 0.343
+avg 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.380
+Replace Multi-Frequency Masking Random Frequency Masking 96 0.330 0.370 0.170 0.262 0.391 0.399 0.303 0.359
+192 0.352 0.392 0.232 0.298 0.406 0.430 0.356 0.395
+336 0.390 0.389 0.276 0.342 0.411 0.432 0.417 0.428
+720 0.452 0.438 0.366 0.392 0.432 0.447 0.420 0.439
+avg 0.381 0.397 0.261 0.324 0.410 0.427 0.374 0.405
+Multi-Patch Masking 96 0.302 0.348 0.168 0.257 0.377 0.408 0.282 0.343
 192 0.336 0.367 0.228 0.297 0.404 0.423 0.343 0.379
@@
 720 0.450 0.438 0.370 0.388 0.431 0.459 0.413 0.443
-avg 0.378 0.4 0.261 0.319 0.408 0.432 0.375 0.407
-w/o Aqp 96 0.297 0.345 0.164 0.252 0.379 0.399 0.276 0.336
+avg 0.378 0.400 0.261 0.319 0.408 0.432 0.375 0.407
+Other pre-training tasks Aug 96 0.304 0.357 0.178 0.266 0.376 0.405 0.281 0.348
+192 0.343 0.379 0.254 0.318 0.409 0.429 0.345 0.389
+336 0.373 0.400 0.299 0.354 0.435 0.453 0.382 0.417
+720 0.444 0.432 0.387 0.408 0.452 0.471 0.434 0.452
+avg 0.366 0.392 0.279 0.336 0.418 0.439 0.360 0.401
+Noise 96 0.303 0.355 0.172 0.261 0.370 0.405 0.280 0.341
+192 0.342 0.376 0.221 0.292 0.403 0.427 0.350 0.384
+336 0.368 0.393 0.272 0.325 0.420 0.439 0.385 0.410
+720 0.423 0.422 0.367 0.386 0.442 0.462 0.403 0.432
+avg 0.359 0.387 0.258 0.316 0.409 0.433 0.355 0.392
+w/o TS-Register 96 0.297 0.345 0.164 0.252 0.379 0.399 0.276 0.336
 192 0.334 0.367 0.221 0.290 0.419 0.420 0.350 0.380
@@
 720 0.424 0.416 0.364 0.379 0.435 0.448 0.400 0.432
-avg 0.354 0.378 0.256 0.312 0.418 0.427 0.355 0.39
-Predict Task 96 0.301 0.348 0.166 0.255 0.380 0.407 0.295 0.359
+avg 0.354 0.378 0.256 0.312 0.418 0.427 0.355 0.390
+Prediction Task 96 0.301 0.348 0.166 0.255 0.380 0.407 0.295 0.359
 192 0.343 0.374 0.221 0.291 0.410 0.426 0.372 0.406
@@
 720 0.424 0.420 0.366 0.384 0.458 0.476 0.418 0.446
-avg 0.36 0.384 0.257 0.314 0.422 0.438 0.372 0.41
-Reconstruct Task 96 0.329 0.371 0.175 0.265 0.374 0.399 0.296 0.355
+avg 0.360 0.384 0.257 0.314 0.422 0.438 0.372 0.410
+Reconstruction Task 96 0.329 0.371 0.175 0.265 0.374 0.399 0.296 0.355
 192 0.363 0.391 0.233 0.304 0.407 0.422 0.354 0.389
@@
 avg 0.387 0.403 0.269 0.327 0.412 0.428 0.361 0.399
-H.4 Zero-shot results
-Table 11: Full results in zero-shot setting
-Model ROSE MOIRAI Timer PatchTST GPT4TS
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1 96 0.512 0.460 0.660 0.476 0.596 0.519 0.906 0.578 0.743 0.545
-192 0.512 0.462 0.707 0.500 0.668 0.560 0.892 0.573 0.759 0.563
-336 0.523 0.470 0.730 0.515 1.580 0.871 0.876 0.567 0.748 0.570
-720 0.552 0.490 0.758 0.536 1.587 0.882 0.836 0.554 0.780 0.576
-avg 0.525 0.471 0.714 0.507 1.108 0.708 0.877 0.568 0.758 0.564
-ETTm2 96 0.224 0.309 0.216 0.282 0.221 0.300 0.301 0.346 0.275 0.343
-192 0.266 0.333 0.294 0.330 0.303 0.349 0.337 0.367 0.298 0.356
-336 0.310 0.358 0.368 0.373 0.411 0.423 0.376 0.389 0.347 0.386
-720 0.395 0.407 0.494 0.439 0.505 0.473 0.456 0.434 0.426 0.417
-avg 0.299 0.352 0.343 0.356 0.360 0.386 0.367 0.384 0.337 0.376
-ETTh1 96 0.382 0.408 0.405 0.397 0.414 0.439 0.583 0.502 0.573 0.521
-192 0.400 0.420 0.458 0.428 0.458 0.465 0.600 0.516 0.594 0.529
-336 0.404 0.426 0.509 0.454 0.475 0.479 0.594 0.525 0.682 0.567
-720 0.420 0.447 0.529 0.494 0.521 0.511 0.569 0.544 0.543 0.524
-avg 0.401 0.425 0.475 0.443 0.467 0.474 0.586 0.522 0.598 0.535
-ETTh2 96 0.298 0.362 0.303 0.338 0.313 0.366 0.401 0.418 0.369 0.411
-192 0.336 0.385 0.369 0.384 0.394 0.415 0.412 0.425 0.382 0.423
-336 0.353 0.399 0.397 0.410 0.420 0.439 0.410 0.427 0.401 0.442
-720 0.395 0.432 0.447 0.450 0.440 0.459 0.415 0.437 0.424 0.439
-avg 0.346 0.394 0.379 0.396 0.392 0.420 0.409 0.427 0.394 0.429
+From Scratch 96 0.301 0.357 0.171 0.260 0.419 0.439 0.315 0.389
+192 0.358 0.385 0.223 0.294 0.438 0.457 0.366 0.405
+336 0.390 0.396 0.282 0.336 0.484 0.484 0.424 0.435
+720 0.436 0.427 0.366 0.380 0.540 0.538 0.495 0.473
+avg 0.371 0.391 0.261 0.318 0.470 0.480 0.400 0.425
+A.10.4 Zero-shot results
+Table 15: Full results in zero-shot setting.
+Models ROSE_512 Timer MOIRAI Chronos TimesFM Moment
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 96 0.382 0.408 0.414 0.439 0.405 0.397 0.494 0.409 0.432 0.405 0.706 0.561
+192 0.400 0.420 0.440 0.455 0.458 0.428 0.561 0.443 0.492 0.438 0.716 0.579
+336 0.404 0.426 0.455 0.463 0.509 0.454 0.580 0.460 0.519 0.458 0.705 0.583
+720 0.420 0.447 0.496 0.496 0.529 0.494 0.605 0.495 0.512 0.477 0.705 0.597
+avg 0.401 0.425 0.451 0.463 0.475 0.443 0.560 0.452 0.489 0.444 0.708 0.580
+ETTh2 96 0.298 0.362 0.305 0.355 0.303 0.338 0.306 0.338 0.311 0.345 0.373 0.416
+192 0.336 0.385 0.365 0.406 0.369 0.384 0.396 0.394 0.401 0.397 0.384 0.422
+336 0.353 0.399 0.378 0.413 0.397 0.410 0.423 0.417 0.436 0.430 0.386 0.426
+720 0.395 0.432 0.414 0.457 0.447 0.450 0.442 0.439 0.437 0.450 0.425 0.454
+avg 0.346 0.394 0.366 0.408 0.379 0.396 0.392 0.397 0.396 0.405 0.392 0.430
+ETTm1 96 0.512 0.460 0.440 0.422 0.660 0.476 0.514 0.443 0.366 0.374 0.679 0.544
+192 0.512 0.462 0.505 0.458 0.707 0.500 0.608 0.475 0.413 0.401 0.690 0.550
+336 0.523 0.470 0.570 0.490 0.730 0.515 0.690 0.507 0.445 0.429 0.701 0.557
+720 0.552 0.490 0.659 0.534 0.758 0.536 0.733 0.555 0.513 0.470 0.719 0.569
+avg 0.525 0.471 0.544 0.476 0.714 0.507 0.636 0.495 0.434 0.419 0.697 0.555
+ETTm2 96 0.224 0.309 0.203 0.285 0.216 0.282 0.202 0.293 0.189 0.257 0.230 0.308
+192 0.266 0.333 0.265 0.327 0.294 0.330 0.286 0.348 0.277 0.325 0.285 0.338
+336 0.310 0.358 0.319 0.361 0.368 0.373 0.355 0.386 0.350 0.381 0.339 0.369
+720 0.395 0.407 0.405 0.410 0.494 0.439 0.409 0.425 0.464 0.448 0.423 0.424
+avg 0.299 0.352 0.360 0.386 0.343 0.356 0.313 0.363 0.320 0.353 0.319 0.360
+Weather 96 0.200 0.260 0.190 0.236 0.188 0.250 0.209 0.244 - - 0.216 0.271
+192 0.239 0.288 0.261 0.293 0.237 0.284 0.254 0.288 - - 0.264 0.306
+336 0.279 0.315 0.332 0.340 0.282 0.323 0.301 0.332 - - 0.313 0.336
+720 0.340 0.357 0.385 0.381 0.359 0.345 0.388 0.374 - - 0.369 0.380
+avg 0.265 0.305 0.292 0.312 0.267 0.300 0.288 0.310 - - 0.291 0.323
+Electricity 96 0.209 0.307 0.210 0.312 0.212 0.301 0.194 0.266 - - 0.844 0.761
+192 0.219 0.315 0.239 0.337 0.225 0.320 0.218 0.289 - - 0.850 0.762
+336 0.236 0.330 0.284 0.372 0.245 0.333 0.244 0.321 - - 0.862 0.766
+720 0.273 0.328 0.456 0.479 0.282 0.358 0.324 0.371 - - 0.888 0.774
+avg 0.234 0.320 0.297 0.375 0.241 0.328 0.245 0.312 - - 0.861 0.766
+Traffic 96 0.572 0.407 0.526 0.368 - - 0.562 0.378 - - 1.390 0.800
+192 0.575 0.406 0.561 0.385 - - 0.579 0.412 - - 1.403 0.802
+336 0.588 0.411 0.614 0.412 - - 0.594 0.420 - - 1.415 0.804
+720 0.618 0.422 0.749 0.464 - - 0.723 0.472 - - 1.437 0.808
+avg 0.588 0.412 0.613 0.407 - - 0.615 0.421 - - 1.411 0.804

### v3

@@
-ROSE: Register-assisted General Time Series Forecasting with Decomposed Frequency Learning
+Towards a General Time Series Forecasting Model with Unified Representation and Adaptive Transfer
 Abstract
-With the increasing collection of time series data from various domains, there arises a strong demand for general time series forecasting models pre-trained on a large number of time-series datasets to support a variety of downstream prediction tasks.
-Enabling general time series forecasting faces two challenges: how to obtain unified representations from multi-domian time series data, and how to capture domain-specific features from time series data across various domains for adaptive transfer in downstream tasks.
-To address these challenges, we propose a Register Assisted General Time Series Forecasting Model with Decomposed Frequency Learning (ROSE), a novel pre-trained model for time series forecasting.
-ROSE employs Decomposed Frequency Learning for the pre-training task, which decomposes coupled semantic information in time series with frequency-based masking and reconstruction to obtain unified representations across domains.
-We also equip ROSE with a Time Series Register, which learns to generate a register to capture domain-specific representations during pre-training and enhances domain-adaptive transfer by selecting related register tokens on downstream tasks.
-After pre-training on large-scale time series data, ROSE achieves state-of-the-art forecasting performance on 7 real-world benchmarks.
-Remarkably, it demonstrates competitive or superior few-shot and zero-shot abilities.
+With the growing availability of multi-domain time series data, there is an increasing demand for general forecasting models pre-trained on multi-source datasets to support diverse downstream prediction scenarios.
+Existing time series foundation models primarily focus on scaling up pre-training datasets and model sizes to enhance generalization performance.
+In this paper, we take a different approach by addressing two critical aspects of general forecasting models: (1) how to derive unified representations from heterogeneous multi-domain time series data, and (2) how to effectively capture domain-specific features to enable adaptive transfer across various downstream scenarios.
+To address the first aspect, we propose Decomposed Frequency Learning as the pre-training task, which leverages frequency-based masking and reconstruction to decompose coupled semantic information in time series, resulting in unified representations across domains.
+For the second aspect, we introduce the Time Series Register, which captures domain-specific representations during pre-training and enhances adaptive transferability to downstream tasks.
+Our model achieves the state-of-the-art forecasting performance on seven real-world benchmarks, demonstrating remarkable few-shot and zero-shot capabilities.
+Machine Learning, ICML
 1 Introduction
-Time Series Forecasting plays a pivotal role across numerous domains, such as energy, smart transportation, weather, and economics (Qiu et al., 2024).
-However, training specific models in deep learning for each dataset is costly and requires tailored parameter tuning, whose prediction accuracy may be limited due to data scarcity (Liu et al., 2024).
-One solution is to pre-train a general model on diverse time series datasets and fine-tune it with a few data for different downstream scenarios or direct predict without fine-tuning.
-Following this idea, foundation models for time series forecasting have recently raised growing attention, continually scaling up the pre-training datasets and model sizes to improve the generalization performance (Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024).
-However, an excessively large scale also causes increasing costs in training and inference, which may go against the original intention of a general model, especially in resource-constrained situations.
-In addition to scaling up, general time series forecasting models can also be designed from the perspectives of pre-training tasks and downstream transfer adaptation.
-From these two angles, we identify the following challenges.
+Time series forecasting plays a crucial role in various domains, including energy, smart transportation, weather, and economics (Qiu et al., 2024; Wu et al., 2025a; Qiu et al., 2025b).
+However, training deep learning models for each specific dataset is resource-intensive and requires tailored parameter tuning.
+This approach often suffers from limited prediction accuracy due to data scarcity (Liu et al., 2024; Wu et al., 2021b, 2024, 2025c).
+A promising solution is to pre-train a general model on diverse time series datasets, which can then be fine-tuned with minimal data for different downstream scenarios or even used directly without fine-tuning.
+Following this idea, foundation models for time series forecasting have gained significant attention.
+Recent efforts have focused on scaling up pre-training datasets and model sizes to enhance generalization performance (Woo et al., 2024; Goswami et al., 2024; Ansari et al., 2024).
+However, excessive scaling introduces high computational costs during training and inference, undermining the practicality of general models, particularly in resource-constrained settings.
+Beyond scaling, the design of general time series forecasting models can also be approached through pre-training tasks and downstream transfer adaptation.
+From these two perspectives, we identify the following challenges.
+Figure 1: (a) Pre-training on multi-domain datasets that exhibit combined frequency.
+Existing general time series forecasting models only extract generalized representations for direct transfer to various downstream target domains.
+We propose to learn generalized and specific representations during pre-training, and adaptively transfer them to each target domain.
+(b) The t-SNE visualization of the hidden representations after direct transfer and adaptive transfer: In direct transfer, representations of different domains are mixed, but in adaptive transfer, they show a clear clustering pattern.
+The detailed experiment setting is in the Appendix A.1.3
 Obtaining a unified representation from time series data across various domains is challenging.
-Time series from each domain involve complex temporal patterns, composed of multiple frequency components combined with each other (Zhou et al., 2022), which is frequency superposition.
+Time series from each domain involve complex temporal patterns, composed of multiple frequency components combined with each other (Zhou et al., 2022; Wu et al., 2025b), which is frequency superposition.
 As shown in Figure 1(a), different frequency components contain distinct semantic information.
@@
 Adaptive transferring information from multi-domain time series to specific downstream scenarios presents a challenge.
-Multi-source time series data originate from various domains (Woo et al., 2024), whose data exhibit domain-specific information (Liu et al., 2024).
+Multi-source time series data originate from various domains (Woo et al., 2024), whose data exhibit domain-specific information (Liu et al., 2024; Miao et al., 2024; Zhao et al., 2023; Guo et al., 2014).
 Information from the same or similar domain as the target domain is useful for improving the model’s effectiveness in the target task (Chen et al., 2023a).
-However, as shown in Figure 1(a), existing time series pre-training frameworks (Woo et al., 2024; Liu et al., 2024; Zhou et al., 2024) focus mainly on learning generalized representation during pre-training and overlook domain-specific representation.
-Thus, they only transfer the same generalized representation to different target domains, called direct transfer, which limits the model’s effectiveness in specific downstream tasks.
-Therefore, it is necessary to learn domain-specific information during pre-training and adaptively transfer the specific representations to each target domain, called adaptive transfer.
+However, as shown in Figure 1(a), existing time series pre-training frameworks (Woo et al., 2024; Liu et al., 2024; Zhou et al., 2024) primarily focus on learning generalized time series representations during pre-training while overlooking domain-specific representations, called direct transfer.
+While generalized representations are essential, directly transferring them to specific downstream tasks without incorporating domain-specific information leaves room for improvement.
+Therefore, it is necessary to learn domain-specific information during pre-training and adaptively transfer the specific representations to target domain, called adaptive transfer.
 Realizing adaptive transfer poses two difficulties: 1) capturing domain-specific information in pre-training.
 2) adaptive use of domain-specific information in various downstream tasks.
-Figure 1: (a) Pre-training on multi-domain datasets that exhibit combined frequency.
-Existing general time series forecasting models only extract generalized representations for direct transfer to various downstream target domains.
-We propose to learn generalized and specific representations during pre-training, and adaptively transfer them to each target domain.
-(b) The t-SNE visualization of the hidden representations after direct transfer and adaptive transfer: In direct transfer, representations of different domains are mixed, but in adaptive transfer, they show a clear clustering pattern.
-To address these challenges, we propose a register assisted general time series forecasting model with decomposed frequency learning (ROSE).
+To address these challenges, we propose a register assisted general time series forecasting model with decomposed frequency learning, namely ROSE.
 First, we propose Decomposed Frequency Learning that learns generalized representations to solve the issue with coupled semantic information.
@@
   • Our experiments with 7 real-world benchmarks demonstrate that ROSE achieves state-of-the-art performance in full-shot setting and achieves competitive or superior results in few-shot setting, along with impressive transferability in zero-shot setting.
+Figure 2: The model architecture of ROSE.
 2 Related Work
 2.1 Traditional Time Series Forecasting
-The statistical time series forecasting models like ARIMA (Box and Jenkins, 1968), despite their theoretical support, are limited in modeling nonlinearity.
+The statistical time series forecasting models like ARIMA (Box & Jenkins, 1968), despite their theoretical support, are limited in modeling nonlinearity.
 With the rise of deep learning, many RNN-based models (Cirstea et al., 2019; Wen et al., 2017; Salinas et al., 2020) have been proposed, modeling the sequential data with an autoregressive process.
-CNN-based models (Luo and Wang, 2024; Liu et al., 2022a) have also received widespread attention due to their ability to capture local features.
+CNN-based models (Luo & Wang, 2024; Liu et al., 2022b) have also received widespread attention due to their ability to capture local features.
 MICN (Wang et al., 2022) utilizes TCN to capture both local and global features, while TimesNet (Wu et al., 2022) focuses on modeling 2D temporal variations.
 However, both RNNs and CNNs struggle to capture long-term dependencies.
-Transformer-based models (Zhou et al., 2022; Nie et al., 2022; Wu et al., 2021; Liu et al., 2023; Chen et al., 2024), with their attention mechanism, can capture long dependencies and extract global information, leading to widespread applications in long-time series prediction.
+Transformer-based models (Zhou et al., 2022; Nie et al., 2022; Wu et al., 2021a; Liu et al., 2023; Chen et al., 2024), with their attention mechanism, can capture long dependencies and extract global information, leading to widespread applications in long-time series prediction.
 However, this case-by-case paradigm requires meticulous hyperparameter design for different datasets, and its predictive performance can also be affected by data scarcity.
-2.2 Pre-trained models for times seires forecasting
-LLMs for time series forecasting: Recent studies have shown that Large Language Models (LLMs) can enhance time series forecasting with limited fine-tuning (Zhou et al., 2024), prompting (Jin et al., 2024; Cao et al., 2024), and modality alignment (Jin et al., 2024).
-GPT4TS (Zhou et al., 2024) fine-tunes a subset of LLM parameters, and achieves competitive performance by leveraging pre-trained text knowledge.
-TimeLLM (Jin et al., 2024) transforms time series into text to align with LLM representations.
-Despite their positive impact, LLMs’ high inference costs limit their practicality.
-Time series foundation model: Pre-training with multiple sources time series has recently received widespread attention (Rasul et al., 2023; Dooley et al., 2024; Garza and Mergenthaler-Canseco, 2023; Kamarthi and Prakash, 2023).
+2.2 Time Series Forecasting Foundation Model
+Pre-training with multiple sources time series has recently received widespread attention (Rasul et al., 2023; Dooley et al., 2024; Garza & Mergenthaler-Canseco, 2023; Kamarthi & Prakash, 2023).
 MOMENT (Goswami et al., 2024) and MOIRAI (Woo et al., 2024) adopt a BERT-style pre-training approach, while Timer (Liu et al., 2024), Chronos (Ansari et al., 2024) and TimsFM (Das et al., 2023a) use a GPT-style pre-training approach, giving rise to improved performance in time series prediction.
@@
 3 Methodology
-Problem Definition.
-Given a multivariate time series $\mathbf{X}_{t}=\{\mathbf{x}_{t-L:t}^{i}\}_{i=1}^{C}$, where each $\mathbf{x}_{t-L:t}^{i}\in\mathbb{R}^{L}$ is a sequence of observations.
-$L$ denotes the look-back window and $C$ denotes the number of channels.
-The forecasting task is to predict the future values $\mathbf{\hat{Y}}_{t}=\{\mathbf{\hat{x}}_{t:t+F}^{i}\}_{i=1}^{C}$, where $F$ denotes the forecast horizon.
-$\mathbf{Y}_{t}=\{\mathbf{x}_{t:t+F}^{i}\}_{i=1}^{C}$ is the ground truth of future.
-The general time series forecasting model is pre-trained with multi-source datasets $\mathbf{D}_{\text{pre-train}}=\{(\mathbf{X}_{t}^{j},\mathbf{Y}_{t}^{j})\}^{N}_{j=1}$, where $N$ is the number of datasets.
-For the downstream task, the model is fine-tuned with a training dataset $\mathbf{D}_{\text{train}}=\{(\mathbf{X}_{t}^{\text{train}},\mathbf{Y}_{t}^{\text{train}})\}$, and is tested with $\mathbf{D}_{\text{test}}=\{(\mathbf{X}_{t}^{\text{test}},\mathbf{Y}_{t}^{\text{test}})\}$ to predict $\mathbf{\hat{Y}}_{t}^{\text{test}}$, where $\mathbf{D}_{\text{pre-train}}$, $\mathbf{D}_{\text{train}}$ and $\mathbf{D}_{\text{test}}$ are pairwise disjoint.
-Alternatively, the model could be directly tested using $\mathbf{D}_{\text{test}}$ without fine-tuning with $\mathbf{D}_{\text{train}}$ to predict $\mathbf{\hat{Y}}_{t}^{\text{test}}$.
 3.1 Architecture
-Figure 2: The model architecture of ROSE.
-The time series forecasting paradigm of ROSE contains two steps: pre-training and fine-tuning.
-First, ROSE is pre-trained on large-scale datasets from various domains, with two pre-training tasks: reconstruction and prediction.
-We set up the reconstruction task to help the model understand time series comprehensively and set the prediction task to enhance the model’s few-shot and zero-shot abilities.
-Then, ROSE is fine-tuned with a target dataset in the downstream scenario.
-As shown in Figure 2, ROSE employs the encoder-decoder architecture to model time series.
-The backbone consists of multiple Transformer layers to process sequential information and effectively capture temporal dependencies (Vaswani et al., 2017).
-The reconstruction decoder and prediction decoder use the same structure as the Transformer encoder.
-They are used for reconstruction and prediction tasks respectively.
+As illustrated in Figure 2, ROSE adopts an encoder-decoder architecture for time series modeling.
+Its backbone comprises multiple Transformer layers, which effectively process sequential information and capture temporal dependencies (Vaswani et al., 2017).
+Both the reconstruction decoder and prediction decoder share the same structure as the Transformer encoder and are designed for reconstruction and prediction tasks, respectively.
+The reconstruction task enables the model to gain a comprehensive understanding of time series, while the prediction task enhances its few-shot and zero-shot capabilities.
 ROSE is pre-trained in a channel-independent way, which is widely used in time series forecasting (Nie et al., 2022).
@@
 Register tokens that capture domain-specific information will be introduced in Section 3.3.
+Figure 3: An illustration of decomposed frequency learning.
+Based on the sampled thresholds, we randomly apply low/high-frequency masking to the time series in the frequency domain and then transform it back to the time domain for reconstruction.
 3.2 Decomposed Frequency Learning
@@
 In contrast, existing frequency masking methods (Zhang et al., 2022), which randomly mask frequencies of a single time series once, show limited forecasting effectiveness due to the lack of common pattern learning from heterogeneous time series that come from various domains.
-Figure 3: An illustration of decomposed frequency learning.
-Based on the sampled thresholds, we randomly apply low/high-frequency masking to the time series in the frequency domain and then transform it back to the time domain for reconstruction.
 Multi-frequency masking.
-As shown in the green part of Figure 3, given a time series $\mathbf{x}\in\mathbb{R}^{L}$, we utilize the Real Fast Fourier Transform (rFFT) (Brigham and Morrow, 1967) to transform it into the frequency domain, giving rise to $\mathbf{x}_{\mathrm{freq}}\in\mathbb{C}^{L/2+1}$.
+As shown in the green part of Figure 3, given a time series $\mathbf{x}\in\mathbb{R}^{L}$, we utilize the Real Fast Fourier Transform (rFFT) (Brigham & Morrow, 1967) to transform it into the frequency domain, giving rise to $\mathbf{x}_{\mathrm{freq}}\in\mathbb{C}^{L/2+1}$.
 $\mathbf{x}_{\mathrm{freq}}=\mathrm{rFFT}(\mathbf{x}).$ (1)
@@
 As shown in the yellow part of Figure 3, after obtaining the $K_{\mathrm{f}}$ masked sequences $\mathbf{X}_{\mathrm{mask}}$, we divide each sequence $\mathbf{x}_{\mathrm{mask}}^{i}$ into $P$ non-overlapping patches, and use a linear layer to transforming them into $P$ patch tokens, and thus we get $\mathcal{X}_{\mathrm{mp}}=\{\mathbf{X}_{\mathrm{mp}}^{i}\}_{i=1}^{K_{\mathrm{f}}}$ to capture general information, where each $\mathbf{X}_{\mathrm{mp}}^{i}\in\mathbb{R}^{P\times D}$, and $D$ is the dimension for each patch token.
-We replicate the register tokens $\mathbf{X}_{\mathrm{u}}$ $K_{\mathrm{f}}$ times to get $\mathcal{X}_{\mathrm{u}}\in\mathbb{R}^{K_{\mathrm{f}}\times N_{\mathrm{r}}\times D}$, where $\mathbf{X}_{\mathrm{u}}\in\mathbb{R}^{N_{\mathrm{r}}\times D}$ is obtained by inputting the original sequence into the TS-Register, as detailed in Section 3.3.
+We replicate the register tokens $\mathbf{X}_{\mathrm{u}}$ $K_{\mathrm{f}}$ times to get $\mathcal{X}_{\mathrm{u}}\in\mathbb{R}^{K_{\mathrm{f}}\times N_{\mathrm{r}}\times D}$, where $\mathbf{X}_{\mathrm{u}}$ is obtained by inputting the original sequence into the TS-Register, as detailed in Section 3.3.
 Then, we concatenate the patch tokens $\mathcal{X}_{\mathrm{mp}}$ with the register tokens $\mathcal{X}_{\mathrm{u}}$, and feed them into the Transformer encoder to obtain the representation of each masked series.
-These representations are then aggregated to yield a unified representation $\mathbf{S}_{\mathrm{m}}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$.
-The aggregator is the averaging operation.
-$\mathbf{S}_{\mathrm{m}}=\mathrm{Aggregator(Encoder(Concatenate(\mathcal{X}_{mp},\mathcal{X}_{u})))}.$ (4)
+These representations are then averaging aggregated to yield a unified representation $\mathbf{S}_{\mathrm{m}}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$.
+$\mathbf{S}_{\mathrm{m}}=\mathrm{Aggregator(Encoder(Concat(\mathcal{X}_{mp},\mathcal{X}_{u})))}.$ (4)
 Reconstruction task.
@@
 As frequency domain masking affects the overall time series, we compute the Mean Squared Error (MSE) reconstruction loss for the entire time series.
-$\mathcal{L}_{\mathrm{reconstruction}}=\mathrm{||\mathbf{x}-\hat{\mathbf{x}}||_{2}^{2}}.$ (5)
+[…]
 3.3 Time Series Register
-By decomposed frequency learning, we can obtain the general representations, and additionally, we propose the TS-Register that learns register tokens as the domain-specific information from the multi-domain datasets for adaptive transfer.
+By decomposed frequency learning, we can obtain the general representations.
+Additionally, we propose the TS-Register that learns domain-specific information from the multi-domain datasets for adaptive transfer.
 It clusters domain-specific information from the multi-domain datasets into register tokens and stores such domain-specific information in the register during pre-training.
@@
 $\mathcal{L}_{\text{register}}=\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{\delta}\|^{2}_{2},\ \ \delta=\underset{j=1:H}{\arg\operatorname{min}}\left\|\mathbf{x}_{\mathrm{e}}-\mathbf{e}_{j}\right\|_{2}.$ (6)
-To update the cluster center vectors in the register $\mathbf{E}$ that represents the domain information of the pre-trained datasets, we set the loss function shown in Equation 6 that minimizes the distance between the data-dependent embedding $\mathbf{x}_{\mathrm{e}}$ and the cluster center $\mathbf{e}_{\delta}$.
+To update the cluster center vectors in the register $\mathbf{E}$ that represents the domain information of the pre-trained datasets, we set the loss function shown in Equation 6 that minimizes the distance between the embedding $\mathbf{x}_{\mathrm{e}}$ and the cluster center $\mathbf{e}_{\delta}$.
 To solve the problem that the gradient of the $\arg\operatorname{min}$ function cannot be backpropagated, we use the stop gradient operation to pass the gradient of $\mathbf{e}_{\delta}$ directly to $\mathbf{x}_{\mathrm{e}}$.
@@
 3.4 Training
-To improve the model’s prediction ability in zero-shot and few-shot setting, we co-train supervised prediction with self-supervised reconstruction that uses multi-frequency masking to learn unified features that are more applicable to the downstream prediction task.
-We normalize time series by employing the REVIN (Kim et al., 2021) that is commonly used by the state-of-the-art time series models, which first normalizes each input sample and subsequently applies inverse normalization to recover the model’s output.
+To improve the prediction performance in zero-shot and few-shot settings, we co-train supervised prediction with self-supervised reconstruction that uses multi-frequency masking to learn unified features that are more applicable to the downstream prediction task.
 Prediction task.
 The input time series $\mathbf{x}\in\mathbb{R}^{L}$ is sliced into $P$ non-overlapping patches and then mapped to $\mathbf{X}_{\mathrm{p}}\in\mathbb{R}^{P\times D}$.
-Based on common forecasted needs (Qiu et al., 2024), we set up four prediction heads mapping to prediction lengths of $\{96,192,336,720\}$ to accomplish the prediction task.
+Based on common forecasting needs (Qiu et al., 2024), we set up four prediction heads mapping to prediction lengths of $\{96,192,336,720\}$ to accomplish the prediction task.
 Patch tokens $\mathbf{X}_{\mathrm{p}}$ are concatenated with the register tokens $\mathbf{X}_{\mathrm{u}}$ and then successively fed into the Transformer encoder to yield the representation $\mathbf{S}\in\mathbb{R}^{(N_{\mathrm{r}}+P)\times D}$:
@@
 With the ground truth $\mathbf{Y}_{F}$, the prediction loss $\mathcal{L}_{\text{prediction}}$ is shown in Equation 11.
-$\mathcal{L}_{\text{prediction}}=\sum_{F\in\{96,192,336,720\}}||{\mathbf{Y}_{F}-{\mathbf{\hat{Y}}_{F}}}||_{2}^{2}.$ (11)
+[…]
 Pre-training.
@@
 The datasets are crucial for pre-training a general time series forecasting model.
-In light of this, we gather a considerable amount of publicly available datasets from various domains, including energy, nature, health, transport, web, and economics, etc.
+In light of this, we gather many publicly available datasets from various domains, including energy, nature, health, transport, web, economics, etc.
 The details of these datasets are shown in the Appendix A.1.1.
@@
 Evaluation datasets.
-To conduct comprehensive and fair comparisons for different models, we conducted experiments on seven well-known forecasting benchmarks as the target datasets, including Weather, Traffic, Electricity, and ETT (4 subsets), which cover multiple domains.
+To conduct comprehensive and fair comparisons for different models, we conduct experiments on seven well-known forecasting benchmarks as the target datasets, including Weather, Traffic, Electricity, and ETT (4 subsets), which cover multiple domains.
 Baselines.
-We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023), and two LLM-based models: GPT4TS (Zhou et al., 2024) and $\textbf{S}^{2}$-LLM (Pan et al., 2024).
+We select the state-of-the-art models as baselines in full-shot and few-shot setting, including four specific models: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023), and two LLM-based models: GPT4TS (Zhou et al., 2024) and $\textbf{S}^{2}$IP-LLM (Pan et al., 2024).
 In addition, we select five foundation models for comparison in zero-shot setting, including Timer (Liu et al., 2024), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., 2024), TimesFM (Das et al., 2023b), and Moment (Goswami et al., 2024).
@@
 Consistent with previous works, we adopted Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics.
-For fair comparison, all methods fix the look-back window $L$ = 512 and predict the future values with lengths $F=\{96,192,336,720\}$.
+Due to ROSE mostly aims at long-term predictions, for fair comparison, all methods fix the look-back window $L$ = 512 and predict the future values with lengths $F=\{96,192,336,720\}$.
 More implementation details are presented in the Appendix A.1.3.
 4.1 In Distribution Forecasting
-Setting. In full-shot setting, we fine-tune pre-trained ROSE and the baselines with the full downstream data. In few-shot setting, we fine-tune all models with only 10% train data.
+Setting.
+In full-shot setting, we utilize full downstream data to fine-tune pre-trained ROSE and baselines.
+In few-shot setting, we fine-tune all models with only 10% train data.
+The "Drop Last" issue is reported by several researchers (Qiu et al., 2024, 2025a; Li et al., 2025).
+That is, in some previous works evaluating the model on test set with drop-last=True setting may cause additional errors related to test batch size.
+In our experiment, to ensure fair comparison in the future, we set the drop last to False for all baselines to avoid this issue.
 Full-shot results.
@@
 Table 1: The results for ROSE in full-shot setting and 10% few-shot setting, compared with other methods in full-shot setting. The average results of all predicted lengths are listed here.
-Models ROSE ROSE (10%) ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$-LLM
+Models ROSE ROSE (10%) ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$IP-LLM
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 Traffic 0.390 0.264 0.418 0.278 0.397 0.282 0.391 0.264 0.623 0.333 0.433 0.305 0.414 0.294 0.405 0.286
-Few-shot results. The results under the 10% few-shot setting are presented in Table 13 in Appendix A.10.2. ROSE outperforms advanced models when training data is scarce in the target domain.
-Figure 4: The forecasting results of ROSE obtained by training from scratch and fine-tuning from the pre-trained model. The right, upper corner is the best case.
+Few-shot results.
+The results under the 10% few-shot setting are presented in Table 13 in Appendix A.10.2.
+ROSE outperforms advanced models when training data is scarce in the target domain.
 Figure 4 shows the performance of pre-trained ROSE and ROSE trained from scratch on ETTh1 and ETTm2 with different fine-tuning data percentages, noting the best baselines in full-shot setting.
@@
 Specifically, the pre-trained ROSE exceeds SOTA performance with only 1% train data for ETTh1 and 2% for ETTm2.
-Moreover, compared to the ROSE trained from scratch, the pre-trained ROSE exhibits a slower decline in predictive performance with the reduction of fine-tuning data, demonstrating the impressive generalization ability of ROSE through pre-training.
+Moreover, compared to the ROSE trained from scratch, the pre-trained ROSE exhibits a slower decline in prediction performance with the reduction of fine-tuning data, demonstrating the impressive generalization ability of ROSE through pre-training.
+Figure 4: The forecasting results of ROSE obtained by training from scratch and fine-tuning from the pre-trained model. The right, upper corner is the best case.
 4.2 Zero-shot Forecasting
@@
 Results.
-As shown in Table 2, ROSE significantly outperforms across the majority of datasets, achieving an average reduction of 15% in Mean Squared Error (MSE).
+As shown in Table 2, ROSE significantly outperforms across the majority of datasets, achieving an average reduction of 15% in MSE.
 In comparison to Timer and Moirai, ROSE achieves average MSE reductions of 9% and 6%, respectively, and demonstrates a remarkable 43% relative improvement over Moment.
-Notably, ROSE stands out not only for its superior performance but also for its exceptional lightweight and efficient design, which sets it apart from other foundational models.
+Notably, ROSE stands out not only for its superior performance but also for its exceptionally lightweight and efficient design, which sets it apart from other foundational models.
 Detailed analysis of these aspects will be presented in Section 4.3.
-Table 2: The results for ROSE and other foundation models in zero-shot setting.
+Table 2: The results for ROSE and other foundation models in the zero-shot setting.
 The average results of all predicted lengths are listed here.
@@
 Models ROSE Timer MOIRAI Chronos TimesFM Moment
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 0.401 0.425 0.451 0.463 0.475 0.443 0.560 0.452 0.489 0.444 0.708 0.580
-ETTh2 0.346 0.394 0.366 0.408 0.379 0.396 0.392 0.397 0.396 0.405 0.392 0.430
-ETTm1 0.525 0.471 0.544 0.476 0.714 0.507 0.636 0.495 0.434 0.419 0.697 0.555
-ETTm2 0.299 0.352 0.360 0.386 0.343 0.356 0.313 0.363 0.320 0.353 0.319 0.360
-Weather 0.265 0.305 0.292 0.312 0.267 0.300 0.288 0.310 - - 0.291 0.323
-Electricity 0.234 0.320 0.297 0.375 0.241 0.328 0.245 0.312 - - 0.861 0.766
-Traffic 0.588 0.412 0.613 0.407 - - 0.371 0.370 - - 1.411 0.804
+Metric MSE MSE MSE MSE MSE MSE
+ETTh1 0.401 0.451 0.475 0.560 0.489 0.708
+ETTh2 0.346 0.366 0.379 0.392 0.396 0.392
+ETTm1 0.525 0.544 0.714 0.636 0.434 0.697
+ETTm2 0.299 0.360 0.343 0.313 0.320 0.319
+Weather 0.265 0.292 0.267 0.288 - 0.291
+Electricity 0.234 0.297 0.241 0.245 - 0.861
+Traffic 0.588 0.613 - 0.615 - 1.411
 4.3 Model Analysis
@@
 Importantly, ROSE uses the least number of parameters among foundation models, with its parameter count approaching that of specific models, while exhibiting superior zero-shot performance.
-This is attributed to our proposed decomposed frequency learning that enhances the model’s comprehension of time series.
+This is attributed to our proposed decomposed frequency learning that enhances the comprehension of time series.
 Concurrently, the TS-Register achieves the adaptive transfer thus efficiently adapting to downstream tasks without the need of scaling up to achieve strong generalizability.
@@
 ROSE 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.380
-Replace Multi-Frequency Masking Random Freq Masking 0.381 0.397 0.261 0.324 0.410 0.427 0.374 0.405
+Random Freq Masking 0.381 0.397 0.261 0.324 0.410 0.427 0.374 0.405
 Multi-Patch Masking 0.356 0.379 0.259 0.316 0.404 0.426 0.349 0.389
@@
 Future efforts will concentrate on expanding pre-training datasets and extending ROSE’s applicability across diverse time series analysis tasks, e.g., classification.
-We provide our code at https://anonymous.4open.science/r/ROSE-A235.
-6 Reproducibility
-Our work meets reproducibility requirements.
-Specifically, you can download our evaluation datasets in a standardized format from the public link (Wang et al., 2024a): https://drive.google.com/drive/folders/13Cg1KYOlzM5C7K8gK8NfC-F3EYxkM3D2, and we provide our code at https://anonymous.4open.science/r/ROSE-A235.
+We provide our code at https://github.com/decisionintelligence/ROSE.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Acknowledgements
+This work was partially supported by National Natural Science Foundation of China (62372179). Chenjuan Guo is the corresponding author of the work.
 Appendix A Appendix
@@
 A.1.1 Pre-training Datasets
-We use multi-source datasets in pre-training which contain subsets of Monash [Godahewa et al., 2021], UEA [Bagnall et al., 2018] and UCR [Dau et al., 2019] time series datasets, as well as some other time series classical datasets [Zhang et al., 2017, Wang et al., 2024b, Liu et al., 2022b, McCracken and Ng, 2016, Taieb et al., 2012].
+We use multi-source datasets in pre-training which contain subsets of Monash (Godahewa et al., 2021), UEA (Bagnall et al., 2018) and UCR (Dau et al., 2019) time series datasets, as well as some other time series classical datasets (Zhang et al., 2017; Wang et al., 2024; Liu et al., 2022a; McCracken & Ng, 2016; Taieb et al., 2012).
 The final list of all pre-training datasets is shown in Table 5.
@@
 For all pre-training datasets, we split them into univariate sequences and train them in a channel-independent manner.
+Table 5: List of pretraining datasets.
+Domain Dataset Frequency Time Pionts Source
+Energy Aus. Electricity Demand Half Hourly 1155264 Monash(Godahewa et al., 2021)
+Wind 4 Seconds 7397147 Monash(Godahewa et al., 2021)
+Wind Farms Minutely 172178060 Monash(Godahewa et al., 2021)
+Solar 10 Minutes 7200720 Monash(Godahewa et al., 2021)
+Solar Power 4 Seconds 7397222 Monash(Godahewa et al., 2021)
+London Smart Meters Half Hourly 166527216 Monash(Godahewa et al., 2021)
+Nature Phoneme - 2160640 UCR(Dau et al., 2019)
+EigenWorms - 27947136 UEA(Bagnall et al., 2018)
+PRSA Hourly 4628448 (Zhang et al., 2017)
+Temperature Rain Daily 23252200 Monash(Godahewa et al., 2021)
+StarLightCurves - 9457664 UCR(Dau et al., 2019)
+Worms 0.033 Seconds 232200 UCR(Dau et al., 2019)
+Saugeen River Flow Daily 23741 Monash(Godahewa et al., 2021)
+Sunspot Daily 73924 Monash(Godahewa et al., 2021)
+Weather Daily 43032000 Monash(Godahewa et al., 2021)
+KDD Cup 2018 Daily 2942364 Monash(Godahewa et al., 2021)
+US Births Daily 7305 Monash(Godahewa et al., 2021)
+Health MotorImagery 0.001 Seconds 72576000 UEA(Bagnall et al., 2018)
+SelfRegulationSCP1 0.004 Seconds 3015936 UEA(Bagnall et al., 2018)
+SelfRegulationSCP2 0.004 Seconds 3064320 UEA(Bagnall et al., 2018)
+AtrialFibrillation 0.008 Seconds 38400 UEA(Bagnall et al., 2018)
+PigArtPressure - 624000 UCR(Dau et al., 2019)
+PIGCVP - 624000 UCR(Dau et al., 2019)
+TDbrain 0.002 Seconds 79232703 (Wang et al., 2024)
+Transport Pems03 5 Minute 9382464 (Liu et al., 2022a)
+Pems04 5 Minute 5216544 (Liu et al., 2022a)
+Pems07 5 Minute 24921792 (Liu et al., 2022a)
+Pems08 5 Minute 3035520 (Liu et al., 2022a)
+Pems-bay 5 Minute 16937700 (Liu et al., 2022a)
+Pedestrian_Counts Hourly 3132346 Monash(Godahewa et al., 2021)
+Web Web Traffic Daily 116485589 Monash(Godahewa et al., 2021)
+Economic FRED_MD Monthly 77896 (McCracken & Ng, 2016)
+Bitcoin Daily 75364 Monash(Godahewa et al., 2021)
+NN5 Daily 87801 (Taieb et al., 2012)
 A.1.2 Evaluation Datasets
@@
 Split Ratio 6:2:2 6:2:2 6:2:2 6:2:2 7:1:2 7:1:2 7:1:2
-Table 6: List of pretraining datasets.
-Domain Dataset Frequency Time Pionts Source
-Energy Aus. Electricity Demand Half Hourly 1155264 Monash[Godahewa et al., 2021]
-Wind 4 Seconds 7397147 Monash[Godahewa et al., 2021]
-Wind Farms Minutely 172178060 Monash[Godahewa et al., 2021]
-Solar 10 Minutes 7200720 Monash[Godahewa et al., 2021]
-Solar Power 4 Seconds 7397222 Monash[Godahewa et al., 2021]
-London Smart Meters Half Hourly 166527216 Monash[Godahewa et al., 2021]
-Nature Phoneme - 2160640 UCRDau et al. [2019]
-EigenWorms - 27947136 UEA[Bagnall et al., 2018]
-PRSA Hourly 4628448 [Zhang et al., 2017]
-Temperature Rain Daily 23252200 Monash[Godahewa et al., 2021]
-StarLightCurves - 9457664 UCR[Dau et al., 2019]
-Worms 0.033 Seconds 232200 UCR[Dau et al., 2019]
-Saugeen River Flow Daily 23741 Monash[Godahewa et al., 2021]
-Sunspot Daily 73924 Monash[Godahewa et al., 2021]
-Weather Daily 43032000 Monash[Godahewa et al., 2021]
-KDD Cup 2018 Daily 2942364 MonashGodahewa et al. [2021]
-US Births Daily 7305 Monash[Godahewa et al., 2021]
-Health MotorImagery 0.001 Seconds 72576000 UEA[Bagnall et al., 2018]
-SelfRegulationSCP1 0.004 Seconds 3015936 UEA[Bagnall et al., 2018]
-SelfRegulationSCP2 0.004 Seconds 3064320 UEA[Bagnall et al., 2018]
-AtrialFibrillation 0.008 Seconds 38400 UEA[Bagnall et al., 2018]
-PigArtPressure - 624000 UCR[Dau et al., 2019]
-PIGCVP - 624000 UCR[Dau et al., 2019]
-TDbrain 0.002 Seconds 79232703 [Wang et al., 2024b]
-Transport Pems03 5 Minute 9382464 [Liu et al., 2022b]
-Pems04 5 Minute 5216544 [Liu et al., 2022b]
-Pems07 5 Minute 24921792 [Liu et al., 2022b]
-Pems08 5 Minute 3035520 [Liu et al., 2022b]
-Pems-bay 5 Minute 16937700 [Liu et al., 2022b]
-Pedestrian_Counts Hourly 3132346 Monash[Godahewa et al., 2021]
-Web Web Traffic Daily 116485589 Monash[Godahewa et al., 2021]
-Economic FRED_MD Monthly 77896 [McCracken and Ng, 2016]
-Bitcoin Daily 75364 Monash[Godahewa et al., 2021]
-NN5 Daily 87801 [Taieb et al., 2012]
 A.1.3 Setting
-We implemented ROSE in PyTorch [Paszke et al., 2019] and all the experiments were conducted on 8 NVIDIA A800 80GB GPU.
-We used ADAM [Kingma and Ba, 2014] with an initial learning rate of $5\times 10^{-4}$ and implemented learning rate decay using the StepLR method to implement learning rate decaying pre-training.
+We implemented ROSE in PyTorch (Paszke et al., 2019) and all the experiments were conducted on 8 NVIDIA A800 80GB GPU.
+We used ADAM (Kingma & Ba, 2014) with an initial learning rate of $5\times 10^{-4}$ and implemented learning rate decay using the StepLR method to implement learning rate decaying pre-training.
 By default, ROSE contains 3 encoder layers and 3 decoder layers with head number of 16 and the dimension of latent space $D=256$.
@@
 The number of register tokens $N_{\mathrm{r}}$ and patch tokens $P$ is the same as in pre-training, and the parameter $k=3$ in TopK is set when selection vectors are performed in the register.
+The t-SNE visualization.
+We select three datasets (Pems08, PSRA, Electricity) from transport, nature and energy domains respectively and compare the differences in hidden representations between direct transfer and adaptive transfer.
+Specifically, direct transfer refers to the case where domain specific information is not considered, while adaptive transfer considers domain specific information that is learned by register tokens.
+We visualized the output of the encoder’s hidden representations using t-SNE.
 A.1.4 Baselines
-We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer [Liu et al., 2023], PatchTST [Nie et al., 2022], TimesNet [Wu et al., 2022], and DLinear [Zeng et al., 2023], and two LLM-based models: GPT4TS [Zhou et al., 2024] and $\textbf{S}^{2}$-LLM [Pan et al., 2024].
-In addition, we selected five foundation models for comparison in zero-shot setting, including Timer [Liu et al., 2024], MOIRAI [Woo et al., 2024], Chronos [Ansari et al., 2024], TimesFM [Das et al., 2023b] and Moment [Goswami et al., 2024].
+We select the state-of-the-art models as our baselines in full-shot and few-shot setting, including four specific models: iTransformer (Liu et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022), and DLinear (Zeng et al., 2023), and two LLM-based models: GPT4TS (Zhou et al., 2024) and $\textbf{S}^{2}$IP-LLM (Pan et al., 2024).
+In addition, we selected five foundation models for comparison in zero-shot setting, including Timer (Liu et al., 2024), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., 2024), TimesFM (Das et al., 2023b) and Moment (Goswami et al., 2024).
+The zero-shot experiment of Moment is designed based on the reconstruction task in the pre-training phase.
+Specifically, to ensure consistency between pre-training (reconstruction) and downstream prediction tasks, we actively mask the time periods to be predicted in the input sequence, and directly use the model’s reconstruction output for this part as the prediction value.
+Moment itself is not designed for zero-shot prediction and does not officially support zero-shot forecasting in this manner.
 The specific code base for these models is listed in Table 7:
@@
 LLM-based Model GPT4TS https://github.com/DAMO-DI-ML/NeurIPS2023-One-Fits-All
-S2P-LLM https://github.com/panzijie825/S2IP-LLM
+S2IP-LLM https://github.com/panzijie825/S2IP-LLM
 A.2 Efficiency Analysis
@@
 Chronos 46M 84B 176s
-TimesFM 200M 100B 167.2s
-Moment 385M 1.13B 88s
+TimesFM 200M 100B 10.5s
+Moment 385M 1.13B 13s
 Itransformer 3.8M - 34.18s
@@
 A.6 Short-term Forecasting
-We also try to apply ROSE to short-term forecasting on the M4 [Makridakis, 2018] dataset, which contains the yearly, quarterly and monthly collected univariate marketing data.
-We follow TimesNet’s [Wu et al., 2022] setting and metrics (SMAPE, MASE and OWA) for testing.
+We also try to apply ROSE to short-term forecasting on the M4 (Makridakis, 2018) dataset, which contains the yearly, quarterly and monthly collected univariate marketing data.
+We follow TimesNet’s (Wu et al., 2022) setting and metrics (SMAPE, MASE and OWA) for testing.
 As shown in Table 9, ROSE also exhibits competitive performance on the M4 dataset compared to the baselines.
@@
 It is notable that multi-frequency masking consistently improves these forecasting models.
-Specifically, it achieves average improvements of 6.3%, 3.7%, 1.5% in Autoformer [Wu et al., 2021], TimesNet [Wu et al., 2022], and PatchTST [Nie et al., 2022], respectively.
+Specifically, it achieves average improvements of 6.3%, 3.7%, 1.5% in Autoformer (Wu et al., 2021a), TimesNet (Wu et al., 2022), and PatchTST (Nie et al., 2022), respectively.
 This indicates that multi-frequency Masking can be widely utilized across various time series forecasting models to learn generalized time series representations and improve prediction accuracy.
@@
 Metric MSE MAE MSE MAE MSE MAE MSE MAE
-Autoformer +Multi-frequency Masking 0.600 0.521 0.328 0.365 0.493 0.487 0.452 0.458
+Autoformer
++Multi-frequency Masking
+0.600 0.521 0.328 0.365 0.493 0.487 0.452 0.458
 0.549 0.488 0.306 0.349 0.474 0.478 0.406 0.425
-TimesNet +Multi-frequency Masking 0.400 0.406 0.291 0.333 0.458 0.450 0.414 0.427
+TimesNet
++Multi-frequency Masking
+0.400 0.406 0.291 0.333 0.458 0.450 0.414 0.427
 0.386 0.398 0.282 0.324 0.446 0.438 0.386 0.403
-PatchTST +Multi-frequency Masking 0.353 0.382 0.256 0.317 0.413 0.434 0.331 0.381
+PatchTST
++Multi-frequency Masking
+0.353 0.382 0.256 0.317 0.413 0.434 0.331 0.381
 0.347 0.372 0.252 0.308 0.405 0.424 0.337 0.379
@@
 Metric MSE MAE MSE MAE MSE MAE -
-ETTm1 0.342$\pm$0.003 0.367$\pm$0.002 0.349$\pm$0.003 0.372$\pm$0.002 0.349$\pm$ 0.004 0.383$\pm$0.003 99%
-ETTm2 0.246$\pm$0.002 0.303$\pm$0.004 0.249$\pm$0.002 0.308$\pm$0.002 0.255$\pm$0.002 0.314$\pm$0.003 99%
-ETTh1 0.392$\pm$0.004 0.413$\pm$0.004 0.397$\pm$0.003 0.419$\pm$0.003 0.411$\pm$0.003 0.432$\pm$0.005 99%
-ETTh2 0.330$\pm$0.003 0.374$\pm$0.002 0.335$\pm$0.004 0.380$\pm$0.003 0.348$\pm$0.004 0.390$\pm$0.004 99%
-Traffic 0.391$\pm$0.008 0.266$\pm$0.005 0.418$\pm$0.011 0.278$\pm$0.006 0.404$\pm$0.009 0.283$\pm$0.002 99%
-Weather 0.217$\pm$0.008 0.250$\pm$0.007 0.224$\pm$0.007 0.252$\pm$0.009 0.223$\pm$0.011 0.263$\pm$0.014 99%
-Electricity 0.156$\pm$0.007 0.249$\pm$0.009 0.164$\pm$0.004 0.253$\pm$0.004 0.163$\pm$0.009 0.261$\pm$0.013 99%
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 A.9 Visualization
@@
 A.9.2 Visualization showcase
-To provide a distinct comparison among different models, we present visualizations of the forecasting results on the ETTh2 dataset and the weather dataset in different settings, as shown in Figures 12 to Figures 15, given by the following models: DLinear [Zeng et al., 2023], TimesNet [Wu et al., 2022], iTransfomrer [Liu et al., 2023], and PatchTST [Nie et al., 2022].
+To provide a distinct comparison among different models, we present visualizations of the forecasting results on the ETTh2 dataset and the weather dataset in different settings, as shown in Figures 12 to Figures 15, given by the following models: DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2022), iTransfomrer (Liu et al., 2023), and PatchTST (Nie et al., 2022).
 Among the methods, ROSE demonstrates the most accurate prediction ability.
@@
 A.10.1 Full-shot results
+Table 12 shows the full results of ROSE in full-shot setting for four prediction lengths. ROSE shows the advantage over the specific models and LLM-based models trained with the full training set.
 Table 12: Full results in full-shot setting.
-Models ROSE ITransformer PatchTST Timesnet Dlinear GPT4TS S2P-LLM
+Models ReadyTS ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$IP-LLM
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 A.10.2 Few-shot results
+Table 13 shows the full results of ROSE in 10% few-shot setting for four prediction lengths. ROSE shows the advantage over the specific models and LLM-based models trained with the 10% training set.
 Table 13: Full results in 10% few-shot setting
-Models ROSE ITransformer PatchTST Timesnet Dlinear GPT4TS S2P-LLM
+Models ReadyTS ITransformer PatchTST Timesnet Dlinear GPT4TS $\textbf{S}^{2}$IP-LLM
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 Novelty of decomposed frequency learning.
-Frequency masking is not a new concept, but past approaches randomly mask frequencies of a single time series once  [Chen et al., 2023b, Zhang et al., 2023], which show limited forecasting effectiveness due to the lack of common pattern learning from heterogeneous time series that come from various domains.
+Frequency masking is not a new concept, but past approaches randomly mask frequencies of a single time series once  (Chen et al., 2023b; Zhang et al., 2023), which show limited forecasting effectiveness due to the lack of common pattern learning from heterogeneous time series that come from various domains.
 While the multi-frequency masking we proposed randomly mask either high-frequency or low-frequency components of a time series multiple times as the key to enable learning of common time series patterns, such as trends and various long and short term fluctuations.
-Moreover, different from utilizing frequency masking as a way of data augmentation to enhance the diversity of input data  [Chen et al., 2023b, Zhang et al., 2023], we combine multi-frequency masking with reconstruction task as a novel pre-training framework, that learns a universal and unified feature representation by comprehending the data from various frequency perspectives, thereby enabling it to learn generalized representations.
+Moreover, different from utilizing frequency masking as a way of data augmentation to enhance the diversity of input data  (Chen et al., 2023b; Zhang et al., 2023), we combine multi-frequency masking with reconstruction task as a novel pre-training framework, that learns a universal and unified feature representation by comprehending the data from various frequency perspectives, thereby enabling it to learn generalized representations.
 Difference between frequency-domain masking and time-domain noise addition.
 Multi-frequency masking and reconstruction are not equivalent to the pre-training methods of adding noise and denoising (Noise).
-Due to the sparsity of time series, the process of adding noise and denoising may potentially disrupt the information of original time series [Dong et al., 2024].
+Due to the sparsity of time series, the process of adding noise and denoising may potentially disrupt the information of original time series (Dong et al., 2024).
 In contrast, multi-frequency masking not only preserves the series from such disruption but also helps the model understand temporal patterns from a multi-frequency perspective, thereby helping the model to learn general features better.
@@
 MSE MAE MSE MAE MSE MAE MSE MAE
-ROSE 96 0.287 0.336 0.159 0.247 0.367 0.395 0.273 0.332
+ReadyTS 96 0.287 0.336 0.159 0.247 0.367 0.395 0.273 0.332
 192 0.331 0.362 0.217 0.287 0.399 0.416 0.334 0.376
@@
 avg 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.380
-Replace Multi-Frequency Masking Random Frequency Masking 96 0.330 0.370 0.170 0.262 0.391 0.399 0.303 0.359
+Replace
+Multi-Frequency Masking
+Random Frequency Masking 96 0.330 0.370 0.170 0.262 0.391 0.399 0.303 0.359
 192 0.352 0.392 0.232 0.298 0.406 0.430 0.356 0.395
@@
 avg 0.378 0.400 0.261 0.319 0.408 0.432 0.375 0.407
-Other pre-training tasks Aug 96 0.304 0.357 0.178 0.266 0.376 0.405 0.281 0.348
+Other
+pre-training tasks
+Aug 96 0.304 0.357 0.178 0.266 0.376 0.405 0.281 0.348
 192 0.343 0.379 0.254 0.318 0.409 0.429 0.345 0.389
@@
 A.10.4 Zero-shot results
+Table 15 shows the full results of ROSE and other foundation models in zero-shot setting for four prediction lengths. ROSE exhibits competitive performance.
 Table 15: Full results in zero-shot setting.